### v1

# Krylov Complexity in early universe

Abstract
The Lanczos algorithm offers a method for constructing wave functions for both closed and open systems based on their Hamiltonians. Given that the entire early universe is fundamentally an open system, we apply the Lanczos algorithm to investigate Krylov complexity across different phases of the early universe, including inflation, the radiation-dominated period (RD), and the matter-dominated period (MD). Notably, we find that Krylov complexity differs between the closed and open system approaches. To effectively capture the impact of potentials during the RD and MD phases, we analyze various inflationary potentials—including the Higgs potential, the $R^{2}$ inflationary potential, and chaotic inflationary potential—taking into account the violations of slow-roll conditions. This analysis is conducted in terms of conformal time through the preheating process. Our numerical results indicate that the evolution of Krylov complexity and Krylov entropy is remarkably similar across both methods, regardless of the potential under consideration. Additionally, we rigorously construct what is referred to as an open two-mode squeezed state, utilizing the second kind of Meixner polynomials. Based on this construction, we are the first to calculate the evolution equations for $r_{k}$ and $\phi_{k}$ as they relate to the scale factor. Our findings suggest that dissipative effects lead to a rapid decoherence-like behavior. Moreover, our results indicate that inflation behaves as a strongly dissipative system, while both the radiation-dominated and matter-dominated phases exhibit characteristics of weak dissipation. This research provides new insights into exploring the universe from the perspective of quantum information.
arxiv: 1234.56789
1 Introduction
The idea of complexity was initially introduced in the holographic framework Stanford:2014jda . In this framework, it was observed that the conformal field theory (CFT) reaches thermal equilibrium relatively quickly, while the corresponding bulk gravity (wormhole) takes a much longer time to reach the same state. To address this issue, Refs. Hartman:2013qma ; Liu:2013iza introduced the concept of computational complexity within the holographic framework. This concept suggests that while the emergence of spacetime is related to quantum entanglement, the evolution of space occurs through computational complexity.
The concept of complexity has become increasingly essential in high-energy physics even without the holographic principle. First, the so-called computational complexity describes the minimum number of logical steps required for a task Aaronson:2016vto . Furthermore, the method of calculating the complexity was first proposed by Nielson, $\it e.t.c$, dubbed as the geometrical method Nielsen:2005mkt ; Dowling:2006tnk ; Nielsen:2006cea . Another approach is highly related to the ”Fubini study” distance of the geometry of information Chapman:2017rqy . These two geometrical methods are highly dependent on the choice of the geometry of the parametric manifold. For more precisely capturing the information for the quantum state itself, the concept of Krylov complexity was proposed using the Lanczos algorithm viswanath1994recursio , which is explicitly related to the quantum state without needing the parametric manifold. Since this method is free of ambiguities compared with the previous two ones, there will be similarities and differences between them. According to the concept of Krylov complexity, the universal operator growth was proposed Parker:2018yvk . Once the precise quantum state is given, we could analyze its Krylov complexity, Krylov entropy, and Lanczos coefficient based on the Lanczos algorithm viswanath1994recursio . For the connection between the Krylov complexity and the circuit complexity, Ref. Aguilar-Gutierrez:2023nyk has shown that they cannot be compatible with each other. However, Ref. Caputa:2021sib demonstrated that the Krylov complexity was proportional to the distance of the ”Fubuni-Study” metric.
Recently, the Krylov complexity has been broadly applied to condensed matter and high-energy physics. Ref. Muck:2022xfc has systematically analyzed the Krylov complexity in various polynomials dubbed as the orthogonal basis. Due to the importance of the SKY model in high energy physics, the Krylov complexity can also be implemented into this field Rabinovici:2020ryf ; Jian:2020qpp ; He:2022ryk . Meanwhile, the Krylov complexity can be applied to the generalized coherent state Patramanis:2021lkx , Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , conformal field theory Dymarsky:2021bjq ; Caputa:2021ori , topological phases of matter Caputa:2022eye and even extend into the quantum field theory He:2024hkw ; He:2024xjp . Ref. Dymarsky:2019elm proved that the chaos can be understood as a delocalization in Krylov complexity. Furthermore, one could observe the exponential growth of Krylov complexity n integrable systems with saddle-dominated scrambling, whose universal features can be verified by Huh:2023jxt . More recent developments of Krylov complexity can be found in Refs. Erdmenger:2023wjg ; Hashimoto:2023swv ; Vasli:2023syq ; Gill:2023umm ; Bhattacharjee:2023uwx ; Adhikari:2022whf ; Loc:2024oen ; Caputa:2024vrn ; Basu:2024tgg ; Sasaki:2024puk ; Caputa:2024xkp ; Bhattacharjee:2022qjw ; Sahu:2024opm ; Kim:2021okd ; Chen:2024imd ; Bhattacharjee:2024yxj ; Sanchez-Garrido:2024pcy ; Chattopadhyay:2024pdj . Especailly, the concept of Krylov complexity can be extended into in various fields Nandy:2024htc .
We aim to investigate the Krylov complexity for the whole early universe, including the inflationary period, radiation-dominated period (RD), and matter-domination period (MD). Some research has examined the computational complexity for inflation Choudhury:2020hil ; Bhargava:2020fhl ; Lehners:2020pem ; Bhattacharyya:2020rpy ; Adhikari:2021ked , where the analysis demonstrated that there is oscillation at the very beginning of inflation, and the evolution was similar in post inflation period. Once considering the quantum effects via various quantum frameworks, the complexity will show the irregular oscillation during inflation providing a new method for distinguishing various inflationary models Li:2023ekd . Quite a few researches are investigating the Krylov complexity in the early universe. Ref. Adhikari:2022oxr is the first one who investigate Krylov complexity in inflation by evaluating the values of non-trivial sound speed, where their method is based on the closed system’s method. However, our universe is an open system. In light of the open system’s method, Ref. Li:2024kfm investigated the Krylov complexity within the framework of modified dispersion relation in inflation, indicating that the inflation behaves as a strong dissipative system and the overall trend of Krylov complexity is increasing. The key ingredient for exploring the Krylov complexity with the open system’s method is constructing the wave function in light of Ref. Bhattacharya:2022gbz , whose wave function wave from Li:2024kfm will serve as our starting point in this work.
In previous studies, Krylov complexity was applied to the early universe using the thermal state. Refs Li:2024ljz ; Li:2024iji examined Krylov complexity by considering the contribution of the potential (quadratic term) and the violation of slow-roll conditions. The potential was approximated to be the quadratic term around the minimum in our previous work Li:2024ljz , which may not capture fully precise information. In the upcoming work, we plan to include more models such as the Higgs potential Higgs:1964pj , the potential of Starobinsky inflation ($R^{2}$ inflation) Starobinsky:1980te , and the potential of chaotic inflation Linde:1983gd . These models can cover most inflationary models, especially those favored by current observations, which indicate that the potential shape is a plateau. It has been suggested by Ref. Galante:2014ifa that many inflationary models can be unified into the so-called $\alpha$ attractors, whose potential is similar to $R^{2}$ inflationary potential. Additionally, Ref. Albrecht:1992kf demonstrated a relation between the Bunch-Davies vacuum and the two-mode squeezed state. We will continue to use the two-mode squeezed state to investigate the Krylov complexity for the entire early universe in light of these two factors.
This paper will be organized as follows. In Sec. 2, we will introduce the Lanczos algorithm. Sec. 3 will give a brief introduction to the early universe, including inflation, RD, and MD, and three different kinds of potentials. Sec. 4 will give the two-mode squeezed state, and the numerical solutions of $r_{k}$ in different periods. Sec. 5 will investigate the Krylov complexity in a method of closed system. Sec. 5.2 will investigate the Krylov entropy in a method of closed system. Sec. 6 will give the Lanczos coefficient and Krylov complexity with the approach of open system, where we are the first to derive the evolution equation of $\phi_{k}$ and $r_{k}$ based on the open two-mode squeezed state. Finally, the conclusion and outlook will be given in Sec. 7.
2 Lanczos algorithm with the approach of closed system
In this section, we will mainly introduce the Lanczos algorithm with the approach of closed system. First, the operator $\mathcal{O}(t)$ in Heisenberg picture needs to be defined as follows,
$\partial_{t}\mathcal{O}(t)=i[H,\mathcal{O}(t)]$ (1)
where H is the Hamilton and its solution is
$\mathcal{O}(t)=e^{iHt}\mathcal{O}e^{-iHt}$ (2)
Next, let’s define the Liouvillian super-operator $\mathcal{L}_{X}$ as $\mathcal{L}_{X}Y=[X,Y]$, where [,][,][ , ] denotes the commutator. We apply this to Eq. (2), the operator can also be expressed as
$\mathcal{O}(t)=e^{i\mathcal{L}t}\mathcal{O}=\sum_{n=0}^{\infty}\frac{(it)^{n}}{n!}\mathcal{L}^{n}\mathcal{O}(0)=\sum_{n=0}^{\infty}\frac{(it)^{n}}{n!}\tilde{\mathcal{O}}_{n}$ (3)
where $\mathcal{L}^{n}\mathcal{O}=\tilde{\mathcal{O}}_{n}=[H,\tilde{O}_{n-1}]$ which can serve as the basis for the Hilbert space.
$\begin{split}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}),\mathcal{L}^{1}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}_{1}),\mathcal{L}^{2}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}_{2}),\mathcal{L}^{3}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}_{3})...\end{split}$ (4)
However, there are not orthogonal bases. In order to construct an orthogonal basis, we use the Lanczos algorithm. The first two operators become
$\mathcal{O}_{0}=|\tilde{\mathcal{O}}_{0})=\mathcal{O},~{}~{}~{}|\mathcal{O}_{1})=b_{1}^{-1}\mathcal{L}|\mathcal{O}_{0})$ (5)
where $b_{1}=\sqrt{(\tilde{\mathcal{O}}_{0}\mathcal{L}|\mathcal{L}\tilde{\mathcal{O}}_{0})}$ desribing the normalized vector. And the next orthogonal basis can be written by
$|\mathcal{O}_{n})=b_{n}^{-1}|A_{n})$ (6)
with
$|A_{n})=\mathcal{L}|\mathcal{O}_{n-1})-b_{n-1}|\mathcal{O}_{n-2}),~{}~{}~{}b_{n}=\sqrt{(A_{n}|A_{n})}$ (7)
where $b_{n}$ represents the Lanczos coefficients. This iterative relation will stop and produce the finite orthogonal Krylov basis if $b_{n}=0$. Subsequently, we can express Eq. (3) as follows,
$\mathcal{O}(t)=e^{i\mathcal{L}t}\mathcal{O}=\sum_{n=0}^{\infty}(i)^{n}\phi_{n}(t)|\mathcal{O}_{n})$ (8)
[…]
$\partial_{t}\phi_{n}(t)=b_{n}\phi_{n-1}-b_{n-1}\phi_{n+1}$ (9)
Next, we can define the Krylov complexity using the wave function $\phi_{n}$ as follows
$K=\sum_{n}n|\phi_{n}|^{2}$ (10)
Additionally, the Lanczos coefficient can be bounded via Parker:2018yvk ,
$b_{n}\leq\alpha n+\eta$ (11)
where $\alpha$ and $\eta$ encodes different information in various models. It should be noticed that $b_{n}=\alpha n+\eta$, meaning that the system is a maximal chaotic dynamical system. For the case of maximal chaos, we could build the relation between the Lyapunov index and $\gamma$ and $\alpha$ as follows,
$\lambda=2\alpha$ (12)
In this work, we will investigate the Krylov complexity for the entire early universe, thus the formula of scale factor $a(\eta)$ should be introduced.
3 Some setup of early universe
In this section, we will first give some basis of scale factor in distinctive periods. Following the notation of our previous work Li:2024ljz , one of the core goal of this work is examining the impact of inflationary potential for the Krylov complexity in RD and MD due to the violation of slow-roll conditions. The Mukhannov-Sasaki variable contains the contribution of potential which cannot show the explicit contribution to the Krylov complexity. That is why we just perturb the inflaton field explicitly in this paper. Thereafter, we will give some very important inflationary potentials, namely, the chaotic inflationary potential, Starobinksy’s inflationary potential ($R^{2}$ inflation) and Higgs potential.
3.1 Inflation, RD and MD
We will follow the notation of Baumann:2009ds to introduce the scale factor. The background metric is the Friedman-Lemaitre-Robertson-Walker metric,
$ds^{2}=a^{2}(\eta)(-d\eta^{2}+d\vec{x}^{2}),$ (13)
where $a(\eta)$ is the scale factor. Here, we will assume a single component universe, which means that the universe will experice the inflation, RD and MD, respectively. We could use equation of state to characterize the universe via
$w_{I}=\frac{P_{I}}{\rho_{I}},$ (14)
where $P_{I}$ and $\rho_{I}$ is the pressure and energy density of the corresponding periods parameterized by $I$. For obtaining the relation between the conformal time and the scale factor, we need the comoving distance,
$\chi_{ph}(\eta)=\int_{t_{i}}^{t}\frac{dt}{a}=\int_{\ln a_{i}}^{\ln a}(aH)^{-1}d\ln a$ (15)
where $\chi_{ph}(\eta)$ is the comoving distance in the event horizon of particle. In light of $w_{I}$, the Hubble radius can be replaced by
$(aH)^{-1}=H_{0}^{-1}a^{\frac{1}{2}(1+3\omega)}$ (16)
where we use $\omega$ to take place of $\omega_{I}$ for simplify. Integrating equation (15) yields the relation between the scale factor and the conformal time
$\eta=\frac{2H_{0}^{-1}}{(1+3\omega)}a^{\frac{1}{2}(1+3\omega)}$ (17)
Consequently, we can derive the relation for various time periods
$\eta=\begin{cases}&-(aH_{0})^{-1},\ \ (\omega=-1),\ \ \rm inflation\\ &aH_{0}^{-1},\ \ (\omega=\frac{1}{3}),\ \ \rm RD\\ &2\sqrt{a}H_{0}^{-1},\ \ (\omega=0),\ \ \rm MD\end{cases}$ (18)
One could see that the various values of $\omega$ describe the distinctive periods.
In the following text, we will discuss the fundamental basis for the different periods in the evolution of the universe. Initially, the universe is created by a singularity, followed by a period of exponential expansion known as inflation. During this phase, the curvature perturbation transitions from the quantum level to the classical level after the horizon exits, leaving an imprint on the Cosmic Microwave Background (CMB). Subsequently, the universe enters the Radiation Dominated (RD) era, characterized by an energy density given by the equation $\rho_{\rm rad}(t)=\rho_{\rm rad_{0}}\frac{a_{0}^{4}}{a(t)^{4}}$, where $\rho_{\rm rad_{0}}$ and $a_{0}$ represent the initial values of the energy density and the scale factor of RD. During the RD era, particles are relativistic, but as the temperature decreases, they become increasingly non-relativistic. This leads the universe into the matter-dominated (MD) epoch, where the energy density is given by $\rho_{\rm mat_{0}}(t)=\rho_{\rm mat_{0}}\frac{a_{0}^{3}}{a(t)^{3}}$. The transition from the RD to the MD epoch occurs when the energy density of radiation becomes equal to the energy density of matter. Once entering MD, there is a process called preheating where the particles will be generated via this process. First, we give the full action for the single field inflation,
[…]
where $V(\phi)$ is the potential of inflation. Following the method of Kofman:1997yn ; Shtanov:1994ce , the equation of motion for background $\phi(t)$ in RD and MD is
$\ddot{\phi}(t)+3H\dot{\phi}(t)+V(\phi),_{\phi}=0,$ (20)
where $\ddot{\phi}(t)=\partial_{t}^{2}\phi(t)$ and the similar work can be found in our previous works Liu:2019xhn ; Liu:2021rgq . Defining with the effective mass $m_{\rm eff}=\frac{d^{2}V}{d\phi^{2}}$ and changing variable as $\tilde{\phi}(t)=a^{3/2}\phi(t)$, Eq. (20) will become
$\ddot{\tilde{\phi}}(t)+m_{\rm eff}^{2}\tilde{\phi}(t)=0,$ (21)
the solution of Eq. (21) is $\phi(t)=a^{-3/2}\cos(m_{\rm eff}t)$ where it is the solution of $\phi(t)$ in preheating period and we have neglecte the term of $\frac{9}{4}H^{2}\tilde{\phi}+\frac{3}{2}\dot{H}\tilde{\phi}$ since it is much smaller compared with potential. It should be noticed that the solution is in terms of physical time $t$. In this paper, we will use $d\eta=\frac{dt}{a}$ to transform into the conformal time for the latter investigations, which is denoted by Eq. (18).
3.2 Various inflationary potentials
There are lots of inflationary models that represent distinctive inflationary potentials. Current observations favor the concave potentials Planck:2018jri , where the potential of $R^{2}$ inflation belongs to this potential. Additionally, the $R^{2}$ inflation agrees with the observations from inflation to large scales. From the theoretical aspect, we could see that many inflationary models can be unified into the so-called $\alpha$ attractors inflationary models Galante:2014ifa , where the other models and $R^{2}$ inflationary model belong to the same kind of inflationary models. Being armed with these two reasons, we will introduce $R^{2}$ inflation. According to the current standard model, the Higgs field is the only scalar field in nature. Therefore, the Higgs potential is our other option Higgs:1964pj . The final choice is the chaotic inflation whose potential is the quadratic term Linde:1983gd . Finally, we could see that our choices could cover the most inflationary models.
Here, we list the formulas of distinctive inflationary potentials as follows,
$\begin{cases}&\ V_{1}=\frac{1}{2}m^{2}\phi^{2},\\ &\ V_{2}=\frac{3}{4}M^{2}M_{p}^{2}(1-e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}})^{2},\\ &\ V_{3}=\frac{1}{2}m^{2}\phi^{2}+\frac{1}{4}\lambda\phi^{4},\end{cases}$ (22)
where $M_{p}=(8\pi G)^{-\frac{1}{2}}$ is the definition of the reduced Planck mass. Additionally, $M$ represents the energy scale, and its value can be determined based on the amplitude of the observed power spectrum for primordial perturbations. Being armed with the solution of Eq. (20) and $dt=ad\eta$, we could rewrite $V_{\phi\phi}$ as
$\begin{cases}&\ V_{1,\phi\phi}=m^{2},\\ &\ V_{2,\phi\phi}=M^{2}\Bigl{(}-e^{-\sqrt{\frac{2}{3}}\frac{Aa^{-\frac{3}{2}}\sin(m_{\rm eff}a\eta)}{M_{p}}}+2e^{-2\sqrt{\frac{2}{3}}\frac{Aa^{-\frac{3}{2}}\sin(m_{\rm eff}a\eta)}{M_{p}}}\Bigr{)},\\ &\ V_{3,\phi\phi}=m^{2}+3\lambda A^{2}a^{-3}\sin^{2}(m_{\rm eff}a\eta)\end{cases}$ (23)
where $\phi(t)=Aa^{-\frac{3}{2}}\sin(m_{\text{eff}}t)$. For the effective mass $m_{\text{eff}}$, we have $m_{\text{eff}}^{2}=\frac{d^{2}V(x)}{dx^{2}}$, and it is evident that $m_{\text{eff}}=\sqrt{V_{i,\phi\phi}}$ in our work $(i=1,2,3)$. According to effective mass for the distinctive inflationary models, we need the iterations for $V_{2}$ and $V_{3}$ in terms of $\log_{10}a$ followed by Li:2024kfm . To determine whether the functions $V_{2}$ and $V_{3}$ converge within the defined interval, we performed iterations at multiple sampling points ranging from $\log_{10}a=0$ to $\log_{10}a=6$ with the step $0.05$. Our findings reveal that all points ranging from $\log_{10}a=0$ to $\log_{10}a=6$ are all convergent.
Figure 1: The numeric of $V_{,\phi\phi}$ at first 50 iterations of potential $V_{2}$ and $V_{3}$ in $\log_{10}a=0.2$ and $\log_{10}a=2.8$, in which we set $M_{p}=1$, $m=10^{-6}$; $M^{2}=10^{-12}$ at $V_{2}$, and $\lambda A^{2}=10^{-4}$ at $V_{3}$.
As an illustration, Fig. 1 clearly shows the iterations for potential $V_{2}$ and $V_{3}$, where we have performed $50$ times iterations for two specific values of $\log_{10}a$. Let us mention the range of parameters as shown in Fig. 1, the potential can be constraint by the COBE normalization Planck:2015sxf , whose values are bounded below $10^{-10}$ in Planck units, and the inflationary field is a large field and its maximal value is around $100$. Consequently, the effective mass for various potentials is around $10^{-6}$ and other parameters are set according to this logic.
Figure 2: The numeric of $V_{,\phi\phi}$ in terms of $\log_{10}a$ for potential $V_{2}$, and $V_{3}$ in RD and MD, in which we replace its exact value with the third iteration of $V_{2}$ and $V_{3}$. We set $\theta_{0}=0$, $M_{p}=1$, $m=10^{-6}$; $M^{2}=10^{-12}$ at $V_{2}$, and $\lambda A^{2}=10^{-4}$ at $V_{3}$.
Fig. 2 depicts the evolution of $V_{,\phi\phi}$ with respect to $\log_{10}a$ in RD and MD, using the third iteration of $V_{2}$ and $V_{3}$ via Fig. 1.The evolution of inflation is neglected due to the violation of slow-roll conditions. From Fig. 2, we could see that the effective mass of $R^{2}$ inflation will decrease in RD and then it will grow to the constant ($10^{-12}$) with the oscillations. For the potential of the Higgs field, it is different which shows that there is a peak in RD and after that, it will decrease into some constants. But, it indicates that the effective mass of $V_{2}$ and $V_{2}$ are of almost the same order based on the COBE normalization Planck:2015sxf .
In this section, we provide an introduction to inflation, RD, and MD, which are characterized by the corresponding scale factor. We then present three essential inflationary potentials that encompass most types of inflationary models. Additionally, we provide an overview of preheating, during which the background of the inflaton field oscillates around the minimum. Subsequently, we incorporate the background solution for the inflaton into the numerical calculation of the effective mass for these three distinct potentials.
4 Evolution of two-mode squeezed state
In this section, we will follow the method of Li:2024kfm for investigating the evolution of $r_{k}$ and $\phi_{k}$ in inflation, RD and MD. We have evaluated the quantum circuit complexity in various inflationary frameworks Liu:2021nzx ; Li:2021kfq ; Li:2023ekd .
4.1 Two-mode squeezed state
For obtaining the two-mode squeezed state, we should introduce the unitary operator as follows,
$\mathcal{U}_{k}=\hat{\mathcal{S}}_{k}(\phi_{k},r_{k})\hat{\mathcal{R}}_{k}(\theta_{k})$ (24)
with
$\hat{\mathcal{S}}_{k}=\exp\bigg{[}r_{k}(\eta)\big{(}e^{-2i\phi_{k}(\eta)}\hat{c}_{\vec{k}}\hat{c}_{-\vec{k}}-e^{2i\phi_{k}(\eta)}\hat{c}^{\dagger}_{-\vec{k}}\hat{c}_{\vec{k}}^{\dagger}\big{)}\bigg{]}$ (25)
^$\hat{\mathcal{R}}_{k}(\theta_{k})=\exp\bigg{[}-i\theta_{k}(\eta)\big{(}\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}+\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}\big{)}\bigg{]}$ (26)
where $\hat{\mathcal{S}}_{k}$ is the squeezed parameter and ^$\hat{\mathcal{R}}_{k}(\theta_{k})$ is the rotation parameter that will be neglected since it only contributes the phase factor of wave function. Then, by applying $\hat{S}_{\vec{k}}(r_{k},\phi_{k})$ to the vacuum state, we can express the wave function of the two-mode squeezed state as follows
$|\psi\rangle=\hat{S}_{\vec{k}}(r_{k},\phi_{k})|0;0\rangle_{\vec{k},-\vec{k}}=\frac{1}{\cosh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}e^{2in\phi_{k}}\tanh^{n}r_{k}|n;n\rangle_{\vec{k},-\vec{k}}.$ (27)
Being armed with this wave function, we could proceed the numeric of $\phi_{k}$ and $r_{k}$. But the evolution of $\phi_{k}$ and $r_{k}$ is based on the total action (19).
4.2 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$
In this work, we will evaluate the impact of potential for the Krylov complexity, where the Mukhanov-Sasaki variable contains the contribution of potential. Consequently, we cannot observe the explicit influence from potential. Therefore, we will directly perturb the total action into the second order,
[…]
$\pi(\eta,\vec{x})=\frac{\delta L}{\delta f^{\prime}(\eta,\vec{x})}=f^{\prime}-\frac{a^{\prime}}{a}f.$ (29)
According to Eq. (28) and Eq. (29), we could obtain
[…]
Using the Fourier decomposition of operator $\hat{f}(\eta,\vec{x})$ and $\hat{\pi}(\eta,\vec{x})$,
$\hat{f}(\eta,\vec{x})=\int\frac{d^{3}k}{(2\pi)^{3/2}}\sqrt{\frac{1}{2k}}(\hat{c}_{\vec{k}}^{\dagger}f_{\vec{k}}^{*}(\eta)e^{-i\vec{k\cdot}\vec{x}}+\hat{c}_{\vec{k}}f_{\vec{k}}e^{i\vec{k\cdot}\vec{x}}),$ (31)
$\hat{\pi}(\eta,\vec{x})=i\int\frac{d^{3}k}{(2\pi)^{3/2}}\sqrt{\frac{k}{2}}(\hat{c}_{\vec{k}}^{\dagger}u_{\vec{k}}^{*}(\eta)e^{-i\vec{k\cdot}\vec{x}}-\hat{c}_{\vec{k}}u_{\vec{k}}e^{i\vec{k\cdot}\vec{x}}).$ (32)
Substituting equations (31) and (32) into equation (30) to obtain the Hamiltonian operator,
$\hat{H}_{k}=(k+\frac{a^{2}V_{,\phi\phi}}{2k})(\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}+\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{\vec{-k}})+(\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a})\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{\vec{k}}^{\dagger}+(\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a})\hat{c}_{-\vec{k}}\hat{c}_{\vec{k}}.$ (33)
[…]
$i\frac{d}{d\eta}|\psi\rangle=\hat{H}_{k}|\psi\rangle,$ (34)
we could obtain the evolution of equation of $r_{k}$ and $\phi_{k}$
$\begin{split}&r^{\prime}_{k}=\frac{a^{2}V_{,\phi\phi}}{2k}\sin{2\phi_{k}}-\frac{a^{\prime}}{a}\cos{2\phi_{k}},\\ &\phi_{k}^{{}^{\prime}}=-k-\frac{a^{2}V_{,\phi\phi}}{2k}+(\frac{a^{\prime}}{a}\sin{2\phi_{k}}+\frac{a^{2}V_{,\phi\phi}}{2k}\cos{2\phi_{k}})\coth{2r_{k}}.\end{split}$ (35)
In inflation, Eq. (35) will nicely recover into the case of
$\begin{split}&r_{k}^{{}^{\prime}}=-\frac{a^{\prime}}{a}\cos(2\phi_{k}),\\ &\phi_{k}^{{}^{\prime}}=-k+\frac{a^{\prime}}{a}\sin(2\phi_{k})\coth(2r_{k}).\end{split}$ (36)
It is very difficult to obtain the numeric of Eq. (35), thus we could make a variable transformation, $y=\log_{10}a$, to numerically solve Eq. (35),
$\begin{split}\frac{dr_{k}}{dy}=&\frac{\ln(10)10^{\frac{y}{2}(5+3\omega)}V_{,\phi\phi}}{2kH_{0}}\sin{2\phi_{k}}-\ln(10)\cos{2\phi_{k}}\\ \frac{d\phi_{k}}{dy}=&-\frac{\ln(10)10^{\frac{y}{2}(1+3\omega)}k}{H_{0}}-\frac{\ln(10)10^{\frac{y}{2}(5+3\omega)}V_{,\phi\phi}}{2kH_{0}}+\Bigl{(}\ln(10)\sin{2\phi_{k}}\\ &+\frac{\ln(10)10^{\frac{y}{2}(5+3\omega)}V_{,\phi\phi}}{2kH_{0}}\cos{2\phi_{k}}\Bigr{)}\coth{2r_{k}}\end{split}$ (37)
where we have utilized Eq. (17) and $\frac{a^{\prime}}{a}=H_{0}a^{-\frac{1}{2}(1+3\omega)}$.
Figure 3: The numerical solution of $R_{k}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$. $k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.. The effect of potential is always ignored in slow-roll inflation. But in RD and MD periods, the influence can be seen by using different potential energies as shown in these panels.
Before analyzing the figure of $r_{k}$, it’s important to note that the Krylov complexity depends only on $r_{k}$. Therefore, we only provide the numerical values of $r_{k}$ in Fig. 3 based on Eq. (37). In this equation, we have set $H_{0}=1$ for simplicity and used various values of $k$ in different time periods. The first panel of Fig. 3 illustrates the evolution of $r_{k}$ during inflation. There is an oscillation before the horizon exits, and then $r_{k}$ will grow linearly, consistent with the standard case of Li:2024kfm . As we transition to the radiation-dominated era (RD), we observe that $r_{k}$ will increase to some constant value that oscillates around this value. Interestingly, different potentials produce almost the same trend for $r_{k}$, indicating that the specific potential does not significantly impact the evolution of $r_{k}$. This observation also applies to the matter-dominated era (MD), where the evolution trend is similar to RD. However, we notice that varying momentum significantly affects $r_{k}$, leading to oscillations as the value of $k$ decreases. Although we cannot distinguish between the various potentials using Fig. 3, our numerical data still demonstrates that the evolution of $r_{k}$ is significantly different whether there is a potential or not. It’s worth noting that we cannot provide the values of $r_{k}$ without a potential, as this is not physically meaningful. The potential cannot be ignored due to the violation of slow-roll conditions.
In this section, we have analyzed the evolution of $r_{k}$ for different inflationary potentials within a single-field inflationary framework. Our research indicates that the value of $r_{k}$ increases during inflation. The behavior of $r_{k}$ in radiation-dominated (RD) and matter-dominated (MD) eras is similar and will oscillate around constant values. Our numerical analysis also shows that various inflationary models produce a similar trend for $r_{k}$. Additionally, we have observed that the scale factor and momentum have a significant impact on the evolution of $r_{k}$. All of the details are encoded in Fig. 3.
5 Krylov complexity in closed system
In this section, we will investigate the Krylov complexity in the method of closed system Liu:2021nzx ; Li:2021kfq ; Li:2023ekd , consisting of the Krylov complexity and Krylov entropy.
5.1 Krylov complexity
For obtaining the Krylov complexity, the most essential quantity is the wave function. First, the two-mode squeezed state wave function can be represented by Krylov basis as follows,
$\mathcal{O}(t)=\sum_{n=0}^{\infty}(i)^{n}\phi_{n}(t)|\mathcal{O}_{n})=\frac{1}{\cosh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}e^{2in\phi_{k}}\tanh^{n}r_{k}|n;n\rangle_{\vec{k},-\vec{k}}.$ (38)
Then, we could find the explicit wave function $\phi_{n}$
$\phi_{n}=\frac{1}{\cosh{r_{k}}}(-e^{2i\phi_{k}}\tanh{r_{k}})^{n}.$ (39)
We can define the Krylov compplexity in terms of $\phi_{n}$
$K=\sum_{n}n\left|\phi_{n}\right|^{2}\\ =\cosh^{2}r_{k}\tanh^{2}r_{k}=\sinh^{2}r_{k}.$ (40)
It is explicitly showing that the Krylov complexity is independent of $\phi_{k}$, that is why we only list the figure of $r_{k}$.
Figure 4: The numerical of Krylov complexity in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$. $k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.
Being armed with Eq. (40), we could give its numeric in Fig. 4, indicating the evolution of Krylov complexity in inflation, RD, and MD. Fig. 4 clearly shows that the Krylov complexity is enhanced in inflation which is consistent with the standard case of the modified dispersion within the framework Li:2024kfm . Similar work investigated the Krylov complexity for the exponential expansion for de-Sitter spacetime corresponding to the inflation Adhikari:2022oxr . Ref. Adhikari:2022oxr mentioned that the Krylov complexity equals the average particle number proportional to the volume. Thus we could see that this simple analysis is consistent with the CV conjuncture Stanford:2014jda , in which the complexity is equaled to the volume of wormhole under the framework of holography. However, the method of Adhikari:2022oxr is for the closed system.
In the case of RD and MD, it is observed that the Krylov complexity will increase to certain constant values that fluctuate around this value. Additionally, different patterns of inflationary models tend to produce similar trends of Krylov complexity in different periods. Our analysis aligns with the findings of Barbon:2019wsy , where the Krylov complexity will saturate some constant values even after the scrambling time in a closed system. The scrambling time will be assessed in the context of an open system, particularly with regard to the Lanczos coefficient, Lyapunov index, $\it e.t.c$.
5.2 Krylov Entropy
In a quantum system, entropy is defined as a measure of the disorder of a system. The Krylov entropy (K-entropy) can be used to test the level of disorder of curvature perturbation. Following Barbon:2019wsy , the K-entropy has the following form
$\begin{split}S_{K}&=-\sum_{n=0}^{\infty}\left|\phi_{n}\right|^{2}\ln\left|\phi_{n}\right|^{2}\\ &=\cosh^{2}r_{k}\ln(\cosh^{2}r_{k})-\sinh^{2}r_{k}\ln(\sinh^{2}r_{k})\end{split}$ (41)
where $\phi_{n}$ is the wave function (27). Once given the analytic formula of K-entropy, we could numerically simulate it in various epochs as shown in Fig. 5.
Figure 5: The numerical of Krylov entropy in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$. $k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.
Fig. 5 indicates a complete evolution of K-entropy in inflation, RD and MD, which is similar to Krylov complexity. We could see that the curvature of perturbation will be more and more disorder in inflation. For RD and MD, the disorder will be enhancing and then saturates to some constant values and the total trend is enhancing, where the trend is also highly related the scale factor. From another aspect, the preheating will generate the particles occurred at MD that leads to the growth of K-entropy.
Let’s discuss Section 5. In this section, we provide a comprehensive analysis of Krylov complexity throughout the early universe. Our numerical data clearly indicates that different potentials yield almost identical patterns of Krylov complexity and K-entropy. Specifically, Krylov complexity increases during inflation and then stabilizes at constant values during RD and MD. This same behavior is observed for K-entropy. Additionally, we must examine the Lanczos coefficient (equivalent to the Lyapunov index in maximal chaos), which remains consistent in both the closed and open system methods.
6 Krylov complexity with the approach of open system
The preheating process in MD and RD is a highly non-equilibrium process that creates particles Kofman:1997yn ; Shtanov:1994ce . In addition, Ref. Cheung:2007st demonstrated that the universe’s evolution breaks the time invariance when the temporal direction is reversed, leading to a violation of energy conservation. Due to these reasons, the open system approach is considered to be more realistic compared to the closed system method. We will follow the method of Bhattacharya:2022gbz to investigate the Krylov complexity.
6.1 Lanczos coefficient and dissipation coefficient
Firstly, the general operator in Heisenberg’s picture has the following form,
$\mathcal{O}(\eta)=e^{i\mathcal{L}_{o}\eta}$ (42)
where the Lindbladian $\mathcal{L}_{o}$, also known as the Hamiltonian, acts on the Krylov basis, we can obtain
$\mathcal{L}_{o}|\mathcal{O}_{n})=-ic_{n}|\mathcal{O}_{n})+b_{n+1}|\mathcal{O}_{n+1})+b_{n}|\mathcal{O}_{n-1})$ (43)
where $c_{n}$ encodes the information of the open system and $b_{n}$ is the Lanczos coefficient. Due to this extra term, we could divide the total Hamiltonian (also known as Lindbladian) into two parts which can be represented by
$\mathcal{L}_{o}=\mathcal{H}_{k}=\mathcal{H}_{close}+\mathcal{H}_{open}$ (44)
In our work, $\mathcal{L}_{o}$ is the Hamiltonian Eq. (33). In light of (43), we can easily obtain the open part (generating $c_{n}$) of Hamiltonian,
$\mathcal{H}_{open}=(k+\frac{a^{2}V_{,\phi\phi}}{2k})(\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}+\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger})$ (45)
and the close part (generating $b_{n}$ and $b_{n-1}$) is
$\mathcal{H}_{close}=(\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a})\hat{c}_{\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}^{\dagger}+(\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a})\hat{c}_{\vec{k}}\hat{c}_{-\vec{k}}$ (46)
where $|\mathcal{O}_{n})=\Bigl{|}n_{\vec{k}};n_{-\vec{k}}\Bigr{\rangle}$. By substituting Eq. (45) and Eq. (46) into Eq. (43), we can obtain the equations for $c_{n}$ and $b_{n}$ as
$c_{n}=i(2n+1)(k+\frac{a^{2}V_{,\phi\phi}}{2k}).$ (47)
$b_{n}=n\sqrt{[(\frac{a^{2}V_{,\phi\phi}}{2k})^{2}+(\frac{a^{\prime}}{a})^{2}]}.$ (48)
Here, we will discuss more about $b_{n}$ and $c_{n}$. Ref. Bhattacharjee:2022lzy has introduced the $a_{n}$ corresponding to $c_{n}$ in our work and also stated that $c_{n}$ is highly related to the dissipative coefficient. Meanwhile, we could see that $b_{n}$ is the same for the closed system and open system since it is precisely determined by the Hamiltonian.
In Ref. Parker:2018yvk , it has clearly indicated that the dynamical system is an infinite, non-integrable, many-body system as $b_{n}=\alpha n+\gamma$, where $\alpha$ and $\gamma$ are two parameters determined by various models. In our case, we could see that $b_{n}\propto n$, thus the whole early universe is an ideal chaotic system.
6.1.1 $b_{n}$, $c_{n}$ and $\mu_{2}$
[…]
$\partial_{\eta}\phi+2b_{n}\partial_{n}\phi+\tilde{c}_{n}\phi=0,$ (49)
where we have defined $\tilde{c}_{n}=-ic_{n}$, meanwhile we have used the $b_{n+1}\approx b_{n}$ and $\phi_{n+1}-\phi_{n-1}\approx 2\partial_{n}\phi$ in the continous limit. $b_{n}$ and $c_{n}$ are defined in Eq. (47) and Eq. (48), respectively. The generic formulas of $b_{n}$ and $c_{n}$ are
$b_{n}^{2}=|1-\mu_{1}^{2}|n(n-1+\beta),\ \ c_{n}=i\mu_{2}(2n+\beta)$ (50)
where $\mu_{1}$, $\mu_{2}$ and $\beta$ are determined by various models and $\mu_{2}$ plays a role of dissipative coefficient as shown in Ref. Li:2024kfm . In our work, one can explicitly find $\beta=1$ and combine Eqs. (47), (48), and (50), we find that the parameters $\mu_{1}$ and $\mu_{2}$ have the following forms,
$|1-\mu_{1}^{2}|=(\frac{a^{2}V_{,\phi\phi}}{2k})^{2}+(\frac{a^{\prime}}{a})^{2},\mu_{2}=k+\frac{a^{2}V_{,\phi\phi}}{2k}.$ (51)
It should be noticed that $\mu_{2}$ is mainly determined by $k$ in inflation due to the slow-roll conditions. Once entering RD and MD (taking the contribution of the effective mass into account), we could see that it will be smaller than unity since the effective mass of these three types potentials is much more tiny compared with momentum $k$. First, we discuss the Lanczos coefficient $b_{n}$. Fig. 6 clearly indicates the three distinctive potentials of $b_{n}$ in inflation, RD and MD, in which the parameters of $k$, $H_{0}$ are set to be the same with the previous plots. As Bhattacharya:2022gbz confirmed, the Lyapunov index is equavillent to $b_{n}$ whose relation is $\lambda=2\alpha$ with $b_{n}=\alpha n+\delta$ ($\delta=0$ in our case), where it is explicitly showing that $\lambda=2\alpha$, namely $\lambda=b_{n}$ with $n=2$ which will be shown in Fig. 6. The physical meaning of $\lambda$ represents the chaotic feature of dynamical system. Consequently, $b_{n}$ in our case also manifest the chaotic feature for the early universe.
Figure 6: The plot of $b_{n}=\lambda$ with $n=2$.The numerical of $b_{n}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$. $k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.
Figure 6 illustrates the evolution trend $b_{n}$ in inflation, radiation domination (RD), and matter domination (MD). Based on the discussion regarding $b_{n}$, chaos will be heightened due to the exponential expansion of the background space. As the energy decreases, $b_{n}$ will also decline since the temperature decreases. During MD, particles are generated through a preheating process known as parametric resonance, which leads to an increase in $b_{n}$. Finally, we need to analyze the dissipative coefficient $\mu_{2}$ that is related to $c_{n}$, where we could confirm the previous results via Li:2024kfm . Fig. 7 clearly shows the evolution in RD and MD. The $\mu_{2}$ in inflation is not shown in Fig. 7 since $\mu_{2}=k=1$ in light of the values of $k$ for inflation. Ref. Li:2024kfm claimed that inflation is a strong dissipative system as $\mu_{2}\geq 1$, RD and MD are weak dissipative systems as $\mu_{2}\ll 1$, which is explicitly confirmed by Fig. 7.
To summarize the chaotic features of the early universe: First, chaos is enhanced by the exponential expansion. Then, as the energy scale diminishes, chaos decreases. Finally, the generation of particles during MD through the preheating process increases chaos once again.
Figure 7: The numerical of dissipative coefficient $\mu_{2}$ in terms of $\log_{10}a$ for RD and MD, where we set $H_{0}=1$, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity. The value of $mu_{2}$ indicates the RD and MD are weak dissipative systems.
6.2 General discussion of $\phi_{n}$
Before the Krylov complexity, we discuss the general properties of $\phi_{n}$ since the definition of Krylov complexity is explicitly represented by $C_{K}=\sum_{n=1}^{+\infty}n|\phi_{n}|^{2}$. We will follow the general discussion of $\phi_{n}$ via Refs. Li:2024kfm ; Bhattacharjee:2022lzy . Eq. (49) can be rewritten by
$\partial_{\eta}\phi+n(\chi\mu\phi+2\alpha\partial_{n}\phi)=0,$ (52)
where we have defined $\tilde{c}_{n}=n\chi\mu$ and $b_{n}=n\alpha$. As $\eta$ approaches infinity, the solution of Eq. (52) has the form $\phi(n)\propto e^{-n\xi}$ with $\xi=\frac{2\alpha}{\mu\chi}$, where this form will appear in RD or MD due to their weak dissipative nature. In the weak dissipative region, the corresponding Krylov complexity behaves as follows: it increases exponentially at the onset, with the average position given by $C_{k}\propto e^{2\alpha\eta}$ before reaching $\eta_{*}$. After this point, the Krylov complexity saturates to a constant value $\xi$ when $\eta>\eta_{*}$, where $\eta_{*}\sim\frac{1}{2\alpha}\ln{\frac{2\alpha}{\chi\mu}}$. It is straightforward to demonstrate that $\exp(2a\eta_{*})=\xi$. After some simple algebra, we could obtain the formula $\xi$ in our case as follows,
$\xi=\frac{2\sqrt{(\frac{a^{2}V_{,\phi\phi}}{2k})^{2}+(\frac{a^{\prime}}{a})^{2}}}{(2+\frac{1}{n})(k+\frac{a^{2}V_{,\phi\phi}}{2k})}$ (53)
From this formula, we can observe that $\xi$ is not strictly constant; it varies with respect to $a$ and $k$. Consequently, a significant change in the background will have a substantial impact on Krylov complexity. In this section, we examined the general properties of Krylov complexity without providing the precise formula for $\phi_{n}$. To fully capture the information concerning Krylov complexity within the framework of the open system method, it is essential to have the exact wave function for $\phi_{n}$.
6.3 Wave function within the method of open system
Although we constructed the wave function for an open system in reference Li:2024kfm , based on the framework established in Bhattacharjee:2022lzy , our approach is too simplistic to capture the complete information of the open system. Notably, our wave function, as shown in Li:2024kfm , can effectively reduce to the case of a closed system, establishing a correspondence between $\eta$ and $r_{k}$. To clearly illustrate this issue, we will first present our constructed wave function as follows:
$\phi_{n}=\frac{\rm sech\eta}{1+\mu_{2}\tanh\eta}|1-\mu_{1}^{2}|^{\frac{n}{2}}(\frac{\tanh\eta}{1+\mu_{2}\tanh\eta})^{n}.$ (54)
To perform the Taylor expansion around $\mu_{2}$ to the first order, we obtain the expression
$\phi_{n}\approx|1-\mu_{1}^{2}|^{n/2}\,\text{sech}(\eta)\,\tanh^{n}(\eta)+\mathcal{O}(\mu_{2}^{n}).$
Next, we can neglect the term involving $\mu_{1}^{2}$, which enables us to find that this approximation matches the wave function given in equation (39) for $\phi_{k}=0$. Additionally, we observe the correspondence $\eta\sim r_{k}$. In principle, the wave function $\phi_{n}$ should incorporate the information of both $r_{k}$ and $\phi_{k}$ to accurately reproduce the wave function represented in (39) as the leading order of the open system’s method. To achieve this, we will follow the guidelines provided in Appendix D of Parker:2018yvk .
For deriving the two-mode squeezed state in open system, we first introduce the concept of orthogonal polynomial sequence (OPS) $\{P_{n}(x)\}_{n\geq 0}$, it satisfied with the following recurrence relation,
$P_{n+1}(x)=(x-\tilde{c}_{n})P_{n}(x)-b_{n}^{2}P_{n-1}(x),\ \ \ \ \mbox{for}\ \ n\geq 1$ (55)
where $x$ represents the Hamiltonian, $\tilde{c}_{n}$ is associated with the dissipative coefficient, and $b_{n}$ refers to the Lanczos coefficient. The sequence begins with the initial conditions $P_{0}(x)=1$ and $P_{1}(x)=x-c_{0}$. For an open system, we have $P_{n}(x;\alpha,\eta)=\det(x-\mathcal{L}_{n})$, where $\mathcal{L}_{n}$ is the Hamiltonian for the $n$-th quantum state. This determinant encodes information about $b_{n}$ as represented by the corresponding matrix. Based on the definition of $L$, we can express it in terms of $x^{n}$ or $e_{n}$, which will be used to derive the wave function using the open system method.
To obtain the similar recurrence relation as (43), we introduce the the natural orthonormal basis $\{e_{n}\}$, and meanwhile $P_{n}$ can be represented in terms of $e_{n}$
$P_{n}(x)=\Bigl{(}\prod_{k=1}^{n}b_{k}\Bigr{)}e_{n},\ \ \ \ \mbox{and}\ \ \ x^{n}=\mathcal{L}^{n}e_{0},$ (56)
where $x$ is the parameter of $P_{n}$, $\mathcal{L}$ represents our Hamiltonian expressed in terms of creation and annihilation operators, and $|e_{n}\rangle\equiv|\mathcal{O}_{n}\rangle=\Bigl{|}n_{\vec{k}};n_{-\vec{k}}\Bigr{\rangle}$. By combining Equation (55) with Equation (56), we can derive the following recurrence relation.
$b_{n+1}e_{n+1}+b_{n}e_{n-1}=(x-\tilde{c}_{n})e_{n}.$ (57)
Here, $x$ functions as the Hamiltonian. Given the recurrence of $P_{n}$, it aligns with the second kind of Meixner polynomials, denoted as $M_{n}(x;\delta,\eta_{1})$. Its generating function is
$\sum_{n=0}^{\infty}M_{n}(x;\delta,\eta_{1})\frac{t^{n}}{n!}=((1+\delta t^{2})+t^{2})^{-\eta/2}\exp(x\arctan(\frac{t}{1+\delta t}))$ (58)
The inverse polynomial of $\{M_{n}(x;\delta,\eta_{1})\}_{n\geq 0}$ is $\{Q_{n}(x;\delta,\eta_{1})\}_{n\geq 0}$, which is defined as
$Q_{n}(x)=\sum_{k=0}^{n}\mu_{n,k}x^{k},\ \ \ \ \mbox{and}\ \ \ x^{n}=\sum_{k=0}^{n}\mu_{n,k}P_{k}(x).$ (59)
where $\mu_{n,k}$ refers to a triangular linear transform with matrix elements that relate $x$ to $P_{n}$. This formula demonstrates that the polynomials $P_{n}$ can be expressed in terms of $x$. All three orthogonal polynomial sequences (OPS) mentioned are examples of Sheffer orthogonal polynomials, which are defined as a set of polynomials $\{P_{n}(x)\}_{n\geq 0}$ characterized by a generating function,
$\sum_{n=0}^{\infty}P_{n}(x)\frac{t^{n}}{n!}=f(t)\exp(xg(t)).$ (60)
And the inverse polynomials are give by
$\sum_{n=0}^{\infty}Q_{n}(x)\frac{t^{n}}{n!}=\frac{1}{f(g^{\langle-1\rangle}(t))}\exp(xg^{\langle-1\rangle}(t)),$ (61)
where $g^{\langle-1\rangle}(t)$ stands for the compositional inverse of $g(t)$, which have the following relationship
$g\bigl{(}g^{\langle-1\rangle}(t)\bigr{)}=t.$ (62)
Back to the Meixner polynomials of the second kind, we find that
$f(t)=((1+\delta t)^{2}+t^{2})^{-\eta/2},\ \ \ \ \mbox{and}\ \ \ g(t)=\arctan(\frac{t}{1+\delta t}).$ (63)
Our target function is $Q_{n}(x)$, and the rationale will be provided in the final derivations. To obtain $Q_{n}(x)$, we need to know the formula for $g^{\langle-1\rangle}$. The compositional inverse polynomial of $g(t)$ is
$t=g\bigl{(}g^{\langle-1\rangle}(t)\bigr{)}=\arctan(\frac{g^{\langle-1\rangle}(t)}{1+\delta g^{\langle-1\rangle}(t)}),$ (64)
Next,
$\frac{g^{\langle-1\rangle}(t)}{1+\delta g^{\langle-1\rangle}(t)}=\tan(t)$ (65)
And after a few simple calculations,
$g^{\langle-1\rangle}(t)=\frac{\tan(t)}{1-\delta\tan(t)},$ (66)
where $\delta$ is a parameter determined by various models. Combine Eqs. (66) and (61), we obtain the inverse polynomials of $\{M_{n}(x;\delta,\eta_{1})\}_{n\geq 0}$ as following,
$\sum_{n=0}^{\infty}Q_{n}(x)\frac{t^{n}}{n!}=\frac{\sec(t)^{\eta_{1}}}{(1-\delta\tan(t))^{\eta_{1}}}\exp(x\frac{\tan(t)}{1-\delta\tan(t)}).$ (67)
After the changing the variable as $t=i\eta$ and $\delta=i\mu_{2}$, where $\eta$ (conformal time) and $\mu_{2}$ (dissipation coefficient) are our defined parameters. Making use of $\tan i\eta=i\tanh\eta$ and $\cos i\eta=\cosh\eta$, the generating function of the inverse polynomials become
$\sum_{n=0}^{\infty}Q_{n}(x)\frac{(i\eta)^{n}}{n!}=\frac{\rm sech(\eta)^{\eta_{1}}}{(1+\mu_{2}\tanh(\eta))^{\eta_{1}}}\exp(x\frac{i\tanh(\eta)}{1+\mu_{2}\tanh(\eta)}).$ (68)
Based on Eqs. (56) and (59), we could obtain
$(e_{d}|\mathcal{L}^{n}|e_{0})=\mu_{n,d}\prod_{k=1}^{n}b_{k}.$ (69)
Furthermore, we could have
$(e_{d}|e^{i\mathcal{L}\eta}|e_{0})=\prod_{k=1}^{d}b_{k}\sum_{n=0}^{\infty}\frac{(i\eta)^{n}}{n!}\mu_{n,d}.$ (70)
In light of (59), we could obtain
$\sum_{n=0}^{\infty}\frac{(i\eta)^{n}}{n!}\mu_{n,d}=\frac{\rm sech\eta}{1+\mu_{2}\tanh\eta}\frac{1}{d!}\frac{(i\tanh\eta)^{d}}{(1+\mu_{2}\tanh\eta)^{d}},$ (71)
Then we utilize $b_{n}=\alpha n=\sqrt{|1-\mu_{1}^{2}|}n$ in our case, we naturally get the following formula,
$(e_{n}|e^{i\mathcal{L}\eta}|e_{0})=\frac{\rm sech\eta}{1+\mu_{2}\tanh\eta}|1-\mu_{1}^{2}|^{\frac{n}{2}}\frac{(i\tanh\eta)^{n}}{(1+\mu_{2}\tanh\eta)^{n}}$ (72)
thus, we obtain the wave function
$|\mathcal{O}(t))=e^{i\mathcal{L}\eta}|e_{0})=\frac{\rm sech\eta}{1+\mu_{2}\tanh\eta}\sum_{n=0}^{\infty}|1-\mu_{2}^{2}|^{\frac{n}{2}}\frac{(-1\exp(2i\phi_{k}(\eta))\tanh\eta)^{n}}{(1+\mu_{2}\tanh\eta)^{n}}|e_{n})$
where
$|\mathcal{O}_{n})=|e_{n})$
In our calculations, we multiply a phase factor $ie^{2i\phi_{k}(\eta)}$ which does not alter the wave function. Finally, we can set $r_{k}=\eta$, allowing us to denote the two-mode squeezed state within the open system as,
$\phi_{n}(\eta)=(-1)^{n}e^{2in\phi_{k}}\frac{\rm sechr_{k}}{1+\mu_{2}\tanh r_{k}}|1-\mu_{1}^{2}|^{\frac{n}{2}}\frac{(\tanh r_{k})^{n}}{(1+\mu_{2}\tanh r_{k})^{n}}.$ (73)
This wave function is our central result in this paper whose leading order of $\phi_{n}$ in terms of $\mu_{2}$ is consistent with the wave function described in (39). The idea of changing the variable is borrowed from the construction of coherent states, where a key component is the generation of the displacement operator, defined as $D(\eta)=e^{iHt}$, with $H$ being the corresponding Hamiltonian, where one can change $t$ in terms of $r_{k}$ and $\phi_{k}$. In Ref. Caputa:2021sib , it is stated that the generalized coherent state is derived by constructing a generalized displacement operator according to their group representation theory.
6.4 The power of wave function (73)
We discuss the two-mode squeezed state in an open system, as described in equation (73). Unlike the two-mode squeezed state derived from closed system methods, our wave function (73) is more realistic, as it considers a vast open environment. Additionally, although our derivation pertains to the early universe, this wave function can also be applied to fields such as quantum optics, quantum information, and condensed matter physics, where information is encoded in the parameters $\mu_{1}$ and $\mu_{2}$ determined by the specific Hamiltonian. Finally, we observe that our calculations hold true by requiring that the recurrence relation (55) which is equivalent to the Lanczos algorithm. This connection highlights the power of the Lanczos algorithm, which provides a natural method for constructing the wave function according to their Hamiltonian, even for open systems.
6.5 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$ in open system’s method
According to the standard procedure, we need to derive the evolution equations for $r_{k}$ and $\phi_{k}$ in relation to conformal time $\eta$. The detailed calculations for this derivation can be found in Appendix A. The evolution can be summarized as follows:
$\begin{split}r_{k}^{\prime}=&\ \frac{|1-\mu_{1}^{2}|^{\frac{1}{2}}\sinh 2r_{k}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}\sin 2\phi_{k}-\frac{a^{\prime}}{a}\cos 2\phi_{k}\bigr{)}-\sinh 2r_{k}\mu_{2}^{\prime}}{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})}\\ \phi_{k}^{\prime}=&\ -\bigl{(}k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr{)}+\frac{1}{2}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}\cos 2\phi_{k}+\frac{a^{\prime}}{a}\sin 2\phi_{k}\bigr{)}\\ &\ \Bigl{[}|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}+|1-\mu_{1}^{2}|^{-\frac{1}{2}}(\tanh^{-1}r_{k}+\mu_{2})\Bigr{]},\end{split}$ (74)
where prime denotes the varying with respect to the conformal time $\eta$. The same as the closed system, we change variable as $y=\log_{10}a$, the evolution equation of $r_{k}$ and $\phi_{k}$ will become as,
$\begin{split}\frac{H_{0}}{\ln 10}10^{-\frac{y}{2}(1+3\omega)}\frac{dr_{k}}{dy}=&\ \frac{|1-\mu_{1}^{2}|^{\frac{1}{2}}\sinh 2r_{k}\bigl{(}\frac{10^{2y}V_{,\phi\phi}}{2k}\sin 2\phi_{k}-H_{0}10^{-\frac{y}{2}(1+3\omega)}\cos 2\phi_{k}\bigr{)}-\sinh 2r_{k}\mu_{2}^{\prime}}{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})}\\ \frac{H_{0}}{\ln 10}10^{-\frac{y}{2}(1+3\omega)}\frac{d\phi_{k}}{dy}=&\ -\bigl{(}k+\frac{10^{2y}V_{,\phi\phi}}{2k}\bigr{)}+\frac{1}{2}\bigl{(}\frac{10^{2y}V_{,\phi\phi}}{2k}\cos 2\phi_{k}+H_{0}10^{-\frac{y}{2}(1+3\omega)}\sin 2\phi_{k}\bigr{)}\\ &\ \Bigl{[}|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}+|1-\mu_{1}^{2}|^{-\frac{1}{2}}(\tanh^{-1}r_{k}+\mu_{2})\Bigr{]}\end{split}$ (75)
where
$\begin{split}&\ \mu_{2}=k+\frac{10^{2y}V_{,\phi\phi}}{2k},\ \ \ \ \mu_{2}^{\prime}=H_{0}10^{-\frac{y}{2}(1+3\omega)}\frac{10^{2y}V_{,\phi\phi}}{k}+\frac{10^{2y}V_{,\phi\phi}^{\prime}}{2k}\\ &\ |1-\mu_{1}^{2}|=(\frac{10^{2y}V_{,\phi\phi}}{2k})^{2}+(H_{0}10^{-\frac{y}{2}(1+3\omega)})^{2}.\end{split}$
During the calculations, we accounted for the background of $\phi$ using conformal time $\eta$, as demonstrated by the various potentials in Eq. (23). Consequently, $\mu_{2}$ and $\mu_{1}$ also vary with respect to conformal time, since they include contributions from the potential and the scale factor $a$.
Figure 8: The numerical of $r_{k}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$. $k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.
Fig. 8 clearly illustrates the evolution of $r_{k}$ across various epochs, including inflation, radiation domination (RD), and matter domination (MD). It is evident that the different potentials produce similar trends for $r_{k}$. Unlike in Fig. 3, there is no oscillation of $r_{k}$ once dissipation effects are considered. Additionally, the trend of $r_{k}$ varies, particularly during the MD phase, where the evolution consistently shows an increase. These differences will result in distinctive evolutions of Krylov complexity and Krylov entropy.
6.6 Krylov complexity and Krylov entropy with the approach of open system
In this section, we will continue use the wave function (73) to explore the Krylov complexity and Krylov entropy. All of these calculation are exactly the same with our previous study Li:2024kfm , the only differences is that we need to use $r_{k}$ to take place of $\eta$ in the formulas for Krylov complexity and Krylov entropy whose formulas can be expressed as following:
$K=\frac{\rm sech^{2}r_{k}|1-\mu_{1}^{2}|\tanh^{2}r_{k}}{[1+2\mu_{2}\tanh r_{k}+(\mu_{2}^{2}-|1-\mu_{1}^{2}|)\tanh^{2}r_{k}]^{2}}$ (76)
and
$\begin{split}S_{k}=&\ \frac{\rm sech^{2}r_{k}}{A^{2}}\bigl{[}(1+\mu_{2}\tanh r_{k})^{2}\ln(1+\mu_{2}\tanh r_{k})^{2}-\tanh^{2}r_{k}|1-\mu_{1}^{2}|\ln\tanh^{2}r_{k}\\ &\ -[(1+\mu_{2}\tanh r_{k})^{2}-|1-\mu_{1}^{2}|\tanh^{2}r_{k}]\ln\rm sech^{2}r_{k}-\tanh^{2}r_{k}|1-\mu_{1}^{2}|\ln|1-\mu_{1}^{2}|\bigr{]}\end{split}$ (77)
where we have defined $A=1+2\mu_{2}\tanh r_{k}+(\mu_{2}^{2}-|1-\mu_{1}^{2}|)\tanh^{2}r_{k}$ for simplicity. It straightforwardly demonstrates the leading order of Krylov complexity (76) and Krylov entropy (77) will nicely recover into the case of Eq. (40) and Eq. (41) under the weak dissipative approximation, namely, $\mu_{2}\ll 1$.
Fig. 9 illustrates the evolution of Krylov complexity during inflation, radiation domination (RD), and matter domination (MD). During inflation, we observe that Krylov complexity increases exponentially, which is consistent with the closed system’s method. However, the trends in RD and MD differ from those shown in Fig. 4, where the complexity tends to stabilize around certain constant values, oscillating around these values in both RD and MD. In the case of an open system, it is evident that Krylov complexity decreases during RD. For MD, Krylov complexity initially peaks as the scale factor $a$ increases and then declines to stabilize at constant values. Additionally, it is noteworthy that the order of Krylov complexity in an open system is significantly lower than in a closed system, even in RD and MD. Our numerical analysis shows that the term $\rm sech(r_{k})$ in Eq. (76) leads to this outcome, regardless of whether the system is weakly dissipative or not. We have demonstrated that both RD and MD can be approximated as weakly dissipative systems under the condition $\mu<1$. Recalling the meaning of Krylov complexity, it describes the operator growth in a many-body system and our early universe is a perfect chaotic system due to $b_{n}=\alpha n$. Our numeric reveals that the operator growth is much smaller compared with the case of a closed system, which is consistent with reality since the dissipation will lead to faster decreasing of operator growth compared with that without the dissipation effects.
Once one understands the role of dissipation effects, it becomes evident that the order of Krylov entropy in an open system is lower compared to that in a closed system, as illustrated in Figs 10 and 5. The evolution trend is particularly distinctive for RD and MD, as shown in Figure 10, where the Krylov entropy initially peaks before stabilizing at constant values. In the case of MD, the Krylov entropy initially holds steady at certain values before decreasing to some constant level. Overall, the background conditions and dissipative effects have a significant impact on both Krylov complexity and Krylov entropy.
In this section, we utilized the wave function presented in Eq. (73) to investigate Krylov complexity and Krylov entropy. By comparing our findings to those from a closed system, we observed that the dissipative effects not only influence the overall evolution of the system but also reduce the order of Krylov complexity and Krylov entropy. Since Krylov complexity reflects the growth of operators—a fundamental quantum property of a dynamical system—we can conclude that, when the growth of the operator is diminished in the open system compared to the closed system, the dissipative effects lead to a more rapid decoherence-like behavior. Meanwhile, we also could see that the various potentials will almost produce the same trend of Krylov complexity and Krylov entropy in different epochs.
Figure 9: The numerical of Krylov complexity in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=0.01$ at period inflation and RD, $k=0.005$ at period MD for simplicity.
Figure 10: The numerical of Krylov entropy in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=0.01$ at period inflation and RD, $k=0.005$ at period MD for simplicity.
7 Summary and outlook
The definition of Krylov complexity is clear and unambiguous, especially when compared to circuit complexity. Its uniqueness eliminates the need for introducing a parametric manifold to illustrate the concept of complexity. Additionally, Krylov complexity characterizes the growth of operations within a dynamical system, and the order of Krylov complexity can indicate the quantum level of that system. Krylov complexity is rooted in the Lanczos algorithm, a powerful method for assessing the nature of dynamical systems, which can be chaotic, integrable, or free. The key criteria for this assessment is the Lanczos coefficient, whose formula indicates the characteristics of a given dynamical system. For example, the relation $b=\alpha n$ suggests that the system is chaotic. In our paper, we highlight that the Lanczos coefficient effectively reveals that our early universe was a many-body, infinite, and chaotic system. From a cosmological perspective, the universe is considered an entirely open system. Notably, the Lanczos algorithm offers a natural way to construct the wave function of an open system using the second kind of Meixner polynomials. Given these insights, it is particularly important to explore Krylov complexity throughout the early universe, including inflation, RD and MD.
In this paper, we thoroughly examine Krylov complexity in both open and closed system methods. Our primary result is the derivation of the two-mode squeezed state using the open system approach, along with the calculation of its evolution equations for $\phi_{k}(\eta)$ and $r_{k}(\eta)$. Below, we present our key findings.
$(a)$. In Sec. 2, we introduce the Lanczos algorithm and the associated Lanczos coefficients, which apply equally to both open and closed systems. As noted by Parker:2018yvk , the Liouvillian superoperator acts as the Hamiltonian when it can be expressed in terms of creation and annihilation operators. Our exploration of Krylov complexity encompasses the entire universe, including inflation, RD, MD, and the violations of slow-roll conditions, which necessitates considering contributions from potential energy. Sec. 3 discusses the most common potentials, such as the Higgs potential, Starobinsky’s inflationary potential, and chaotic inflationary potential. These models encompass a wide range of inflationary scenarios due to the unifying characteristics of $\alpha$ attractors. To accurately capture the information pertaining to these potentials, we examine the dynamics of the field $\phi$ in relation to the scale factor during the preheating process, as represented in Eq. (21). From these calculations, we determine the effective mass for various inflationary potentials, as illustrated in Fig. 2. Our findings indicate that the potential approaches certain constant values during RD and MD.
$(b)$. In light of the scale factor and various potentials, we examine Krylov complexity using the closed system method. First, we numerically simulate the $r_{k}$ of the two-mode squeezed state, as depicted in Fig. 3. The simulation shows oscillations during inflation, which are enhanced in RD and MD, maintaining some constant values. Notably, the different potentials produce almost identical trends for $r_{k}$. Based on the numerical results for $r_{k}$, we also simulate the corresponding Krylov complexity and Krylov entropy, as illustrated in Figs. 4 and 5. The trends of both Krylov complexity and Krylov entropy are similar across the corresponding epochs. During inflation, we observe exponential growth, consistent with our previous study Li:2024kfm . In the RD and MD phases, the trends show initial enhancement before saturating to constant values. Additionally, various inflationary models exhibit nearly the same trends for Krylov complexity and Krylov entropy, confirming the approximation that the effective mass of the inflaton is approximately constant, as indicated in Li:2024ljz .
$(c)$. Within the method of an open system, we rigorously derive the wave function of an open system using the second kind of Meixner polynomials, which we name it as the open two-mode squeezed state. This derivation is represented in Eq. (73). Throughout this process, we discover that the Lanczos coefficient $b_{n}$ and the dissipative coefficient are precisely determined by the specific Hamiltonian, leading to the conclusion that $b_{n}$ is identical across both methods. Our findings indicate that $b_{n}=\alpha n$, suggesting that our universe behaves as an entirely chaotic system. Consequently, the Lyapunov index is equivalent to $b_{n}$ when $n=2$. Figure 6 illustrates the chaos present across various epochs. During the inflationary period, chaos increases exponentially before diminishing to a certain level, only to rise again due to the preheating process. Additionally, we examine the evolution of the dissipative coefficient in RD and MD. This analysis builds on previous work in Ref. Li:2024kfm , which discusses the strong dissipative nature of the system during inflation. From Fig. 7, we can confirm that the RD and MD systems are weakly dissipative, characterized by $\mu_{2}\ll 1$.
Additionally, we derive the evolution of $r_{k}$ and $\phi_{k}$, noting that the solution for $r_{k}$ is dependent on the solution for $\phi_{k}$ due to their coupling, as described by Eq. (74). Using the wave function in Eq. (73), we calculate the Krylov complexity and Krylov entropy, which are represented in Eqs. (76) and (77). The leading order of these quantities aligns well with the formulas derived from the closed system method. In accordance with standard procedures, we need the numerical values of $r_{k}$ within the open system framework, as illustrated in Fig. 8. Unlike the closed system, there is no oscillation observed for $r_{k}$ during inflation, and the trends of Krylov complexity and Krylov entropy are depicted in Figs. 9 and 10. Compared to the closed system method, we observe that the values of Krylov complexity and Krylov entropy are significantly smaller. Recalling the definition of Krylov complexity, which pertains to the operator growth in a dynamical system, we find that a smaller operator growth suggests a decoherence-like behavior within the Krylov complexity. When dissipative effects are considered, we observe decoherence-like behavior, whether the system is weakly dissipative or not.
While we have achieved a comprehensive understanding of Krylov complexity in the early universe, many ideas still remain to be explored. Here, we outline several future directions.
Firstly, we investigate the Krylov complexity in the context of single-field inflation, and it is quite natural to extend our work into the multi-field framework Liu:2020zzv ; Liu:2019xhn ; Liu:2021rgq ; Zhang:2022bde as well as $f(R)$ gravity Liu:2018hno ; Liu:2020zlr ; Liu:2018htf . This extension allows us to explore the geometrical effects of fields and higher orders of Ricci curvature on Krylov complexity. Our Hamiltonian, given by (33), is Hermitian and includes the component of an open system. The traditional approach to handling open systems in quantum mechanics involves constructing what is known as a Lindbladian Lindblad:1975ef ; Gorini:1975nb , where the corresponding Hamiltonian is non-Hermitian. In our case, a natural method is to expand the action to higher orders with respect to the Mukhanov-Sasaki variable.
Secondly, our results indicate that the Krylov complexity in a weak dissipative system (RD or MD) will not follow the general discussions mentioned in Sec. 6.2. One possible explanation for this discrepancy is the neglect of decoherence in curvature perturbation during horizon exits. Ref. Bhattacharjee:2022lzy suggests that decoherence in circuit complexity may lead to this phenomenon of saturation. A similar concept might apply to Krylov complexity, particularly concerning the role of decoherence in the construction of the wave function.
Thirdly, we have already obtained the open two-mode squeezed state as described in (73). In light this wave function, we can explore several additional avenues. For instance, we can calculate the two-point correlation functions to derive the power spectrum. Since the derivation of this wave function is model-independent, we have the flexibility to utilize different models of this wave function to test its validity. Conversely, we can also constrain the parameters $\mu_{1}$ and $\mu_{2}$ based on data from the Cosmic Microwave Background (CMB). Furthermore, it is important to highlight that Ref. Adhikari:2022oxr indicates the potential application of the CV conjecture to Krylov complexity during inflation; however, their approach is tailored for a closed system. In contrast, our universe is fundamentally an open system, which provides us with the opportunity to further investigate the validity of the CV conjecture for Krylov complexity within the context of open systems.
Acknowledgements.
LH and KH are funded by NSFC grant NO. 12165009, Hunan Natural Science Foundation NO. 2023JJ30487 and NO. 2022JJ40340.
Appendix A The calculation of $r_{k}$ and $\phi_{k}$ within the method of open system
In this appendix, we will give the detailed calculations for $r_{k}$ and $\phi_{k}$ based on the open two-mode squeezed state,
$|\psi\rangle=\frac{\rm{sech}\ r_{k}}{1+\mu_{2}\tanh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}$ (78)
We will write down our Hamiltonian (33) as follows,
$\hat{H}=\bigl{(}k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr{)}(\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}+\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}})+\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a}\bigr{)}\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{\vec{k}}^{\dagger}+\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr{)}\hat{c}_{-\vec{k}}\hat{c}_{\vec{k}}$ (79)
[…]
$\begin{split}\hat{H}|\psi\rangle=&\ \frac{\rm sech~{}r_{k}}{1+\mu_{2}\tanh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}\Bigl{[}\bigl{(}k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr{)}(\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}\\ &\ +\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}})|n,n\rangle_{\vec{k},-\vec{k}}+\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a}\bigr{)}\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{\vec{k}}^{\dagger}|n,n\rangle_{\vec{k},-\vec{k}}+\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr{)}\hat{c}_{-\vec{k}}\hat{c}_{\vec{k}}|n,n\rangle_{\vec{k},-\vec{k}}\Bigr{]}\\ =&\ \frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}\Bigl{[}\bigl{(}k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr{)}(2n+1)\\ &\ |n,n\rangle_{\vec{k},-\vec{k}}+\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a}\bigr{)}(n+1)|n+1,n+1\rangle_{\vec{k},-\vec{k}}+\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr{)}n|n-1,n-1\rangle_{\vec{k},-\vec{k}}\Bigr{]}\\ =&\ \frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl{[}\bigl{(}k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr{)}-|1-\mu_{1}^{2}|^{\frac{1}{2}}e^{2i\phi_{k}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr{)}\Bigr{]}|0,0\rangle_{\vec{k},-\vec{k}}\\ &\ +\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl{[}\bigl{(}k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr{)}-|1-\mu_{1}^{2}|^{\frac{1}{2}}e^{2i\phi_{k}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr{)}\Bigr{]}\\ &\ \sum_{n=1}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}+\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl{[}\bigl{(}2k+\frac{a^{2}V_{,\phi\phi}}{k}\bigr{)}\\ &\ -|1-\mu_{1}^{2}|^{\frac{1}{2}}e^{2i\phi_{k}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr{)}-|1-\mu_{1}^{2}|^{-\frac{1}{2}}e^{-2i\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{-1}\\ &\ \bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a}\bigr{)}\Bigr{]}\sum_{n=1}^{\infty}(-1)^{n}n|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}\end{split}$ (80)
Then, the right hand side is following as,
$\begin{split}i\frac{d}{d\eta}|\psi\rangle=&\ i\frac{d}{d\eta}\Bigl{[}\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}\Bigr{]}\\ =&\ i\Bigl{[}-\frac{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})r_{k}^{\prime}+\sinh 2r_{k}\mu_{2}^{\prime}}{2\cosh^{3}r_{k}(1+\mu_{2}\tanh r_{k})^{2}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}\\ &\ |n,n\rangle_{\vec{k},-\vec{k}}\Bigr{]}+i\Bigl{[}\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}\\ &\ \bigl{[}2i\phi_{k}^{\prime}+\frac{2r_{k}^{\prime}}{\sin 2r_{k}(1+\mu_{2}\tanh r_{k})}+\frac{|1-\mu_{1}^{2}|^{\prime}}{2|1-\mu_{1}^{2}|}-\frac{\tanh r_{k}\mu_{2}^{\prime}}{1-\mu_{2}\tanh r_{k}}\bigr{]}|n,n\rangle_{\vec{k},-\vec{k}}\Bigr{]}\\ =&\ -i\frac{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})r_{k}^{\prime}+\sinh 2r_{k}\mu_{2}^{\prime}}{2\cosh^{3}r_{k}(1+\mu_{2}\tanh r_{k})^{2}}|0,0\rangle_{\vec{k},-\vec{k}}\\ &\ -i\frac{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})r_{k}^{\prime}+\sinh 2r_{k}\mu_{2}^{\prime}}{2\cosh^{3}r_{k}(1+\mu_{2}\tanh r_{k})^{2}}\sum_{n=1}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}\\ &\ |n,n\rangle_{\vec{k},-\vec{k}}+\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl{(}-2\phi_{k}^{\prime}+i\frac{2r_{k}^{\prime}}{\sin 2r_{k}(1+\mu_{2}\tanh r_{k})}+i\frac{|1-\mu_{1}^{2}|^{\prime}}{2|1-\mu_{1}^{2}|}\\ &\ -i\frac{\tanh r_{k}\mu_{2}^{\prime}}{1-\mu_{2}\tanh r_{k}}\Bigr{)}\sum_{n=1}^{\infty}(-1)^{n}n|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}\end{split}$ (81)
[…]
$\begin{split}&\ A_{1}|0,0\rangle_{\vec{k},-\vec{k}}+A_{1}\sum_{n=1}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}\\ &\ +A_{2}\sum_{n=1}^{\infty}(-1)^{n}n|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}\\ =&\ B_{1}|0,0\rangle_{\vec{k},-\vec{k}}+B_{1}\sum_{n=1}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}\\ &\ +B_{2}\sum_{n=1}^{\infty}(-1)^{n}n|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{n}|n,n\rangle_{\vec{k},-\vec{k}}.\end{split}$ (82)
Thus, we obtain $A_{1}=B_{1}$ as
$\begin{split}&\ -i\frac{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})r_{k}^{\prime}+\sinh 2r_{k}\mu_{2}^{\prime}}{2\cosh^{3}r_{k}(1+\mu_{2}\tanh r_{k})^{2}}\\ =\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}&\ \Bigl{[}\bigl{(}k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr{)}-|1-\mu_{1}^{2}|^{\frac{1}{2}}e^{2i\phi_{k}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr{)}\Bigr{]},\end{split}$ (83)
and we’re concerned about the imaginary component,
$\begin{split}&\ -i\frac{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})r_{k}^{\prime}+\sinh 2r_{k}\mu_{2}^{\prime}}{2\cosh^{3}r_{k}(1+\mu_{2}\tanh r_{k})^{2}}\\ =-\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}&\ \Bigl{[}|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}\sin 2\phi_{k}-\frac{a^{\prime}}{a}\cos 2\phi_{k}\bigr{)}\Bigr{]}.\end{split}$ (84)
and therefore, we could obtain the evolution of $r_{k}$ as,
$r_{k}^{\prime}=\frac{|1-\mu_{1}^{2}|^{\frac{1}{2}}\sinh 2r_{k}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}\sin 2\phi_{k}-\frac{a^{\prime}}{a}\cos 2\phi_{k}\bigr{)}-\sinh 2r_{k}\mu_{2}^{\prime}}{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})}.$ (85)
The other side, we have $A_{2}=B_{2}$ showing as
$\begin{split}&\ \frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl{[}-2\phi_{k}^{\prime}+i\frac{2r_{k}^{\prime}}{\sin 2r_{k}(1+\mu_{2}\tanh r_{k})}+i\frac{|1-\mu_{1}^{2}|^{\prime}}{2|1-\mu_{1}^{2}|}-i\frac{\tanh r_{k}\mu_{2}^{\prime}}{1-\mu_{2}\tanh r_{k}}\Bigr{]}\\ =&\ \frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl{[}\bigl{(}2k+\frac{a^{2}V_{,\phi\phi}}{k}\bigr{)}-|1-\mu_{1}^{2}|^{\frac{1}{2}}e^{2i\phi_{k}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr{)}\\ &\ -|1-\mu_{1}^{2}|^{-\frac{1}{2}}e^{-2i\phi_{k}}\Bigl{(}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr{)}^{-1}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a}\bigr{)}\Bigr{]},\end{split}$ (86)
the real component can read as,
$\begin{split}-\frac{2\phi_{k}^{\prime}\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}&\ =\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl{[}\bigl{(}2k+\frac{a^{2}V_{,\phi\phi}}{k}\bigr{)}\\ -\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}\cos 2\phi_{k}+\frac{a^{\prime}}{a}\sin 2\phi_{k}\bigr{)}&\ \Bigl{(}|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}+|1-\mu_{1}^{2}|^{-\frac{1}{2}}(\tanh^{-1}r_{k}+\mu_{2})\Bigr{)}\Bigr{]}.\end{split}$ (87)
Finally, we could obtain the evolution equation of $\phi_{k}$ as follows,
$\begin{split}\phi_{k}^{\prime}=&\ -\bigl{(}k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr{)}+\frac{1}{2}\bigl{(}\frac{a^{2}V_{,\phi\phi}}{2k}\cos 2\phi_{k}+\frac{a^{\prime}}{a}\sin 2\phi_{k}\bigr{)}\\ &\ \Bigl{[}|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}+|1-\mu_{1}^{2}|^{-\frac{1}{2}}(\tanh^{-1}r_{k}+\mu_{2})\Bigr{]}.\end{split}$ (88)
One can easily check that Eqs. (85) and (88) could nicely recover into Eq. (35) as $\mu_{1}=\mu_{2}=0$.

### v2

@@
 Notably, we find that Krylov complexity differs between the closed and open system approaches.
-To effectively capture the impact of potentials during the RD and MD phases, we analyze various inflationary potentials—including the Higgs potential, the $R^{2}$ inflationary potential, and chaotic inflationary potential—taking into account the violations of slow-roll conditions.
+To effectively capture the impact of potentials during the RD and MD phases, we analyze various inflationary potentials, including the Higgs potential, the $R^{2}$ inflationary potential, and chaotic inflationary potential, which is taking into account the violations of slow-roll conditions.
 This analysis is conducted in terms of conformal time through the preheating process.
-Our numerical results indicate that the evolution of Krylov complexity and Krylov entropy is remarkably similar across both methods, regardless of the potential under consideration.
+Our numerical results indicate that the evolution of Krylov complexity and Krylov entropy is remarkably similar within distinctive potentials in RD and MD.
 Additionally, we rigorously construct what is referred to as an open two-mode squeezed state, utilizing the second kind of Meixner polynomials.
@@
 This research provides new insights into exploring the universe from the perspective of quantum information.
-arxiv: 1234.56789
 1 Introduction
@@
 Due to the importance of the SKY model in high energy physics, the Krylov complexity can also be implemented into this field Rabinovici:2020ryf ; Jian:2020qpp ; He:2022ryk .
-Meanwhile, the Krylov complexity can be applied to the generalized coherent state Patramanis:2021lkx , Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , conformal field theory Dymarsky:2021bjq ; Caputa:2021ori , topological phases of matter Caputa:2022eye and even extend into the quantum field theory He:2024hkw ; He:2024xjp .
+Meanwhile, the Krylov complexity can be applied to the generalized coherent state Patramanis:2021lkx , Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , conformal field theory Dymarsky:2021bjq ; Caputa:2021ori ; Kundu:2023hbk , topological phases of matter Caputa:2022eye and even extend into the quantum field theory He:2024hkw ; He:2024xjp .
 Ref.
 Dymarsky:2019elm proved that the chaos can be understood as a delocalization in Krylov complexity.
-Furthermore, one could observe the exponential growth of Krylov complexity n integrable systems with saddle-dominated scrambling, whose universal features can be verified by Huh:2023jxt .
+Furthermore, one could observe the exponential growth of Krylov complexity in integrable systems with saddle-dominated scrambling, whose universal features can be verified by Huh:2023jxt .
 More recent developments of Krylov complexity can be found in Refs.
-Erdmenger:2023wjg ; Hashimoto:2023swv ; Vasli:2023syq ; Gill:2023umm ; Bhattacharjee:2023uwx ; Adhikari:2022whf ; Loc:2024oen ; Caputa:2024vrn ; Basu:2024tgg ; Sasaki:2024puk ; Caputa:2024xkp ; Bhattacharjee:2022qjw ; Sahu:2024opm ; Kim:2021okd ; Chen:2024imd ; Bhattacharjee:2024yxj ; Sanchez-Garrido:2024pcy ; Chattopadhyay:2024pdj .
+Erdmenger:2023wjg ; Hashimoto:2023swv ; Vasli:2023syq ; Gill:2023umm ; Bhattacharjee:2023uwx ; Adhikari:2022whf ; Loc:2024oen ; Caputa:2024vrn ; Basu:2024tgg ; Sasaki:2024puk ; Caputa:2024xkp ; Bhattacharjee:2022qjw ; Sahu:2024opm ; Kim:2021okd ; Chen:2024imd ; Bhattacharjee:2024yxj ; Sanchez-Garrido:2024pcy ; Chattopadhyay:2024pdj ; Balasubramanian:2024ghv ; Bhattacharya:2024hto ; Mohan:2023btr .
 Especailly, the concept of Krylov complexity can be extended into in various fields Nandy:2024htc .
@@
 […]
-$\partial_{t}\phi_{n}(t)=b_{n}\phi_{n-1}-b_{n-1}\phi_{n+1}$ (9)
+$\partial_{t}\phi_{n}(t)=b_{n}\phi_{n-1}-b_{n+1}\phi_{n+1}$ (9)
 Next, we can define the Krylov complexity using the wave function $\phi_{n}$ as follows
@@
 For the potential of the Higgs field, it is different which shows that there is a peak in RD and after that, it will decrease into some constants.
-But, it indicates that the effective mass of $V_{2}$ and $V_{2}$ are of almost the same order based on the COBE normalization Planck:2015sxf .
+But, it indicates that the effective mass of $V_{2}$ and $V_{3}$ are of almost the same order based on the COBE normalization Planck:2015sxf .
 In this section, we provide an introduction to inflation, RD, and MD, which are characterized by the corresponding scale factor.
@@
 where we have utilized Eq. (17) and $\frac{a^{\prime}}{a}=H_{0}a^{-\frac{1}{2}(1+3\omega)}$.
-Figure 3: The numerical solution of $R_{k}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$.
-$k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity..
+Figure 3: The numerical solution of $r_{k}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD, $k=0.005$ at MD for simplicity.
 The effect of potential is always ignored in slow-roll inflation.
-But in RD and MD periods, the influence can be seen by using different potential energies as shown in these panels.
+But in RD and MD periods, the impact of potential is taken into account.
 Before analyzing the figure of $r_{k}$, it’s important to note that the Krylov complexity depends only on $r_{k}$.
@@
 There is an oscillation before the horizon exits, and then $r_{k}$ will grow linearly, consistent with the standard case of Li:2024kfm .
-As we transition to the radiation-dominated era (RD), we observe that $r_{k}$ will increase to some constant value that oscillates around this value.
+As we transit to RD, we observe that $r_{k}$ will increase to some constant value that oscillates around this value.
 Interestingly, different potentials produce almost the same trend for $r_{k}$, indicating that the specific potential does not significantly impact the evolution of $r_{k}$.
-This observation also applies to the matter-dominated era (MD), where the evolution trend is similar to RD.
+This observation also applies to MD, where the evolution trend is similar to RD.
 However, we notice that varying momentum significantly affects $r_{k}$, leading to oscillations as the value of $k$ decreases.
@@
 Our research indicates that the value of $r_{k}$ increases during inflation.
-The behavior of $r_{k}$ in radiation-dominated (RD) and matter-dominated (MD) eras is similar and will oscillate around constant values.
+The behavior of $r_{k}$ in RD and MD eras is similar and will oscillate around constant values.
 Our numerical analysis also shows that various inflationary models produce a similar trend for $r_{k}$.
@@
 It is explicitly showing that the Krylov complexity is independent of $\phi_{k}$, that is why we only list the figure of $r_{k}$.
-Figure 4: The numerical of Krylov complexity in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$.
-$k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.
+Figure 4: The numerical of Krylov complexity in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.
 Being armed with Eq.
@@
 where $\phi_{n}$ is the wave function (27). Once given the analytic formula of K-entropy, we could numerically simulate it in various epochs as shown in Fig. 5.
-Figure 5: The numerical of Krylov entropy in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$.
-$k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.
+Figure 5: The numerical of Krylov entropy in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.
 Fig.
@@
 In our case, we could see that $b_{n}\propto n$, thus the whole early universe is an ideal chaotic system.
-6.1.1 $b_{n}$, $c_{n}$ and $\mu_{2}$
+6.1.1 $b_{n}$ and $\mu_{2}$
 […]
@@
 Consequently, $b_{n}$ in our case also manifest the chaotic feature for the early universe.
-Figure 6: The plot of $b_{n}=\lambda$ with $n=2$.The numerical of $b_{n}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$.
-$k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.
+Figure 6: The plot of $b_{n}=\lambda$ with $n=2$.The numerical of $b_{n}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.
 Figure 6 illustrates the evolution trend $b_{n}$ in inflation, radiation domination (RD), and matter domination (MD).
@@
 Finally, the generation of particles during MD through the preheating process increases chaos once again.
-Figure 7: The numerical of dissipative coefficient $\mu_{2}$ in terms of $\log_{10}a$ for RD and MD, where we set $H_{0}=1$, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.
-The value of $mu_{2}$ indicates the RD and MD are weak dissipative systems.
+Figure 7: The numerical of dissipative coefficient $\mu_{2}$ in terms of $\log_{10}a$ for RD and MD, where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.
+The value of $\mu_{2}$ indicates the RD and MD are weak dissipative systems.
 6.2 General discussion of $\phi_{n}$
@@
 In the weak dissipative region, the corresponding Krylov complexity behaves as follows: it increases exponentially at the onset, with the average position given by $C_{k}\propto e^{2\alpha\eta}$ before reaching $\eta_{*}$.
-After this point, the Krylov complexity saturates to a constant value $\xi$ when $\eta>\eta_{*}$, where $\eta_{*}\sim\frac{1}{2\alpha}\ln{\frac{2\alpha}{\chi\mu}}$.
+After this point, the Krylov complexity saturates to a constant value $\xi$ when $\eta>\eta_{*}$, where $\eta_{*}\sim\frac{1}{2\alpha}\ln{\frac{2\alpha}{\chi\mu}}$ dubbed as the scrambling time.
+It is evident that $\eta_{*}$ determines the range of exponential growth of Krylov complexity, determined by the dynamical system itself.
 It is straightforward to demonstrate that $\exp(2a\eta_{*})=\xi$.
@@
 Here, $x$ functions as the Hamiltonian. Given the recurrence of $P_{n}$, it aligns with the second kind of Meixner polynomials, denoted as $M_{n}(x;\delta,\eta_{1})$. Its generating function is
-$\sum_{n=0}^{\infty}M_{n}(x;\delta,\eta_{1})\frac{t^{n}}{n!}=((1+\delta t^{2})+t^{2})^{-\eta/2}\exp(x\arctan(\frac{t}{1+\delta t}))$ (58)
+$\sum_{n=0}^{\infty}M_{n}(x;\delta,\eta_{1})\frac{t^{n}}{n!}=((1+\delta t)^{2}+t^{2})^{-\eta/2}\exp(x\arctan(\frac{t}{1+\delta t}))$ (58)
 The inverse polynomial of $\{M_{n}(x;\delta,\eta_{1})\}_{n\geq 0}$ is $\{Q_{n}(x;\delta,\eta_{1})\}_{n\geq 0}$, which is defined as
@@
 thus, we obtain the wave function
-$|\mathcal{O}(t))=e^{i\mathcal{L}\eta}|e_{0})=\frac{\rm sech\eta}{1+\mu_{2}\tanh\eta}\sum_{n=0}^{\infty}|1-\mu_{2}^{2}|^{\frac{n}{2}}\frac{(-1\exp(2i\phi_{k}(\eta))\tanh\eta)^{n}}{(1+\mu_{2}\tanh\eta)^{n}}|e_{n})$
+$|\mathcal{O}(t))=e^{i\mathcal{L}\eta}|e_{0})=\frac{\rm sech\eta}{1+\mu_{2}\tanh\eta}\sum_{n=0}^{\infty}|1-\mu_{1}^{2}|^{\frac{n}{2}}\frac{(-1\exp(2i\phi_{k}(\eta))\tanh\eta)^{n}}{(1+\mu_{2}\tanh\eta)^{n}}|e_{n})$
 where
@@
 Consequently, $\mu_{2}$ and $\mu_{1}$ also vary with respect to conformal time, since they include contributions from the potential and the scale factor $a$.
-Figure 8: The numerical of $r_{k}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$.
-$k=1$ at period inflation, $k=0.01$ at period RD, $k=0.005$ at period MD for simplicity.
+Figure 8: The numerical of $r_{k}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.
 Fig.
@@
 Meanwhile, we also could see that the various potentials will almost produce the same trend of Krylov complexity and Krylov entropy in different epochs.
-Figure 9: The numerical of Krylov complexity in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=0.01$ at period inflation and RD, $k=0.005$ at period MD for simplicity.
-Figure 10: The numerical of Krylov entropy in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=0.01$ at period inflation and RD, $k=0.005$ at period MD for simplicity.
+Figure 9: The numerical of Krylov complexity in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.
+Figure 10: The numerical of Krylov entropy in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.
 7 Summary and outlook

### v3

@@
 Abstract
-The Lanczos algorithm offers a method for constructing wave functions for both closed and open systems based on their Hamiltonians.
-Given that the entire early universe is fundamentally an open system, we apply the Lanczos algorithm to investigate Krylov complexity across different phases of the early universe, including inflation, the radiation-dominated period (RD), and the matter-dominated period (MD).
-Notably, we find that Krylov complexity differs between the closed and open system approaches.
-To effectively capture the impact of potentials during the RD and MD phases, we analyze various inflationary potentials, including the Higgs potential, the $R^{2}$ inflationary potential, and chaotic inflationary potential, which is taking into account the violations of slow-roll conditions.
-This analysis is conducted in terms of conformal time through the preheating process.
-Our numerical results indicate that the evolution of Krylov complexity and Krylov entropy is remarkably similar within distinctive potentials in RD and MD.
-Additionally, we rigorously construct what is referred to as an open two-mode squeezed state, utilizing the second kind of Meixner polynomials.
-Based on this construction, we are the first to calculate the evolution equations for $r_{k}$ and $\phi_{k}$ as they relate to the scale factor.
-Our findings suggest that dissipative effects lead to a rapid decoherence-like behavior.
-Moreover, our results indicate that inflation behaves as a strongly dissipative system, while both the radiation-dominated and matter-dominated phases exhibit characteristics of weak dissipation.
-This research provides new insights into exploring the universe from the perspective of quantum information.
-1 Introduction
-The idea of complexity was initially introduced in the holographic framework Stanford:2014jda .
-In this framework, it was observed that the conformal field theory (CFT) reaches thermal equilibrium relatively quickly, while the corresponding bulk gravity (wormhole) takes a much longer time to reach the same state.
-To address this issue, Refs.
-Hartman:2013qma ; Liu:2013iza introduced the concept of computational complexity within the holographic framework.
-This concept suggests that while the emergence of spacetime is related to quantum entanglement, the evolution of space occurs through computational complexity.
-The concept of complexity has become increasingly essential in high-energy physics even without the holographic principle.
-First, the so-called computational complexity describes the minimum number of logical steps required for a task Aaronson:2016vto .
-Furthermore, the method of calculating the complexity was first proposed by Nielson, $\it e.t.c$, dubbed as the geometrical method Nielsen:2005mkt ; Dowling:2006tnk ; Nielsen:2006cea .
-Another approach is highly related to the ”Fubini study” distance of the geometry of information Chapman:2017rqy .
-These two geometrical methods are highly dependent on the choice of the geometry of the parametric manifold.
-For more precisely capturing the information for the quantum state itself, the concept of Krylov complexity was proposed using the Lanczos algorithm viswanath1994recursio , which is explicitly related to the quantum state without needing the parametric manifold.
-Since this method is free of ambiguities compared with the previous two ones, there will be similarities and differences between them.
-According to the concept of Krylov complexity, the universal operator growth was proposed Parker:2018yvk .
-Once the precise quantum state is given, we could analyze its Krylov complexity, Krylov entropy, and Lanczos coefficient based on the Lanczos algorithm viswanath1994recursio .
-For the connection between the Krylov complexity and the circuit complexity, Ref.
-Aguilar-Gutierrez:2023nyk has shown that they cannot be compatible with each other.
+The Lanczos algorithm offers a framework for constructing wave functions in closed and open quantum systems from their Hamiltonians.
+Since the early universe is inherently an open system, we employ this algorithm to investigate Krylov complexity across various cosmological phases: inflation, radiation domination (RD), and matter domination (MD).
+Our results highlight a clear distinction in Krylov complexity between the closed- and open-system methodologies.
+To accurately capture the influence of potentials during RD and MD, we examine a set of inflationary potentials, including the Higgs potential, $R^{2}$ inflation, and chaotic inflation, while incorporating violations of slow-roll conditions.
+This study is conducted in conformal time through the preheating stage.
+Numerically, we find that the evolution of Krylov complexity and Krylov entropy shows remarkable similarity across different potentials during RD and MD.
+Furthermore, we rigorously construct an open two-mode squeezed state using the second kind of Meixner polynomial.
+Based on this construction, we derive for the first time the evolution equations for the squeezing parameter $r_{k}$ and phase $\phi_{k}$ in terms of the scale factor.
+Our analysis indicates that dissipative effects lead to rapid decoherence-like behavior.
+In addition, we observe that the inflationary universe behaves as a strongly dissipative system, whereas during the RD and MD epochs the universe exhibits weak dissipative characteristics.
+This work opens new perspectives for studying the universe from a quantum-informational viewpoint.
+I Introduction
+The concept of complexity has become increasingly important in high-energy physics, although a unified definition is still lacking.
+Nonetheless, several computational frameworks have been proposed to evaluate it.
+One major approach is computational complexity, which quantifies the minimum number of logical steps required to perform a given task Aaronson:2016vto .
+Another significant line of research stems from Nielsen and collaborators, who introduced a geometric method for assessing complexity Nielsen:2005mkt ; Dowling:2006tnk ; Nielsen:2006cea .
+A related framework connects complexity to the Fubini–Study distance in the geometry of information space Chapman:2017rqy .
+However, both geometric approaches depend strongly on the choice of metric on the underlying parameter manifold.
+A more recent development is the notion of Krylov complexity, which circumvents the need to construct a parameter manifold altogether viswanath1994recursio .
+By operating directly in the operator Hilbert space, this approach avoids certain ambiguities inherent in geometric formulations, though the precise connections and distinctions between Krylov complexity and other frameworks remain an active area of research.
+Within this paradigm, the concept of universal operator growth has been proposed Parker:2018yvk , suggesting that the late-time dynamics of quantum operators exhibit universal features encoded in their Krylov complexity.
+Given a specific quantum operator, one can explicitly compute its Krylov complexity, associated Krylov entropy, and the sequence of Lanczos coefficients using the recursive Lanczos algorithm viswanath1994recursio .
+The relationship between Krylov complexity and circuit complexity has been examined from contrasting viewpoints: while Ref.
+Aguilar-Gutierrez:2023nyk argues that the two notions are generally incompatible.
 However, Ref.
-Caputa:2021sib demonstrated that the Krylov complexity was proportional to the distance of the ”Fubuni-Study” metric.
-Recently, the Krylov complexity has been broadly applied to condensed matter and high-energy physics.
-Ref.
-Muck:2022xfc has systematically analyzed the Krylov complexity in various polynomials dubbed as the orthogonal basis.
-Due to the importance of the SKY model in high energy physics, the Krylov complexity can also be implemented into this field Rabinovici:2020ryf ; Jian:2020qpp ; He:2022ryk .
-Meanwhile, the Krylov complexity can be applied to the generalized coherent state Patramanis:2021lkx , Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , conformal field theory Dymarsky:2021bjq ; Caputa:2021ori ; Kundu:2023hbk , topological phases of matter Caputa:2022eye and even extend into the quantum field theory He:2024hkw ; He:2024xjp .
-Ref.
-Dymarsky:2019elm proved that the chaos can be understood as a delocalization in Krylov complexity.
-Furthermore, one could observe the exponential growth of Krylov complexity in integrable systems with saddle-dominated scrambling, whose universal features can be verified by Huh:2023jxt .
-More recent developments of Krylov complexity can be found in Refs.
+Caputa:2021sib shows that, under certain conditions, Krylov complexity is proportional to the distance measured by the Fubini–Study metric.
+Recently, Krylov complexity has found broad applications across both condensed matter and high-energy physics.
+A systematic study employing various orthogonal polynomial bases was carried out in Ref.
+Muck:2022xfc , highlighting the flexibility and robustness of the Krylov framework.
+Given the central role of the Sachdev–Ye–Kitaev (SYK) model in understanding quantum chaos and holography, Krylov complexity has been actively explored in this context as well Rabinovici:2020ryf ; Jian:2020qpp ; He:2022ryk .
+Beyond the SYK model, it has been applied to the analysis of generalized coherent states Patramanis:2021lkx , spin systems such as the Ising and Heisenberg models Cao:2020zls ; Trigueros:2021rwj ; Heveling:2022hth , and conformal field theories Dymarsky:2021bjq ; Caputa:2021ori ; Kundu:2023hbk .
+More recently, the formalism has been extended to topological phases of matter Caputa:2022eye and even to quantum field theory He:2024hkw ; He:2024xjp , underscoring its growing relevance as a universal diagnostic of operator dynamics and quantum complexity.
+It was shown in Ref.
+Dymarsky:2019elm that quantum chaos can be interpreted as delocalization in Krylov space, an insight that has reshaped the understanding of operator growth in chaotic systems.
+Interestingly, exponential growth of Krylov complexity has also been observed in certain integrable models exhibiting saddle-dominated scrambling, with universal features confirmed in Ref.
+Huh:2023jxt .
+Since then, the scope of Krylov complexity has continued to expand rapidly.
+Recent developments include its application to open quantum systems, non-Hermitian dynamics, holography, random matrix theory, and out-of-equilibrium phenomena—see, e.g., Refs.
 Erdmenger:2023wjg ; Hashimoto:2023swv ; Vasli:2023syq ; Gill:2023umm ; Bhattacharjee:2023uwx ; Adhikari:2022whf ; Loc:2024oen ; Caputa:2024vrn ; Basu:2024tgg ; Sasaki:2024puk ; Caputa:2024xkp ; Bhattacharjee:2022qjw ; Sahu:2024opm ; Kim:2021okd ; Chen:2024imd ; Bhattacharjee:2024yxj ; Sanchez-Garrido:2024pcy ; Chattopadhyay:2024pdj ; Balasubramanian:2024ghv ; Bhattacharya:2024hto ; Mohan:2023btr .
-Especailly, the concept of Krylov complexity can be extended into in various fields Nandy:2024htc .
-We aim to investigate the Krylov complexity for the whole early universe, including the inflationary period, radiation-dominated period (RD), and matter-domination period (MD).
-Some research has examined the computational complexity for inflation Choudhury:2020hil ; Bhargava:2020fhl ; Lehners:2020pem ; Bhattacharyya:2020rpy ; Adhikari:2021ked , where the analysis demonstrated that there is oscillation at the very beginning of inflation, and the evolution was similar in post inflation period.
-Once considering the quantum effects via various quantum frameworks, the complexity will show the irregular oscillation during inflation providing a new method for distinguishing various inflationary models Li:2023ekd .
-Quite a few researches are investigating the Krylov complexity in the early universe.
-Ref.
-Adhikari:2022oxr is the first one who investigate Krylov complexity in inflation by evaluating the values of non-trivial sound speed, where their method is based on the closed system’s method.
-However, our universe is an open system.
-In light of the open system’s method, Ref.
-Li:2024kfm investigated the Krylov complexity within the framework of modified dispersion relation in inflation, indicating that the inflation behaves as a strong dissipative system and the overall trend of Krylov complexity is increasing.
-The key ingredient for exploring the Krylov complexity with the open system’s method is constructing the wave function in light of Ref.
-Bhattacharya:2022gbz , whose wave function wave from Li:2024kfm will serve as our starting point in this work.
-In previous studies, Krylov complexity was applied to the early universe using the thermal state.
-Refs Li:2024ljz ; Li:2024iji examined Krylov complexity by considering the contribution of the potential (quadratic term) and the violation of slow-roll conditions.
-The potential was approximated to be the quadratic term around the minimum in our previous work Li:2024ljz , which may not capture fully precise information.
-In the upcoming work, we plan to include more models such as the Higgs potential Higgs:1964pj , the potential of Starobinsky inflation ($R^{2}$ inflation) Starobinsky:1980te , and the potential of chaotic inflation Linde:1983gd .
-These models can cover most inflationary models, especially those favored by current observations, which indicate that the potential shape is a plateau.
+Notably, the framework has also been generalized beyond standard unitary quantum mechanics, with extensions to stochastic dynamics, classical systems, and information-theoretic settings Nandy:2024htc .
+In this work, we aim to investigate Krylov complexity throughout the entire evolution of the early universe, spanning the inflationary epoch, the subsequent RD era, and the MD era.
+Previous studies have explored computational complexity during inflation Choudhury:2020hil ; Bhargava:2020fhl ; Lehners:2020pem ; Bhattacharyya:2020rpy ; Adhikari:2021ked , uncovering an oscillatory behavior at the onset of inflation and a qualitatively similar pattern in the post-inflationary phase.
+When quantum effects are incorporated through various quantum-information-theoretic frameworks, the complexity displays irregular oscillations during inflation, a feature that provides a novel diagnostic for distinguishing between competing inflationary models Li:2023ekd .
+While several recent works have begun to explore Krylov complexity in the context of the early universe, Ref.
+Adhikari:2022oxr stands out as the first to investigate it during inflation, specifically accounting for non-trivial sound speed effects within a closed quantum system framework.
+However, given that the early universe is more realistically modeled as an open quantum system, interacting with unobservable environmental degrees of freedom, a more physically complete treatment requires going beyond unitary evolution.
+Addressing this, Ref.
+Li:2024kfm adopted an open quantum system approach to study Krylov complexity under a modified dispersion relation during inflation.
+Their analysis reveals that inflation behaves as a strongly dissipative process, with Krylov complexity displaying a persistent overall increase.
+A crucial ingredient in this open-system formulation is the explicit construction of the time-evolving wave function, as systematically developed in Ref.
+Bhattacharya:2022gbz .
+In the present work, we take the wave function derived in Ref.
+Li:2024kfm as our starting point to extend the analysis across the full cosmological timeline, from inflation through RD and MD eras.
+In previous studies, Krylov complexity has been applied to the early universe using a thermal state as the reference.
+Refs.
+Li:2024ljz ; Li:2024iji extended this framework by incorporating the contribution of the scalar potential—approximated by its quadratic term, and examined the impact of slow-roll violation on complexity growth.
+In our earlier work Li:2024ljz , the potential was expanded quadratically around its minimum; while analytically convenient, this approximation may not fully capture the dynamics of realistic inflationary scenarios.
 It has been suggested by Ref.
-Galante:2014ifa that many inflationary models can be unified into the so-called $\alpha$ attractors, whose potential is similar to $R^{2}$ inflationary potential.
-Additionally, Ref.
-Albrecht:1992kf demonstrated a relation between the Bunch-Davies vacuum and the two-mode squeezed state.
-We will continue to use the two-mode squeezed state to investigate the Krylov complexity for the entire early universe in light of these two factors.
-This paper will be organized as follows.
+Galante:2014ifa that many inflationary models are unified into the framework of $\alpha$-attractors, whose potential form closely resembles that of $R^{2}$ inflation.
+Moreover, Ref.
+Albrecht:1992kf established a connection between the Bunch-Davies vacuum and the two-mode squeezed state.
+Motivated by these two key insights, we will continue to employ the two-mode squeezed state formalism to study Krylov complexity in the context of the entire early universe.
+To improve physical fidelity, we plan in upcoming work to generalize the analysis to well-motivated inflationary potentials, including the Higgs potential Higgs:1964pj , Starobinsky’s inflation Starobinsky:1980te , and chaotic inflation Linde:1983gd .
+These models encompass most inflationary models and are especially relevant in light of current observational constraints, which strongly favor plateau-like potentials.
+This extension will allow for a more accurate and observationally grounded exploration of quantum complexity during inflation.
+This paper is organized as follows.
 In Sec.
-2, we will introduce the Lanczos algorithm.
+II, we will introduce the Lanczos algorithm.
 Sec.
-3 will give a brief introduction to the early universe, including inflation, RD, and MD, and three different kinds of potentials.
+III will provide a brief review of the early universe, covering inflation, RD, and MD eras, along with three representative inflationary potentials.
+In Sec.
+IV, we will present the two-mode squeezed state formalism and numerically solve for the parameter $r_{k}$ across different cosmological periods.
 Sec.
-4 will give the two-mode squeezed state, and the numerical solutions of $r_{k}$ in different periods.
-Sec.
-5 will investigate the Krylov complexity in a method of closed system.
-Sec.
-5.2 will investigate the Krylov entropy in a method of closed system.
-Sec.
-6 will give the Lanczos coefficient and Krylov complexity with the approach of open system, where we are the first to derive the evolution equation of $\phi_{k}$ and $r_{k}$ based on the open two-mode squeezed state.
-Finally, the conclusion and outlook will be given in Sec.
-7.
-2 Lanczos algorithm with the approach of closed system
-In this section, we will mainly introduce the Lanczos algorithm with the approach of closed system. First, the operator $\mathcal{O}(t)$ in Heisenberg picture needs to be defined as follows,
+V will investigate the Krylov complexity in the framework of a closed-system methodology, which Sec.
+V.2 will examine the corresponding Krylov entropy.
+In Sec.
+VI, we will analyze the Lanczos coefficients and Krylov complexity using an open-system methodology; notably, we will present the evolution equations for $\phi_{k}$ and $r_{k}$ derived within the open two-mode squeezed state formalism in the first time.
+Finally, conclusions and outlook will be given in Sec.
+VII.
+In this work, we will work in Planck units, namely $c=M_{P}=G=1$.
+II Lanczos algorithm with closed-system methodology
+In this section, we will mainly review the Lanczos algorithm with closed-system methodology. First, the operator $\mathcal{O}(t)$ in Heisenberg picture needs to be defined as follows,
 $\partial_{t}\mathcal{O}(t)=i[H,\mathcal{O}(t)]$ (1)
-where H is the Hamilton and its solution is
+where H is the Hamiltonian and its solution is
 $\mathcal{O}(t)=e^{iHt}\mathcal{O}e^{-iHt}$ (2)
-Next, let’s define the Liouvillian super-operator $\mathcal{L}_{X}$ as $\mathcal{L}_{X}Y=[X,Y]$, where [,][,][ , ] denotes the commutator.
-We apply this to Eq.
-(2), the operator can also be expressed as
+Next, let’s define the Liouvillian super-operator $\mathcal{L}_{X}$ as $\mathcal{L}_{X}Y=[X,Y]$, where $[,]$ denotes the commutator. We apply this to Eq. (2), the operator is also expressed as
 $\mathcal{O}(t)=e^{i\mathcal{L}t}\mathcal{O}=\sum_{n=0}^{\infty}\frac{(it)^{n}}{n!}\mathcal{L}^{n}\mathcal{O}(0)=\sum_{n=0}^{\infty}\frac{(it)^{n}}{n!}\tilde{\mathcal{O}}_{n}$ (3)
 where $\mathcal{L}^{n}\mathcal{O}=\tilde{\mathcal{O}}_{n}=[H,\tilde{O}_{n-1}]$ which can serve as the basis for the Hilbert space.
-$\begin{split}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}),\mathcal{L}^{1}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}_{1}),\mathcal{L}^{2}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}_{2}),\mathcal{L}^{3}\mathcal{O}\equiv|{\tilde{\mathcal{O}}}_{3})...\end{split}$ (4)
-However, there are not orthogonal bases. In order to construct an orthogonal basis, we use the Lanczos algorithm. The first two operators become
+[…]
+However, they are not orthogonal bases. In order to construct an orthogonal basis, we use the Lanczos algorithm. The first two operators become
 $\mathcal{O}_{0}=|\tilde{\mathcal{O}}_{0})=\mathcal{O},~~~|\mathcal{O}_{1})=b_{1}^{-1}\mathcal{L}|\mathcal{O}_{0})$ (5)
-where $b_{1}=\sqrt{(\tilde{\mathcal{O}}_{0}\mathcal{L}|\mathcal{L}\tilde{\mathcal{O}}_{0})}$ desribing the normalized vector. And the next orthogonal basis can be written by
+where $b_{1}=\sqrt{(\tilde{\mathcal{O}}_{0}\mathcal{L}|\mathcal{L}\tilde{\mathcal{O}}_{0})}$ describing the normalized vector. And the next orthogonal basis is written by
 $|\mathcal{O}_{n})=b_{n}^{-1}|A_{n})$ (6)
@@
 $|A_{n})=\mathcal{L}|\mathcal{O}_{n-1})-b_{n-1}|\mathcal{O}_{n-2}),~~~b_{n}=\sqrt{(A_{n}|A_{n})}$ (7)
-where $b_{n}$ represents the Lanczos coefficients. This iterative relation will stop and produce the finite orthogonal Krylov basis if $b_{n}=0$. Subsequently, we can express Eq. (3) as follows,
+where $b_{n}$ represents the Lanczos coefficients. This iterative relation will stop and produce the finite orthogonal Krylov basis if $b_{n}=0$. Subsequently, we express Eq. (3) as follows,
 $\mathcal{O}(t)=e^{i\mathcal{L}t}\mathcal{O}=\sum_{n=0}^{\infty}(i)^{n}\phi_{n}(t)|\mathcal{O}_{n})$ (8)
@@
 $\partial_{t}\phi_{n}(t)=b_{n}\phi_{n-1}-b_{n+1}\phi_{n+1}$ (9)
-Next, we can define the Krylov complexity using the wave function $\phi_{n}$ as follows
+Next, we define the Krylov complexity using the wave function $\phi_{n}$ as follows
 $K=\sum_{n}n|\phi_{n}|^{2}$ (10)
-Additionally, the Lanczos coefficient can be bounded via Parker:2018yvk ,
-$b_{n}\leq\alpha n+\eta$ (11)
-where $\alpha$ and $\eta$ encodes different information in various models.
-It should be noticed that $b_{n}=\alpha n+\eta$, meaning that the system is a maximal chaotic dynamical system.
-For the case of maximal chaos, we could build the relation between the Lyapunov index and $\gamma$ and $\alpha$ as follows,
+Additionally, the Lanczos coefficient is bounded via Parker:2018yvk ,
+$b_{n}\leq\alpha n+\eta,$ (11)
+where $\alpha$ and $\eta$ encode different information in various models.
+Note that $b_{n}=\alpha n+\eta$, which indicates the system is a maximally chaotic dynamical system.
+For the case of maximal chaos, we obtain the relationship between the Lyapunov exponent and $\gamma$ and $\alpha$ as follows
 $\lambda=2\alpha$ (12)
 In this work, we will investigate the Krylov complexity for the entire early universe, thus the formula of scale factor $a(\eta)$ should be introduced.
-3 Some setup of early universe
-In this section, we will first give some basis of scale factor in distinctive periods.
-Following the notation of our previous work Li:2024ljz , one of the core goal of this work is examining the impact of inflationary potential for the Krylov complexity in RD and MD due to the violation of slow-roll conditions.
-The Mukhannov-Sasaki variable contains the contribution of potential which cannot show the explicit contribution to the Krylov complexity.
-That is why we just perturb the inflaton field explicitly in this paper.
-Thereafter, we will give some very important inflationary potentials, namely, the chaotic inflationary potential, Starobinksy’s inflationary potential ($R^{2}$ inflation) and Higgs potential.
-3.1 Inflation, RD and MD
-We will follow the notation of Baumann:2009ds to introduce the scale factor. The background metric is the Friedman-Lemaitre-Robertson-Walker metric,
+III Some setup of early universe
+In this section, we first review some fundamental aspects of the scale factor in different cosmological epochs.
+Following the notation of our previous work Li:2024ljz , a central goal of this study is to examine the impact of the inflationary potential on Krylov complexity during the RD and MD eras, particularly due to the violation of slow-roll conditions.
+The Mukhanov-Sasaki variable incorporates contributions from the potential in a way that does not explicitly reveal its influence on Krylov complexity.
+Therefore, in this work, we perturb the inflaton field directly.
+We then introduce several important inflationary potentials: the chaotic inflationary potential, the Starobinsky potential ($R^{2}$ inflation), and the Higgs potential.
+Finally, we employ preheating dynamics to numerically evaluate the effective mass of the inflaton with respect to the scale factor.
+III.1 Inflation, RD and MD
+We will follow the notation of Baumann:2009ds to introduce the scale factor. The background metric is the Friedman-Lemaitre-Robertson-Walker (FLRW) metric,
 $ds^{2}=a^{2}(\eta)(-d\eta^{2}+d\vec{x}^{2}),$ (13)
-where $a(\eta)$ is the scale factor.
-Here, we will assume a single component universe, which means that the universe will experice the inflation, RD and MD, respectively.
-We could use equation of state to characterize the universe via
+where $a(\eta)$ is the scale factor expressed in conformal time $\eta$.
+Alternatively, the FLRW metric can also be written in terms of physical time $t$ as $ds^{2}=-dt^{2}+a(t)^{2}d\vec{x}^{2}$, from which the relation $dt=ad\eta$ follows explicitly.
+Here, we consider a universe dominated by a single component during each epoch, meaning it undergoes successive phases of inflation, RD, and MD.
+The evolution of the scale factor in each era can be characterized by the equation of state parameter $w$ via
 $w_{I}=\frac{P_{I}}{\rho_{I}},$ (14)
-where $P_{I}$ and $\rho_{I}$ is the pressure and energy density of the corresponding periods parameterized by $I$.
-For obtaining the relation between the conformal time and the scale factor, we need the comoving distance,
+where $P_{I}$ and $\rho_{I}$ are the pressure and energy density, respectively, in each epoch $I$.
+To derive the relation between conformal time and the scale factor, we make use of the comoving distance,
 $\chi_{ph}(\eta)=\int_{t_{i}}^{t}\frac{dt}{a}=\int_{\ln a_{i}}^{\ln a}(aH)^{-1}d\ln a$ (15)
-where $\chi_{ph}(\eta)$ is the comoving distance in the event horizon of particle. In light of $w_{I}$, the Hubble radius can be replaced by
+where $\chi_{\mathrm{ph}}(\eta)$ is the comoving particle horizon distance, and we have used the definition of the Hubble parameter $H=\dot{a}/a$, with $\dot{a}=da/dt$.
+In terms of the equation of state parameter $w_{I}$, the Hubble radius can be expressed as
 $(aH)^{-1}=H_{0}^{-1}a^{\frac{1}{2}(1+3\omega)}$ (16)
-where we use $\omega$ to take place of $\omega_{I}$ for simplify. Integrating equation (15) yields the relation between the scale factor and the conformal time
-$\eta=\frac{2H_{0}^{-1}}{(1+3\omega)}a^{\frac{1}{2}(1+3\omega)}$ (17)
-Consequently, we can derive the relation for various time periods
+where, for simplicity, we replace $\omega_{I}$ with $\omega$. Integrating Eq. (15) yields an explicit functional relation between the scale factor and the conformal time,
+$\eta=\frac{2H_{0}^{-1}}{(1+3\omega)}a^{\frac{1}{2}(1+3\omega)}.$ (17)
+Consequently, we derive the relation for various time periods
 $\eta=\begin{cases}&-(aH_{0})^{-1},\ \ (\omega=-1),\ \ \rm inflation\\ &aH_{0}^{-1},\ \ (\omega=\frac{1}{3}),\ \ \rm RD\\ &2\sqrt{a}H_{0}^{-1},\ \ (\omega=0),\ \ \rm MD\end{cases}$ (18)
-One could see that the various values of $\omega$ describe the distinctive periods.
-In the following text, we will discuss the fundamental basis for the different periods in the evolution of the universe.
-Initially, the universe is created by a singularity, followed by a period of exponential expansion known as inflation.
-During this phase, the curvature perturbation transitions from the quantum level to the classical level after the horizon exits, leaving an imprint on the Cosmic Microwave Background (CMB).
-Subsequently, the universe enters the Radiation Dominated (RD) era, characterized by an energy density given by the equation $\rho_{\rm rad}(t)=\rho_{\rm rad_{0}}\frac{a_{0}^{4}}{a(t)^{4}}$, where $\rho_{\rm rad_{0}}$ and $a_{0}$ represent the initial values of the energy density and the scale factor of RD.
-During the RD era, particles are relativistic, but as the temperature decreases, they become increasingly non-relativistic.
-This leads the universe into the matter-dominated (MD) epoch, where the energy density is given by $\rho_{\rm mat_{0}}(t)=\rho_{\rm mat_{0}}\frac{a_{0}^{3}}{a(t)^{3}}$.
-The transition from the RD to the MD epoch occurs when the energy density of radiation becomes equal to the energy density of matter.
-Once entering MD, there is a process called preheating where the particles will be generated via this process.
-First, we give the full action for the single field inflation,
-[…]
+It shows that the various values of $\omega$ describe the distinctive periods.
+In the following, we outline the fundamental evolution of the universe across its distinct epochs.
+The universe emerges from a singularity and undergoes a phase of exponential expansion known as inflation.
+During this period, curvature perturbations transition from quantum to classical after horizon exit, becoming imprinted on the Cosmic Microwave Background (CMB).
+Subsequently, the universe enters the RD era, characterized by an energy density $\rho_{\rm rad}(t)=\rho_{\rm rad_{0}},a_{0}^{4}/a(t)^{4}$, where $\rho_{\rm rad_{0}}$ and $a_{0}$ denote the initial energy density and scale factor at the onset of RD.
+In this epoch, particles are relativistic.
+As the temperature drops, particles become increasingly non-relativistic, leading the universe into the MD era, where the energy density evolves as $\rho_{\rm mat}(t)=\rho_{\rm mat_{0}},a_{0}^{3}/a(t)^{3}$.
+The transition from RD to MD occurs when the energy densities of radiation and matter equalize.
+During the subsequent MD epoch, a preheating process can generate particles.
+We begin by presenting the full action for single-field inflation:
+$S=\int d\eta d^{3}\vec{x}\sqrt{-g}\bigg[\frac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-V(\phi)\bigg],$ (19)
 where $V(\phi)$ is the potential of inflation. Following the method of Kofman:1997yn ; Shtanov:1994ce , the equation of motion for background $\phi(t)$ in RD and MD is
 $\ddot{\phi}(t)+3H\dot{\phi}(t)+V(\phi),_{\phi}=0,$ (20)
-where $\ddot{\phi}(t)=\partial_{t}^{2}\phi(t)$ and the similar work can be found in our previous works Liu:2019xhn ; Liu:2021rgq .
-Defining with the effective mass $m_{\rm eff}=\frac{d^{2}V}{d\phi^{2}}$ and changing variable as $\tilde{\phi}(t)=a^{3/2}\phi(t)$, Eq.
-(20) will become
+where $\ddot{\phi}(t)=\partial_{t}^{2}\phi(t)$, $V(\phi),_{\phi}=\frac{dV(\phi)}{d\phi}$ and the similar work is found in our previous works Liu:2019xhn ; Liu:2021rgq .
+Defining with the effective mass as follows,
+$m_{\rm eff}=\frac{d^{2}V}{d\phi^{2}}=V_{,\phi\phi}.$ (21)
+Here, we explain why the effective mass is defined as $m_{\rm eff}=d^{2}V/d\phi^{2}$.
+For the different inflationary potentials in Eq.
+(23), we take $V_{\rm chaotic}$ as an example, where one can directly show that $m_{\rm eff}^{2}=m^{2}$.
+Based on this, we adopt the same definition of effective mass for all inflationary potentials considered.
+To solve Eq.
+(20), we introduce a new variable $\tilde{\phi}(t)=a^{3/2}\phi(t)$, which transforms Eq.
+(20) into
 $\ddot{\tilde{\phi}}(t)+m_{\rm eff}^{2}\tilde{\phi}(t)=0,$ (22)
-the solution of Eq.
-(21) is $\phi(t)=a^{-3/2}\cos(m_{\rm eff}t)$ where it is the solution of $\phi(t)$ in preheating period and we have neglecte the term of $\frac{9}{4}H^{2}\tilde{\phi}+\frac{3}{2}\dot{H}\tilde{\phi}$ since it is much smaller compared with potential.
-It should be noticed that the solution is in terms of physical time $t$.
-In this paper, we will use $d\eta=\frac{dt}{a}$ to transform into the conformal time for the latter investigations, which is denoted by Eq.
+The solution to Eq.
+(22) is $\phi(t)\propto a^{-3/2}\cos(m_{\rm eff}t)$, which describes the behavior of $\phi(t)$ during the preheating period.
+Here, the terms $\frac{9}{4}H^{2}\tilde{\phi}+\frac{3}{2}\dot{H}\tilde{\phi}$ have been neglected, as they are significantly smaller than the potential contribution.
+Note that this solution is expressed in physical time $t$.
+For the subsequent analysis in this paper, we will transform it into conformal time using $d\eta=dt/a$, as given by Eq.
 (18).
-3.2 Various inflationary potentials
-There are lots of inflationary models that represent distinctive inflationary potentials.
-Current observations favor the concave potentials Planck:2018jri , where the potential of $R^{2}$ inflation belongs to this potential.
-Additionally, the $R^{2}$ inflation agrees with the observations from inflation to large scales.
-From the theoretical aspect, we could see that many inflationary models can be unified into the so-called $\alpha$ attractors inflationary models Galante:2014ifa , where the other models and $R^{2}$ inflationary model belong to the same kind of inflationary models.
-Being armed with these two reasons, we will introduce $R^{2}$ inflation.
-According to the current standard model, the Higgs field is the only scalar field in nature.
-Therefore, the Higgs potential is our other option Higgs:1964pj .
-The final choice is the chaotic inflation whose potential is the quadratic term Linde:1983gd .
-Finally, we could see that our choices could cover the most inflationary models.
-Here, we list the formulas of distinctive inflationary potentials as follows,
-$\begin{cases}&\ V_{1}=\frac{1}{2}m^{2}\phi^{2},\\ &\ V_{2}=\frac{3}{4}M^{2}M_{p}^{2}(1-e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}})^{2},\\ &\ V_{3}=\frac{1}{2}m^{2}\phi^{2}+\frac{1}{4}\lambda\phi^{4},\end{cases}$ (22)
-where $M_{p}=(8\pi G)^{-\frac{1}{2}}$ is the definition of the reduced Planck mass.
-Additionally, $M$ represents the energy scale, and its value can be determined based on the amplitude of the observed power spectrum for primordial perturbations.
-Being armed with the solution of Eq.
-(20) and $dt=ad\eta$, we could rewrite $V_{\phi\phi}$ as
-$\begin{cases}&\ V_{1,\phi\phi}=m^{2},\\ &\ V_{2,\phi\phi}=M^{2}\Bigl{(}-e^{-\sqrt{\frac{2}{3}}\frac{Aa^{-\frac{3}{2}}\sin(m_{\rm eff}a\eta)}{M_{p}}}+2e^{-2\sqrt{\frac{2}{3}}\frac{Aa^{-\frac{3}{2}}\sin(m_{\rm eff}a\eta)}{M_{p}}}\Bigr{)},\\ &\ V_{3,\phi\phi}=m^{2}+3\lambda A^{2}a^{-3}\sin^{2}(m_{\rm eff}a\eta)\end{cases}$ (23)
-where $\phi(t)=Aa^{-\frac{3}{2}}\sin(m_{\text{eff}}t)$.
-For the effective mass $m_{\text{eff}}$, we have $m_{\text{eff}}^{2}=\frac{d^{2}V(x)}{dx^{2}}$, and it is evident that $m_{\text{eff}}=\sqrt{V_{i,\phi\phi}}$ in our work $(i=1,2,3)$.
-According to effective mass for the distinctive inflationary models, we need the iterations for $V_{2}$ and $V_{3}$ in terms of $\log_{10}a$ followed by Li:2024kfm .
-To determine whether the functions $V_{2}$ and $V_{3}$ converge within the defined interval, we performed iterations at multiple sampling points ranging from $\log_{10}a=0$ to $\log_{10}a=6$ with the step $0.05$.
-Our findings reveal that all points ranging from $\log_{10}a=0$ to $\log_{10}a=6$ are all convergent.
-Figure 1: The numeric of $V_{,\phi\phi}$ at first 50 iterations of potential $V_{2}$ and $V_{3}$ in $\log_{10}a=0.2$ and $\log_{10}a=2.8$, in which we set $M_{p}=1$, $m=10^{-6}$; $M^{2}=10^{-12}$ at $V_{2}$, and $\lambda A^{2}=10^{-4}$ at $V_{3}$.
+III.2 Various inflationary potentials
+In this subsection, we will use Eq.
+(22) to establish a relation between the scale factor and various inflationary potentials.
+There exist many inflationary models, each characterized by a distinct potential.
+Current observational data favor concave potentials Planck:2018jri , a category that includes the potential of $R^{2}$ inflation.
+Furthermore, $R^{2}$ inflation remains consistent with observations from inflationary to large scales.
+From a theoretical perspective, many inflationary models can be unified within the framework of $\alpha$-attractor models Galante:2014ifa , which also encompass $R^{2}$ inflation.
+Motivated by these two considerations, we include $R^{2}$ inflation in our analysis.
+According to the Standard Model, the Higgs field is the only fundamental scalar field in nature; therefore, the Higgs potential serves as another natural choice Higgs:1964pj .
+Lastly, we consider chaotic inflation with a quadratic potential Linde:1983gd .
+Collectively, these choices cover a representative range of inflationary models.
+We consider the following inflationary potentials:
+$\begin{cases}&\ V_{\rm chaotic}=\frac{1}{2}m^{2}\phi^{2},\\ &\ V_{\rm R^{2}}=\frac{3}{4}M^{2}M_{p}^{2}(1-e^{-\sqrt{\frac{2}{3}}\frac{\phi}{M_{p}}})^{2},\\ &\ V_{\rm Higss}=\frac{1}{2}m^{2}\phi^{2}+\frac{1}{4}\lambda\phi^{4},\end{cases}$ (23)
+where $M_{p}=(8\pi G)^{-\frac{1}{2}}$ is the reduced Planck mass.
+Here, $M$ denotes an energy scale whose value is fixed by the observed amplitude of the primordial power spectrum.
+Using the solution of Eq.
+(20) and the relation $dt=ad\eta$, we can express $V,_{\phi\phi}$ as
+$\begin{cases}&\ V_{\rm chaotic,\phi\phi}=m^{2},\\ &\ V_{\rm R^{2},\phi\phi}=M^{2}\Bigl(-e^{-\sqrt{\frac{2}{3}}\frac{Aa^{-\frac{3}{2}}\sin(m_{\rm eff}a\eta)}{M_{p}}}+2e^{-2\sqrt{\frac{2}{3}}\frac{Aa^{-\frac{3}{2}}\sin(m_{\rm eff}a\eta)}{M_{p}}}\Bigr),\\ &\ V_{\rm Higss,\phi\phi}=m^{2}+3\lambda A^{2}a^{-3}\sin^{2}(m_{\rm eff}a\eta)\end{cases}$ (24)
+where $\phi(t)=Aa^{-\frac{3}{2}}\sin(m_{\text{eff}}t)$ and we have used $dt=ad\eta$.
+The definition of effective mass in Eq.
+(21) has been employed, and clearly $m_{\text{eff}}=\sqrt{V_{i,\phi\phi}}$ in our work ($\rm i=chaotic,R^{2},Higgs$).
+To simplify the subsequent numerical simulations, we define the variable $y=\log_{10}a$.
+To numerically obtain the effective mass of the distinct inflationary models $V_{\rm R^{2}}$ and $V_{\rm Higgs}$, iterative evaluations in terms of $\log_{10}a$ are required, based on Eq.
+(24).
+Let us clarify the concept of iterations for the effective mass.
+The Newton-Raphson method is adopted for its computational efficiency and fast convergence rate.
+In (24), the effective mass $m_{\rm eff}$ defined by (21) carries the same physical meaning as the left-hand side of (24).
+Thus, the relation (24) naturally lends itself to an iterative scheme, requiring an initial value $m_{0}$ for the effective mass.
+This value ($m_{0}^{2}=10^{-12}$) can be fixed by the COBE normalization Planck:2015sxf .
+In our implementation, we perform 50 iterations for the effective mass, meaning that for each single value of $y$, the calculation is repeated 50 times.
+To check the convergence of $V_{\rm R^{2},\phi\phi}$ and $V_{\rm Higgs,\phi\phi}$ within the defined interval, we performed iterations at multiple sampling points from $\log_{10}a=0$ to $\log_{10}a=6$ with a step size of $0.05$ (this range covers from RD to MD).
+We find that all points in the range $\log_{10}a=0$ to $\log_{10}a=6$ are convergent.
+Figure 1: The numeric of $V_{,\phi\phi}$ at first 50 iterations of $V_{\rm R^{2},\phi\phi}$ and $V_{\rm Higgs,\phi\phi}$ in $\log_{10}a=0.2$ and $\log_{10}a=2.8$, in which we set $M_{p}=1$, $m=10^{-6}$; $M^{2}=10^{-12}$ at $V_{\rm R^{2}}$, and $\lambda A^{2}=10^{-4}$ at $V_{\rm Higss}$.
 As an illustration, Fig.
-1 clearly shows the iterations for potential $V_{2}$ and $V_{3}$, where we have performed $50$ times iterations for two specific values of $\log_{10}a$.
-Let us mention the range of parameters as shown in Fig.
-1, the potential can be constraint by the COBE normalization Planck:2015sxf , whose values are bounded below $10^{-10}$ in Planck units, and the inflationary field is a large field and its maximal value is around $100$.
-Consequently, the effective mass for various potentials is around $10^{-6}$ and other parameters are set according to this logic.
-Figure 2: The numeric of $V_{,\phi\phi}$ in terms of $\log_{10}a$ for potential $V_{2}$, and $V_{3}$ in RD and MD, in which we replace its exact value with the third iteration of $V_{2}$ and $V_{3}$.
-We set $\theta_{0}=0$, $M_{p}=1$, $m=10^{-6}$; $M^{2}=10^{-12}$ at $V_{2}$, and $\lambda A^{2}=10^{-4}$ at $V_{3}$.
+1 shows the iterative results for $V_{\rm R^{2},\phi\phi}$ and $V_{\rm Higgs,\phi\phi}$, where 50 iterations were performed for two specific values of $\log_{10}a$.
+The parameters used in the figure are chosen as follows: the potential amplitude is constrained by the COBE normalization Planck:2015sxf , which bounds its value below $10^{-10}$ in Planck units, while the inflationary field is a large field with a maximum value around $100$ in these units.
+Consequently, the effective mass for the various potentials is of order $10^{-6}$, and all other parameters are set consistently with this scaling.
+Let us now discuss the effects of the distinctive parameters of the various potentials.
+For $V_{\text{chaotic}}$, its effective mass is explicitly set by the inflaton mass, whose value can be fixed via the COBE normalizations.
+As for $V_{R^{2}}$, it contains one parameter $M$, which acts as the amplitude of the effective mass in Eq.
+(24); consequently, the range of $M$ is also constrained by the COBE normalization.
+Finally, the Higgs potential $V_{\text{Higgs}}$ involves three parameters: the Higgs mass $m$, the coupling constant $\lambda$, and the amplitude of the inflaton during preheating $A$.
+Again, using the COBE normalizations, we can determine the upper bounds on $m^{2}$ and $\lambda A^{2}$; values exceeding these orders would violate the observational constraints.
+Thus, appropriate parameter values can be set from observational data.
+As illustrated in Fig.
+2, the effective masses for all the considered potentials lie within $10^{-12}$ in Planck units.
+Figure 2: The numeric of $V_{,\phi\phi}$ in terms of $\log_{10}a$ for potential $V_{\rm R^{2}}$, and $V_{\rm Higgs}$ in RD and MD, in which we replace its exact value with the iteration of $V_{\rm R^{2}}$, and $V_{\rm Higgs}$.
+We set $M_{p}=1$, $m=10^{-6}$; $M^{2}=10^{-12}$ for $V_{\rm R^{2}}$, and $\lambda A^{2}=10^{-4}$ for $V_{\rm Higgs}$.
 Fig.
-2 depicts the evolution of $V_{,\phi\phi}$ with respect to $\log_{10}a$ in RD and MD, using the third iteration of $V_{2}$ and $V_{3}$ via Fig.
-1.The evolution of inflation is neglected due to the violation of slow-roll conditions.
+2 shows the evolution of $V_{,\phi\phi}$ with respect to $\log_{10}a$ during the RD and MD eras, based on the iterative results for $V_{\rm R^{2}}$ and $V_{\rm Higgs}$ from Fig.
+1.
+The evolution of the inflationary phase itself is neglected here due to the violation of slow-roll conditions in these periods.
 From Fig.
-2, we could see that the effective mass of $R^{2}$ inflation will decrease in RD and then it will grow to the constant ($10^{-12}$) with the oscillations.
-For the potential of the Higgs field, it is different which shows that there is a peak in RD and after that, it will decrease into some constants.
-But, it indicates that the effective mass of $V_{2}$ and $V_{3}$ are of almost the same order based on the COBE normalization Planck:2015sxf .
-In this section, we provide an introduction to inflation, RD, and MD, which are characterized by the corresponding scale factor.
-We then present three essential inflationary potentials that encompass most types of inflationary models.
-Additionally, we provide an overview of preheating, during which the background of the inflaton field oscillates around the minimum.
-Subsequently, we incorporate the background solution for the inflaton into the numerical calculation of the effective mass for these three distinct potentials.
-4 Evolution of two-mode squeezed state
-In this section, we will follow the method of Li:2024kfm for investigating the evolution of $r_{k}$ and $\phi_{k}$ in inflation, RD and MD.
-We have evaluated the quantum circuit complexity in various inflationary frameworks Liu:2021nzx ; Li:2021kfq ; Li:2023ekd .
-4.1 Two-mode squeezed state
-For obtaining the two-mode squeezed state, we should introduce the unitary operator as follows,
+2, we observe that the effective mass for $R^{2}$ inflation ($V_{2}$) decreases during RD and then grows, with oscillations, toward a constant value of approximately $10^{-12}$.
+The behavior for the Higgs potential ($V_{\rm Higgs}$) is different: it exhibits a peak during RD, after which it decreases and settles into an approximately constant value.
+Notably, the effective masses for both $V_{\rm R^{2}}$ and $V_{\rm Higgs}$ are of nearly the same order of magnitude, consistent with the COBE normalization constraint Planck:2015sxf .
+In this section, we introduce the inflationary, RD, and MD eras, each described by its corresponding scale factor.
+We also present three representative inflationary potentials that cover a broad range of model types.
+An overview of preheating is given, during which the inflaton background oscillates around its potential minimum.
+Finally, we incorporate this background solution into a numerical evaluation of the effective mass for each of the three potentials.
+IV Evolution of two-mode squeezed state
+As we know, the curvature perturbation will become classical as the horizon exits.
+Ref.
+Grishchuk:1990bj already shown the amplification from quantum level to classical level during inflation is naturally a quantum squeezing process.
+Thus, the two-mode squeezed state is quite a natural choice for investigating the early universe.
+In this section, we will follow the method of Li:2024kfm to investigate the evolution of $r_{k}$ and $\phi_{k}$ during inflation, RD and MD.
+IV.1 Two-mode squeezed state
+First, we will simply recall the two-mode squeezed state. To obtain the two-mode squeezed state, we introduce the unitary operator as follows,
 $\mathcal{U}_{k}=\hat{\mathcal{S}}_{k}(\phi_{k},r_{k})\hat{\mathcal{R}}_{k}(\theta_{k})$ (25)
@@
 $\hat{\mathcal{S}}_{k}=\exp\bigg[r_{k}(\eta)\big(e^{-2i\phi_{k}(\eta)}\hat{c}_{\vec{k}}\hat{c}_{-\vec{k}}-e^{2i\phi_{k}(\eta)}\hat{c}^{\dagger}_{-\vec{k}}\hat{c}_{\vec{k}}^{\dagger}\big)\bigg]$ (26)
-^$\hat{\mathcal{R}}_{k}(\theta_{k})=\exp\bigg{[}-i\theta_{k}(\eta)\big{(}\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}+\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}\big{)}\bigg{]}$ (26)
-where $\hat{\mathcal{S}}_{k}$ is the squeezed parameter and ^$\hat{\mathcal{R}}_{k}(\theta_{k})$ is the rotation parameter that will be neglected since it only contributes the phase factor of wave function.
-Then, by applying $\hat{S}_{\vec{k}}(r_{k},\phi_{k})$ to the vacuum state, we can express the wave function of the two-mode squeezed state as follows
+$\hat{\mathcal{R}}_{k}(\theta_{k})=\exp\bigg[-i\theta_{k}(\eta)\big(\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}+\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}\big)\bigg]$ (27)
+rwhere $\hat{\mathcal{S}}_{k}$ is the squeeze operator and $\hat{\mathcal{R}}_{k}(\theta_{k})$ is the rotation operator, which can be neglected as it only contributes an overall phase factor to the wave function.
+Applying $\hat{S}_{{k}}(r_{k},\phi_{k})$ to the vacuum state then yields the wave function of the two-mode squeezed state:
 $|\psi\rangle=\hat{S}_{\vec{k}}(r_{k},\phi_{k})|0;0\rangle_{\vec{k},-\vec{k}}=\frac{1}{\cosh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}e^{2in\phi_{k}}\tanh^{n}r_{k}|n;n\rangle_{\vec{k},-\vec{k}}.$ (28)
-Being armed with this wave function, we could proceed the numeric of $\phi_{k}$ and $r_{k}$. But the evolution of $\phi_{k}$ and $r_{k}$ is based on the total action (19).
-4.2 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$
-In this work, we will evaluate the impact of potential for the Krylov complexity, where the Mukhanov-Sasaki variable contains the contribution of potential.
-Consequently, we cannot observe the explicit influence from potential.
-Therefore, we will directly perturb the total action into the second order,
-[…]
-[…]
+With this wave function, we can now compute the evolution of the squeezed state, which is governed by the parameters $r_{k}$ and $\phi_{k}$.
+The derivation of the two-mode squeezed state in Eq.
+(28) is model-independent, involving only the action of the squeeze operator on the vacuum.
+Model-specific information is encoded in the parameters $r_{k}$ and $\phi_{k}$, whose evolution is determined by the Hamiltonian.
+This will be confirmed in later sections.
+Furthermore, the two-mode quantum state $|n;n\rangle_{\vec{k},-\vec{k}}$ is a Fock state, and its form is given by
+$|n;n\rangle_{\vec{k},-\vec{k}}=\bigg[\frac{1}{n!}(c_{\vec{k}}^{\dagger}c_{-\vec{k}}^{\dagger})|0,0\rangle_{\vec{k},-\vec{k}}\bigg],$ (29)
+thus, it automatically satisfies both orthogonality and normalization conditions.
+Here, we need to specify the initial Krylov operator $\mathcal{O}_{0}$ in our setup.
+As mentioned, the wave function is a two-mode squeezed state (28); the initial operator must satisfy the Lanczos algorithm given in Eq.
+(5).
+Following Ref.
+Adhikari:2022oxr , we choose $|\mathcal{O}_{0})=|0,0\rangle_{\vec{k},-\vec{k}}$ as our initial operator.
+This choice naturally fulfills the requirements of the Lanczos algorithm, which will be verified in Sec.
+VI.
+IV.2 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$
+Numerical simulations of $\phi_{k}$ and $r_{k}$ require their evolution equations, which are derived from the Hamiltonian within the two-mode squeezed state formalism.
+This work focuses on evaluating the impact of the potential on Krylov complexity.
+The conventional Mukhanov-Sasaki variable is not ideal for this purpose, as it encapsulates the combined effect of the potential and other factors.
+To explicitly track the potential’s role, we instead perform a direct second-order perturbation of the total action.
+$S=\frac{1}{2}\int d\eta d^{3}x\Bigl[f^{\prime 2}-(\partial_{i}f)^{2}+(\frac{a^{\prime}}{a})^{2}f^{2}-2f^{\prime}f\frac{a^{\prime}}{a}-a^{2}V_{,\phi\phi}f^{2}\Bigr]$ (30)
+where $f(\eta,\vec{x})$ is the inflaton perturbation defined via $\phi(\eta,\vec{x})=\bar{\phi}(\eta)+\frac{f(\eta,\vec{x})}{a(\eta)}$, with $\frac{f(\eta,\vec{x})}{a(\eta)}\ll 1$ in Planck units and this perturbation theory will be violated in strong gravitational system.
+In the current investigations, we focus on the perturbation of inflaton, which is responsible for the formation of large-scale structure; for further details, we refer the reader to Baumann:2009ds .
+Under the slow-roll condition during inflation, we neglect the potential term, i.e., we assume $\left(\frac{a^{\prime}}{a}\right)^{2}f^{2}-2f^{\prime}f\frac{a^{\prime}}{a}\gg a^{2}V_{,\phi\phi}f^{2}.$ The Hamiltonian is then constructed as $H=\int d^{3}x\,d\eta\,(\pi f^{\prime}-\mathcal{L})$, where $\mathcal{L}$ denotes the Lagrangian and $\pi$ is the conjugate momentum defined by
 $\pi(\eta,\vec{x})=\frac{\delta L}{\delta f^{\prime}(\eta,\vec{x})}=f^{\prime}-\frac{a^{\prime}}{a}f.$ (31)
-According to Eq. (28) and Eq. (29), we could obtain
-[…]
+According to Eq. (30) and Eq. (31), we obtain
+$H=\frac{1}{2}\int d^{3}xd\eta[\pi^{2}+(\partial_{i}f)^{2}+\frac{a^{\prime}}{a}(\pi f+f\pi)+a^{2}V_{,\phi\phi}f^{2}].$ (32)
 Using the Fourier decomposition of operator $\hat{f}(\eta,\vec{x})$ and $\hat{\pi}(\eta,\vec{x})$,
@@
 $i\frac{d}{d\eta}|\psi\rangle=\hat{H}_{k}|\psi\rangle,$ (36)
-we could obtain the evolution of equation of $r_{k}$ and $\phi_{k}$
+we obtain the evolution of equation of $r_{k}$ and $\phi_{k}$
 $\begin{split}&r^{\prime}_{k}=\frac{a^{2}V_{,\phi\phi}}{2k}\sin{2\phi_{k}}-\frac{a^{\prime}}{a}\cos{2\phi_{k}},\\ &\phi_{k}^{{}^{\prime}}=-k-\frac{a^{2}V_{,\phi\phi}}{2k}+(\frac{a^{\prime}}{a}\sin{2\phi_{k}}+\frac{a^{2}V_{,\phi\phi}}{2k}\cos{2\phi_{k}})\coth{2r_{k}}.\end{split}$ (37)
 In inflation, Eq. (37) will nicely recover into the case of
-$\begin{split}&r_{k}^{{}^{\prime}}=-\frac{a^{\prime}}{a}\cos(2\phi_{k}),\\ &\phi_{k}^{{}^{\prime}}=-k+\frac{a^{\prime}}{a}\sin(2\phi_{k})\coth(2r_{k}).\end{split}$ (36)
-It is very difficult to obtain the numeric of Eq. (35), thus we could make a variable transformation, $y=\log_{10}a$, to numerically solve Eq. (35),
+$\begin{split}&r_{k}^{{}^{\prime}}=-\frac{a^{\prime}}{a}\cos(2\phi_{k}),\\ &\phi_{k}^{{}^{\prime}}=-k+\frac{a^{\prime}}{a}\sin(2\phi_{k})\coth(2r_{k}),\end{split}$ (38)
+where it is consistent with Bhattacharyya:2020rpy for $\frac{z^{\prime}}{z}=\frac{a^{\prime}}{a}$, since $\epsilon$ is constant.
+Directly solving Eq.
+(37) numerically is very difficult.
+Therefore, we introduce the variable transformation $y=\log_{10}a$ and solve Eq.
+(37) numerically in terms of $y$,
 $\begin{split}\frac{dr_{k}}{dy}=&\frac{\ln(10)10^{\frac{y}{2}(5+3\omega)}V_{,\phi\phi}}{2kH_{0}}\sin{2\phi_{k}}-\ln(10)\cos{2\phi_{k}}\\ \frac{d\phi_{k}}{dy}=&-\frac{\ln(10)10^{\frac{y}{2}(1+3\omega)}k}{H_{0}}-\frac{\ln(10)10^{\frac{y}{2}(5+3\omega)}V_{,\phi\phi}}{2kH_{0}}+\Bigl(\ln(10)\sin{2\phi_{k}}\\ &+\frac{\ln(10)10^{\frac{y}{2}(5+3\omega)}V_{,\phi\phi}}{2kH_{0}}\cos{2\phi_{k}}\Bigr)\coth{2r_{k}}\end{split}$ (39)
 where we have utilized Eq. (17) and $\frac{a^{\prime}}{a}=H_{0}a^{-\frac{1}{2}(1+3\omega)}$.
-Figure 3: The numerical solution of $r_{k}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD, $k=0.005$ at MD for simplicity.
+Figure 3: The numerical solution of $r_{k}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ for RD, $k=0.005$ for MD.
 The effect of potential is always ignored in slow-roll inflation.
 But in RD and MD periods, the impact of potential is taken into account.
-Before analyzing the figure of $r_{k}$, it’s important to note that the Krylov complexity depends only on $r_{k}$.
-Therefore, we only provide the numerical values of $r_{k}$ in Fig.
-3 based on Eq.
+Before analyzing the figure for $r_{k}$, it is important to note that the Krylov complexity depends solely on this quantity.
+Therefore, in Fig.
+3, we show only the numerical results for $r_{k}$ based on Eq.
 (39).
-In this equation, we have set $H_{0}=1$ for simplicity and used various values of $k$ in different time periods.
+For simplicity, we set $H_{0}=1$ and examine different values of the comoving wavenumber $k$ across various cosmological epochs.
 The first panel of Fig.
 3 illustrates the evolution of $r_{k}$ during inflation.
-There is an oscillation before the horizon exits, and then $r_{k}$ will grow linearly, consistent with the standard case of Li:2024kfm .
-As we transit to RD, we observe that $r_{k}$ will increase to some constant value that oscillates around this value.
-Interestingly, different potentials produce almost the same trend for $r_{k}$, indicating that the specific potential does not significantly impact the evolution of $r_{k}$.
-This observation also applies to MD, where the evolution trend is similar to RD.
-However, we notice that varying momentum significantly affects $r_{k}$, leading to oscillations as the value of $k$ decreases.
-Although we cannot distinguish between the various potentials using Fig.
-3, our numerical data still demonstrates that the evolution of $r_{k}$ is significantly different whether there is a potential or not.
-It’s worth noting that we cannot provide the values of $r_{k}$ without a potential, as this is not physically meaningful.
-The potential cannot be ignored due to the violation of slow-roll conditions.
-In this section, we have analyzed the evolution of $r_{k}$ for different inflationary potentials within a single-field inflationary framework.
-Our research indicates that the value of $r_{k}$ increases during inflation.
-The behavior of $r_{k}$ in RD and MD eras is similar and will oscillate around constant values.
-Our numerical analysis also shows that various inflationary models produce a similar trend for $r_{k}$.
-Additionally, we have observed that the scale factor and momentum have a significant impact on the evolution of $r_{k}$.
-All of the details are encoded in Fig.
+Prior to horizon exit, $r_{k}$ exhibits oscillatory behavior, after which it grows linearly, consistent with the standard result reported in Li:2024kfm .
+Following the transition to the RD era, $r_{k}$ rises and then saturates, fluctuating around a nearly constant value.
+Notably, different inflationary potentials yield nearly identical trends for $r_{k}$, indicating that the specific form of the potential does not significantly influence its evolution.
+A similar trend is observed during the MD era, where the behavior of $r_{k}$ closely resembles that in the RD era.
+However, we find that the value of the wavenumber $k$ has a pronounced effect: as $k$ decreases, $r_{k}$ displays increasingly prominent oscillations.
+We note that presenting results in RD and MD, the contribution of potential is taken into account since the violation of slow roll condtions.
+In this section, we have analyzed the evolution of $r_{k}$ for different inflationary potentials within the single-field inflationary model.
+Our results indicate that $r_{k}$ grows during the inflationary epoch.
+In both the RD and MD eras, $r_{k}$ exhibits a similar qualitative behavior, saturating and then oscillating around an approximately constant value.
+Numerical analysis further shows that various inflationary models yield qualitatively similar trends for $r_{k}$.
+Additionally, we find that both the scale factor and the comoving wavenumber significantly influence the evolution of $r_{k}$.
+These features are summarized visually in Fig.
 3.
-5 Krylov complexity in closed system
-In this section, we will investigate the Krylov complexity in the method of closed system Liu:2021nzx ; Li:2021kfq ; Li:2023ekd , consisting of the Krylov complexity and Krylov entropy.
-5.1 Krylov complexity
-For obtaining the Krylov complexity, the most essential quantity is the wave function. First, the two-mode squeezed state wave function can be represented by Krylov basis as follows,
-$\mathcal{O}(t)=\sum_{n=0}^{\infty}(i)^{n}\phi_{n}(t)|\mathcal{O}_{n})=\frac{1}{\cosh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}e^{2in\phi_{k}}\tanh^{n}r_{k}|n;n\rangle_{\vec{k},-\vec{k}}.$ (38)
-Then, we could find the explicit wave function $\phi_{n}$
+V Krylov complexity in closed-system methodology
+In this section, we conduct numerical simulations of Krylov complexity using the closed-system methodology Liu:2021nzx ; Li:2021kfq ; Li:2023ekd .
+This framework encompasses both the Krylov complexity and the Krylov entropy.
+The Krylov complexity is determined by the coefficients $r_{k}$ and the phases $\phi_{k}$, which are in turn governed by the Hamiltonian operator given in Eq.
+(35).
+V.1 Krylov complexity
+To obtain the Krylov complexity, the key quantity is the wave function. Specifically, the two-mode squeezed state wave function is expressed in the Krylov basis as follows,
+$\mathcal{O}(\eta)=\sum_{n=0}^{\infty}(i)^{n}\phi_{n}(\eta)|\mathcal{O}_{n})=\frac{1}{\cosh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}e^{2in\phi_{k}}\tanh^{n}r_{k}|n;n\rangle_{\vec{k},-\vec{k}},$ (40)
+where $\mathcal{O}(\eta)$ is the operator of wave function whose original definition can be found in Eq. (8). Then, we find the explicit wave function $\phi_{n}$
 $\phi_{n}=\frac{1}{\cosh{r_{k}}}(-e^{2i\phi_{k}}\tanh{r_{k}})^{n}.$ (41)
-We can define the Krylov compplexity in terms of $\phi_{n}$
+We define the Krylov compplexity in terms of $\phi_{n}$
 $K=\sum_{n}n\left|\phi_{n}\right|^{2}\\ =\cosh^{2}r_{k}\tanh^{2}r_{k}=\sinh^{2}r_{k}.$ (42)
-It is explicitly showing that the Krylov complexity is independent of $\phi_{k}$, that is why we only list the figure of $r_{k}$.
+This result explicitly demonstrates the independence of the Krylov complexity from $\phi_{k}$; hence, only $r_{k}$ is plotted.
 Figure 4: The numerical of Krylov complexity in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.
-Being armed with Eq.
-(40), we could give its numeric in Fig.
-4, indicating the evolution of Krylov complexity in inflation, RD, and MD.
+Based on Eq.
+(42), we show its numerical evolution in Fig.
+4, which depicts the behavior of the Krylov complexity throughout the inflationary, RD, and MD eras.
 Fig.
-4 clearly shows that the Krylov complexity is enhanced in inflation which is consistent with the standard case of the modified dispersion within the framework Li:2024kfm .
-Similar work investigated the Krylov complexity for the exponential expansion for de-Sitter spacetime corresponding to the inflation Adhikari:2022oxr .
-Ref.
-Adhikari:2022oxr mentioned that the Krylov complexity equals the average particle number proportional to the volume.
-Thus we could see that this simple analysis is consistent with the CV conjuncture Stanford:2014jda , in which the complexity is equaled to the volume of wormhole under the framework of holography.
-However, the method of Adhikari:2022oxr is for the closed system.
-In the case of RD and MD, it is observed that the Krylov complexity will increase to certain constant values that fluctuate around this value.
-Additionally, different patterns of inflationary models tend to produce similar trends of Krylov complexity in different periods.
-Our analysis aligns with the findings of Barbon:2019wsy , where the Krylov complexity will saturate some constant values even after the scrambling time in a closed system.
-The scrambling time will be assessed in the context of an open system, particularly with regard to the Lanczos coefficient, Lyapunov index, $\it e.t.c$.
-5.2 Krylov Entropy
-In a quantum system, entropy is defined as a measure of the disorder of a system.
-The Krylov entropy (K-entropy) can be used to test the level of disorder of curvature perturbation.
-Following Barbon:2019wsy , the K-entropy has the following form
+4 clearly reveals a significant growth in the Krylov complexity during the inflationary phase, a result consistent with the modified dispersion case studied in Li:2024kfm .
+This finding is also in agreement with related work exploring Krylov complexity under exponential (de Sitter) expansion corresponding to inflation Adhikari:2022oxr .
+We note that Ref.
+Adhikari:2022oxr identifies the Krylov complexity with the average particle number, which scales proportionally with volume.
+However, we emphasize that the approach adopted in Adhikari:2022oxr is specifically applicable to a closed system.
+During the RD and MD periods, the Krylov complexity is observed to increase toward certain constant values, around which it subsequently fluctuates.
+Furthermore, different inflationary models exhibit broadly similar trends in the evolution of Krylov complexity across the respective epochs.
+Our results are consistent with the analysis in Barbon:2019wsy , which found that in a closed system the Krylov complexity saturates to approximate constant values even after the scrambling time.
+The scrambling time itself will be analyzed in the open-system framework, specifically in relation to the Lanczos coefficients, the Lyapunov exponent, among other quantities.
+V.2 Krylov Entropy
+In quantum systems, entropy is defined as a measure of a system’s disorder.
+The Krylov entropy (K-entropy) serves to quantify the degree of disorder in curvature perturbations.
+Following Barbon:2019wsy , the K-entropy is defined as follows,
 $\begin{split}S_{K}&=-\sum_{n=0}^{\infty}\left|\phi_{n}\right|^{2}\ln\left|\phi_{n}\right|^{2}\\ &=\cosh^{2}r_{k}\ln(\cosh^{2}r_{k})-\sinh^{2}r_{k}\ln(\sinh^{2}r_{k})\end{split}$ (43)
-where $\phi_{n}$ is the wave function (27). Once given the analytic formula of K-entropy, we could numerically simulate it in various epochs as shown in Fig. 5.
+where $\phi_{n}$ is the wave function (28). Once given the analytic formula of K-entropy, we numerically simulate it in various epochs as shown in Fig. 5.
 Figure 5: The numerical of Krylov entropy in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.
 Fig.
-5 indicates a complete evolution of K-entropy in inflation, RD and MD, which is similar to Krylov complexity.
-We could see that the curvature of perturbation will be more and more disorder in inflation.
-For RD and MD, the disorder will be enhancing and then saturates to some constant values and the total trend is enhancing, where the trend is also highly related the scale factor.
-From another aspect, the preheating will generate the particles occurred at MD that leads to the growth of K-entropy.
-Let’s discuss Section 5.
-In this section, we provide a comprehensive analysis of Krylov complexity throughout the early universe.
-Our numerical data clearly indicates that different potentials yield almost identical patterns of Krylov complexity and K-entropy.
-Specifically, Krylov complexity increases during inflation and then stabilizes at constant values during RD and MD.
-This same behavior is observed for K-entropy.
-Additionally, we must examine the Lanczos coefficient (equivalent to the Lyapunov index in maximal chaos), which remains consistent in both the closed and open system methods.
-6 Krylov complexity with the approach of open system
-The preheating process in MD and RD is a highly non-equilibrium process that creates particles Kofman:1997yn ; Shtanov:1994ce .
-In addition, Ref.
-Cheung:2007st demonstrated that the universe’s evolution breaks the time invariance when the temporal direction is reversed, leading to a violation of energy conservation.
-Due to these reasons, the open system approach is considered to be more realistic compared to the closed system method.
-We will follow the method of Bhattacharya:2022gbz to investigate the Krylov complexity.
-6.1 Lanczos coefficient and dissipation coefficient
-Firstly, the general operator in Heisenberg’s picture has the following form,
-$\mathcal{O}(\eta)=e^{i\mathcal{L}_{o}\eta}$ (42)
-where the Lindbladian $\mathcal{L}_{o}$, also known as the Hamiltonian, acts on the Krylov basis, we can obtain
-$\mathcal{L}_{o}|\mathcal{O}_{n})=-ic_{n}|\mathcal{O}_{n})+b_{n+1}|\mathcal{O}_{n+1})+b_{n}|\mathcal{O}_{n-1})$ (43)
-where $c_{n}$ encodes the information of the open system and $b_{n}$ is the Lanczos coefficient.
-Due to this extra term, we could divide the total Hamiltonian (also known as Lindbladian) into two parts which can be represented by
+5 illustrates the complete evolutionary profile of the K-entropy across the inflationary, RD, and MD eras, exhibiting a trend similar to that of the Krylov complexity.
+It can be observed that the curvature perturbations become progressively more disordered during inflation.
+In the subsequent RD and MD phases, the degree of disorder initially increases before saturating to approximately constant values; nonetheless, the overall trend remains one of growth, which is closely tied to the evolution of the scale factor.
+From another perspective, preheating, which generates particles during the MD era, also contributes to the observed increase in the K-entropy.
+We now summarize the findings of Section V.
+In this section, we have conducted a comprehensive analysis of Krylov complexity in the early universe.
+Our numerical results clearly demonstrate that different inflationary potentials produce nearly identical patterns of evolution for both the Krylov complexity and the K-entropy.
+Specifically, the Krylov complexity grows during inflation and subsequently saturates to nearly constant values throughout the RD and MD eras.
+The K-entropy exhibits the same qualitative behavior.
+To characterize the chaotic features of the early universe more precisely, we must examine the Lanczos coefficients, which, in maximal chaos, play a role analogous to the Lyapunov exponent.
+These coefficients are identical in form for both the closed- and open-system approaches.
+In the following, we will employ the open-system methodology to investigate Krylov complexity.
+VI Krylov complexity with open-system methodology
+The preheating process during the MD and RD eras is a highly non-equilibrium phenomenon that generates particles Kofman:1997yn ; Shtanov:1994ce .
+Furthermore, Ref.
+Cheung:2007st demonstrated that the evolution of the universe breaks time-reversal invariance, leading to a violation of energy conservation.
+For these reasons, the open-system methodology is regarded as more realistic than the closed-system approach.
+Several open-system methodologies exist for investigating Krylov complexity.
+Here, we will implement the generalization of the Lanczos algorithm to the Arnoldi iteration via the Heissenberg form Bhattacharjee:2022lzy ; Bhattacharya:2022gbz ; Liu:2022god ; Nizami:2023dkf ; Nandy:2024evd .
+Let us review this Arnoldi iteration step by step.
+We begin by recalling the general Lindblad master equation Lindblad:1975ef :
+$\dot{\rho}=-i[H,\rho]+\sum_{k}[L_{k}\rho L_{k}^{\dagger}-\frac{1}{2}\{L_{k}^{\dagger}L_{k},\rho\}],$ (44)
+where $\rho$ is the density matrix, $H$ is the Hamiltonian, and $L_{k}^{\dagger}$ denotes the jump operator (which in our context corresponds to the annihilation operator).
+Following [81], we consider the open-system dynamics in an infinite-temperature environment.
+Specifically, we identify the system state $\rho$ with the maximally mixed state (infinite temperature state) $\rho_{\infty}$ due to the extremal high temperature for the early universe.
+This assumption defines the inner product for the operator space as $(A,B)=\text{Tr}(\rho_{\infty}A^{\dagger}B)$, which is central to our Krylov complexity calculations.
+Under this approximation, we obtain
+$\sum_{k}[L_{k}\rho_{\infty}L_{k}^{\dagger}-\frac{1}{2}\{L_{k}^{\dagger}L_{k},\rho_{\infty}\}]=0,$ (45)
+In this stationary state, it is reasonable to assume that the operator dynamics is governed by
+$\mathcal{\hat{O}}(t)=\exp(i\mathcal{L}t)\mathcal{\hat{O}}$, (46)
+where the Heisenberg-picture Lindbladian $\mathcal{L}$ is given by Bhattacharya:2022gbz ,
+$\mathcal{L}=\mathcal{L}_{H}+\mathcal{L}_{D},~~\mathcal{L}_{H}\mathcal{\hat{O}}=[H,\mathcal{\hat{O}}],~~\mathcal{L}_{D}=\sum_{k}[L_{k}\mathcal{\hat{O}}L_{k}^{\dagger}-\frac{1}{2}\{L_{k}^{\dagger}L_{k},\mathcal{\hat{O}}\}$] (47)
+where $\mathcal{L}_{H}$ governs the Hermitian dynamics, $\mathcal{L}_{D}$ denotes the dissipative part, and $\mathcal{L}$ behaves as a superoperator on the operator space.
+The inner product is defined as $(A|B):=\mathrm{Tr}[\rho_{\infty}A^{\dagger}B]$, which explicitly connects the operator space structure to the Krylov basis,
+$\rm span(\mathcal{O}_{0},...,\mathcal{O}_{n})=span(\mathcal{V}_{0},...,\mathcal{V}_{n}),$ (48)
+where $\mathcal{V}_{n}$ denotes the original Krylov basis, which is generated via the Arnoldi iteration.
+In this work, we follow the standard construction, beginning with an initial normalized operator $\mathcal{O}_{0}\propto\mathcal{O}$.
+Subsequent Krylov basis vectors are then generated by repeatedly applying the Liouvillian superoperator and performing orthonormalization at each step.
+To compute the physical quantity, an appropriate basis is required.
+Following Bhattacharya:2022gbz , we adopt an orthonormal basis $\{\mathcal{O}_{0},\ldots,\mathcal{O}_{n},\ldots\}$ generated by the open-system Lindbladian:
+$\rm span(\mathcal{O}_{0},...,\mathcal{O}_{n})=span(\mathcal{V}_{0},...,\mathcal{V}_{n}),$ (49)
+where $\mathcal{V}_{n}$ is the original Krylov basis, and we generate this basis by using the Arnoldi iteration, which is started with an initial normalized vector $\mathcal{O}_{0}\propto\mathcal{O}$.
+Thereafter, we construct the subsequent Krylov basis vectors by repeating the following steps:
+$|\mathcal{U}_{k})=\mathcal{L}|\mathcal{O}_{k-1}).$ (50)
+From $j=0$ to n1n-1, we perfom the following iterations: $1.~h_{j,k-1}=(\mathcal{O}_{j}|\mathcal{U}_{k})$, $2.~|\tilde{\mathcal{U}}_{k})=|\mathcal{U}_{k})-\sum_{j=0}^{k-1}h_{j,k-1}|\mathcal{O}_{j})$, $3.~h_{k,k-1}=\sqrt{(\tilde{U}_{k}|\tilde{\mathcal{U}}_{k})}$, it will stop as if $h_{k,k-1}=0$, otherwise, it is defined as follows,
+$|\mathcal{O}_{k})=\frac{|\tilde{\mathcal{U}}_{k})}{h_{k,k-1}}.$ (51)
+Finally, the Lindbladian can transform into an upper Heissenberg form in the span Krylov basis:
+$\mathcal{L}=\begin{pmatrix}h_{0,0}&h_{0,1}&h_{0,2}&\cdots&\cdots&h_{0,n}\\ h_{1,0}&h_{1,1}&h_{1,2}&\cdots&\cdots&h_{1,n}\\ 0&h_{2,1}&h_{2,2}&h_{2,3}&\cdots&\cdots\\ \vdots&0&h_{3,2}&\cdots&\cdots&\cdots\\ 0&\vdots&0&\vdots&\ddots&h_{n-1,n}\\ 0&0&\vdots&0&h_{n,n-1}&h_{n,n}\end{pmatrix}$ (52)
+where $h_{m,n}=(\mathcal{O}_{m}|\mathcal{L}|\mathcal{O}_{n})$.
+As $\mathcal{L}$ is a Hermitian superoperator, the Arnoldi iteration will reduce into the Lanczos algorithm.
+Following conventions of Li:2024kfm , we represent the generalized Lanczos algorithm as
+$\mathcal{L}|\mathcal{O}_{n})=-ic_{n}|\mathcal{O}_{n})+b_{n+1}|\mathcal{O}_{n+1})+b_{n}|\mathcal{O}_{n-1}),$ (53)
+where the coefficient $c_{n}$ encodes the information of the open quantum system dynamics for a given Hamiltonian, corresponding to the diagonal elements $h_{n,n}$ in Eq.
+(52).
+Meanwhile, the Lanczos coefficients $b_{n}$, associated with the off-diagonal elements $h_{n,n-1}$ in Eq.
+(52), characterize the chaotic behavior of the dynamical system.
+The orthogonal basis $\mathcal{O}_{n}=|n,n\rangle_{-\vec{k},\vec{k}}$ spans the Krylov space defined by Eq.
+(49).
+The Liouvillian superoperator $\mathcal{L}$ is equivalent to a Hamiltonian operator that admits a representation in terms of creation and annihilation operators.
+VI.1 Lanczos coefficient and dissipation coefficient
+Based on the foregoing analysis, we separate the total Hamiltonian (the Lindbladian) into two constituent parts:
 $\mathcal{L}_{o}=\mathcal{H}_{k}=\mathcal{H}_{close}+\mathcal{H}_{open}$ (54)
-In our work, $\mathcal{L}_{o}$ is the Hamiltonian Eq. (33). In light of (43), we can easily obtain the open part (generating $c_{n}$) of Hamiltonian,
+In light of (53), we easily obtain the open part (generating $c_{n}$) of Hamiltonian,
 $\mathcal{H}_{open}=(k+\frac{a^{2}V_{,\phi\phi}}{2k})(\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}+\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger})$ (55)
@@
 $\mathcal{H}_{close}=(\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a})\hat{c}_{\vec{k}}^{\dagger}\hat{c}_{-\vec{k}}^{\dagger}+(\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a})\hat{c}_{\vec{k}}\hat{c}_{-\vec{k}}$ (56)
-where $|\mathcal{O}_{n})=\Bigl{|}n_{\vec{k}};n_{-\vec{k}}\Bigr{\rangle}$. By substituting Eq. (45) and Eq. (46) into Eq. (43), we can obtain the equations for $c_{n}$ and $b_{n}$ as
+where $|\mathcal{O}_{n})=\Bigl|n_{\vec{k}};n_{-\vec{k}}\Bigr\rangle$.
+Specifically, the jump operators $L_{k}$ appearing in the dissipative term $L_{D}$ are identified with the annihilation and creation operators in our work, i.e., $L_{k}=c_{k}$ and $L_{-k}=c_{-k}^{\dagger}$.
+By substituting Eq.
+(55) and Eq.
+(56) into Eq.
+(53), we obtain the equations for $c_{n}$ and $b_{n}$ as
 $c_{n}=i(2n+1)(k+\frac{a^{2}V_{,\phi\phi}}{2k}).$ (57)
 $b_{n}=n\sqrt{[(\frac{a^{2}V_{,\phi\phi}}{2k})^{2}+(\frac{a^{\prime}}{a})^{2}]}.$ (58)
-Here, we will discuss more about $b_{n}$ and $c_{n}$.
-Ref.
-Bhattacharjee:2022lzy has introduced the $a_{n}$ corresponding to $c_{n}$ in our work and also stated that $c_{n}$ is highly related to the dissipative coefficient.
-Meanwhile, we could see that $b_{n}$ is the same for the closed system and open system since it is precisely determined by the Hamiltonian.
+Here, we provide further discussion regarding the coefficients $b_{n}$ and $c_{n}$.
 In Ref.
-Parker:2018yvk , it has clearly indicated that the dynamical system is an infinite, non-integrable, many-body system as $b_{n}=\alpha n+\gamma$, where $\alpha$ and $\gamma$ are two parameters determined by various models.
-In our case, we could see that $b_{n}\propto n$, thus the whole early universe is an ideal chaotic system.
-6.1.1 $b_{n}$ and $\mu_{2}$
-[…]
+Bhattacharjee:2022lzy , the coefficient denoted by $a_{n}$ corresponds to our $c_{n}$, and it is also established that $c_{n}$ is associated with the dissipation coefficient.
+Meanwhile, we observe that $b_{n}$ remains identical for both the closed and open systems, as it is determined solely by the Hamiltonian.
+In Ref.
+Parker:2018yvk , it is clearly established that the dynamical system is an infinite, chaotic, many-body system as $b_{n}=\alpha n+\gamma$, where $\alpha$ and $\gamma$ are parameters that depend on the specific model.
+In our case, since $b_{n}\propto n$, the early universe can be regarded as an ideal chaotic system.
+VI.1.1 $b_{n}$ and $\mu_{2}$
+In this work, we adopt the notation of Ref.
+Li:2024kfm to perform numerical simulations of the coefficients $b_{n}$ and $\mu_{2}$.
+We begin by examining the properties of $c_{n}$ within our framework.
+Recall that Eq.
+(9) (which is essentially the Schrödinger equation) involves only the Lanczos coefficients.
+The parameter $c_{n}$ will subsequently appear in the following modified equation:
 $\partial_{\eta}\phi+2b_{n}\partial_{n}\phi+\tilde{c}_{n}\phi=0,$ (59)
-where we have defined $\tilde{c}_{n}=-ic_{n}$, meanwhile we have used the $b_{n+1}\approx b_{n}$ and $\phi_{n+1}-\phi_{n-1}\approx 2\partial_{n}\phi$ in the continous limit.
-$b_{n}$ and $c_{n}$ are defined in Eq.
+Here, we have defined $\tilde{c}_{n}=-ic_{n}$.
+In taking the continuous limit, we have employed the approximations $b_{n+1}\approx b_{n}$ and $\phi_{n+1}-\phi_{n-1}\approx 2\partial_{n}\phi$.
+The coefficients $b_{n}$ and $c_{n}$ are defined in Eq.
 (58) and Eq.
 (57), respectively.
-The generic formulas of $b_{n}$ and $c_{n}$ are
+Their general expressions are given by
 $b_{n}^{2}=|1-\mu_{1}^{2}|n(n-1+\beta),\ \ c_{n}=i\mu_{2}(2n+\beta)$ (60)
-where $\mu_{1}$, $\mu_{2}$ and $\beta$ are determined by various models and $\mu_{2}$ plays a role of dissipative coefficient as shown in Ref.
+where $\mu_{1}$, $\mu_{2}$, $\beta$ are determined by various models and $\mu_{2}$ plays a role of dissipative coefficient as shown in Ref.
 Li:2024kfm .
@@
 $|1-\mu_{1}^{2}|=(\frac{a^{2}V_{,\phi\phi}}{2k})^{2}+(\frac{a^{\prime}}{a})^{2},\mu_{2}=k+\frac{a^{2}V_{,\phi\phi}}{2k}.$ (61)
-It should be noticed that $\mu_{2}$ is mainly determined by $k$ in inflation due to the slow-roll conditions.
-Once entering RD and MD (taking the contribution of the effective mass into account), we could see that it will be smaller than unity since the effective mass of these three types potentials is much more tiny compared with momentum $k$.
-First, we discuss the Lanczos coefficient $b_{n}$.
-Fig.
-6 clearly indicates the three distinctive potentials of $b_{n}$ in inflation, RD and MD, in which the parameters of $k$, $H_{0}$ are set to be the same with the previous plots.
-As Bhattacharya:2022gbz confirmed, the Lyapunov index is equavillent to $b_{n}$ whose relation is $\lambda=2\alpha$ with $b_{n}=\alpha n+\delta$ ($\delta=0$ in our case), where it is explicitly showing that $\lambda=2\alpha$, namely $\lambda=b_{n}$ with $n=2$ which will be shown in Fig.
-6.
-The physical meaning of $\lambda$ represents the chaotic feature of dynamical system.
-Consequently, $b_{n}$ in our case also manifest the chaotic feature for the early universe.
+It should be noted that during inflation, the parameter $\mu_{2}$ is primarily determined by the wavenumber $k$ due to the slow-roll conditions.
+Upon entering the RD and MD eras, which is taking into account the effective mass, we observe that $\mu_{2}$ becomes smaller than unity, because the effective mass associated with these three types of potentials is much smaller than the momentum $k$.
+We first discuss the Lanczos coefficient $b_{n}$.
+Figure 6 clearly illustrates the three distinct behaviors of $b_{n}$ during inflation, RD, and MD, where the parameters $k$ and $H_{0}$ are set to the same values as in the previous plots.
+As established in Ref.
+Bhattacharya:2022gbz , the Lyapunov exponent $\lambda$ is related to $b_{n}$ by $\lambda=2\alpha$, given the linear form $b_{n}=\alpha n+\delta$ (with $\delta=0$ in our case).
+This explicitly shows that $\lambda=2\alpha$, or equivalently $\lambda=b_{n}$ for $n=2$, as will be demonstrated in Figure 6.
+Physically, $\lambda$ characterizes the chaotic nature of the dynamical system.
+Consequently, in our framework, $b_{n}$ also reflects the chaotic behavior of the early universe.
 Figure 6: The plot of $b_{n}=\lambda$ with $n=2$.The numerical of $b_{n}$ in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.
-Figure 6 illustrates the evolution trend $b_{n}$ in inflation, radiation domination (RD), and matter domination (MD).
-Based on the discussion regarding $b_{n}$, chaos will be heightened due to the exponential expansion of the background space.
-As the energy decreases, $b_{n}$ will also decline since the temperature decreases.
-During MD, particles are generated through a preheating process known as parametric resonance, which leads to an increase in $b_{n}$.
-Finally, we need to analyze the dissipative coefficient $\mu_{2}$ that is related to $c_{n}$, where we could confirm the previous results via Li:2024kfm .
-Fig.
-7 clearly shows the evolution in RD and MD.
-The $\mu_{2}$ in inflation is not shown in Fig.
-7 since $\mu_{2}=k=1$ in light of the values of $k$ for inflation.
-Ref.
-Li:2024kfm claimed that inflation is a strong dissipative system as $\mu_{2}\geq 1$, RD and MD are weak dissipative systems as $\mu_{2}\ll 1$, which is explicitly confirmed by Fig.
+Figure 6 illustrates the evolution of the coefficient $b_{n}$ during inflation, RD, and MD.
+Based on the analysis of $b_{n}$, the degree of chaos increases due to the exponential expansion of the background spacetime.
+As the energy scale decreases, $b_{n}$ decreases correspondingly with the temperature.
+During MD, particles are produced via the preheating mechanism known as parametric resonance, leading to an increase in $b_{n}$.
+Finally, we analyze the dissipation coefficient $\mu_{2}$, which is related to $c_{n}$, thereby confirming previous results from Ref.
+Li:2024kfm .
+Figure 7 clearly shows the evolution of $\mu_{2}$ in RD and MD.
+The value of $\mu_{2}$ during inflation is not shown in Fig.
+7 because $\mu_{2}=k=1$ for the chosen wavenumber $k$ in the inflationary scenario.
+As asserted in Ref.
+Li:2024kfm , inflation constitutes a strong dissipative system with $\mu_{2}\geq 1$, while RD and MD are weak dissipative systems with $\mu_{2}\ll 1$, a result explicitly corroborated by Fig.
 7.
@@
 The value of $\mu_{2}$ indicates the RD and MD are weak dissipative systems.
-6.2 General discussion of $\phi_{n}$
-Before the Krylov complexity, we discuss the general properties of $\phi_{n}$ since the definition of Krylov complexity is explicitly represented by $C_{K}=\sum_{n=1}^{+\infty}n|\phi_{n}|^{2}$.
+VI.2 General discussion of $\phi_{n}$
+Before investigating the Krylov complexity, we discuss the general properties of $\phi_{n}$ since the definition of Krylov complexity is explicitly represented by $C_{K}=\sum_{n=1}^{+\infty}n|\phi_{n}|^{2}$.
 We will follow the general discussion of $\phi_{n}$ via Refs.
@@
 Eq.
-(49) can be rewritten by
+(59) is rewritten by
 $\partial_{\eta}\phi+n(\chi\mu\phi+2\alpha\partial_{n}\phi)=0,$ (62)
@@
 It is straightforward to demonstrate that $\exp(2a\eta_{*})=\xi$.
-After some simple algebra, we could obtain the formula $\xi$ in our case as follows,
+After some simple algebra, we obtain the formula $\xi$ in our case as follows,
 $\xi=\frac{2\sqrt{(\frac{a^{2}V_{,\phi\phi}}{2k})^{2}+(\frac{a^{\prime}}{a})^{2}}}{(2+\frac{1}{n})(k+\frac{a^{2}V_{,\phi\phi}}{2k})}$ (63)
-From this formula, we can observe that $\xi$ is not strictly constant; it varies with respect to $a$ and $k$.
+From this formula, we observe that $\xi$ is not strictly constant; it varies with respect to $a$ and $k$.
 Consequently, a significant change in the background will have a substantial impact on Krylov complexity.
 In this section, we examined the general properties of Krylov complexity without providing the precise formula for $\phi_{n}$.
-To fully capture the information concerning Krylov complexity within the framework of the open system method, it is essential to have the exact wave function for $\phi_{n}$.
-6.3 Wave function within the method of open system
-Although we constructed the wave function for an open system in reference Li:2024kfm , based on the framework established in Bhattacharjee:2022lzy , our approach is too simplistic to capture the complete information of the open system.
-Notably, our wave function, as shown in Li:2024kfm , can effectively reduce to the case of a closed system, establishing a correspondence between $\eta$ and $r_{k}$.
+To fully capture the information concerning Krylov complexity within the framework of the open-system methodology, it is essential to have the exact wave function for $\phi_{n}$.
+VI.3 Wave function within open-system methodology
+Although we constructed the wave function for an open system in reference Li:2024kfm , based on the framework established in Bhattacharjee:2022lzy , this wave function cannot capture the complete information since it is the function of the conformal time $\eta$.
 To clearly illustrate this issue, we will first present our constructed wave function as follows:
 $\phi_{n}=\frac{\rm sech\eta}{1+\mu_{2}\tanh\eta}|1-\mu_{1}^{2}|^{\frac{n}{2}}(\frac{\tanh\eta}{1+\mu_{2}\tanh\eta})^{n}.$ (64)
-To perform the Taylor expansion around $\mu_{2}$ to the first order, we obtain the expression
-$\phi_{n}\approx|1-\mu_{1}^{2}|^{n/2}\,\text{sech}(\eta)\,\tanh^{n}(\eta)+\mathcal{O}(\mu_{2}^{n}).$
-Next, we can neglect the term involving $\mu_{1}^{2}$, which enables us to find that this approximation matches the wave function given in equation (39) for $\phi_{k}=0$.
-Additionally, we observe the correspondence $\eta\sim r_{k}$.
-In principle, the wave function $\phi_{n}$ should incorporate the information of both $r_{k}$ and $\phi_{k}$ to accurately reproduce the wave function represented in (39) as the leading order of the open system’s method.
-To achieve this, we will follow the guidelines provided in Appendix D of Parker:2018yvk .
+To perform the Taylor expansion around $\mu_{2}$ to the leading order $(\mu_{2}=0)$, we obtain the expression
+$\phi_{n}\approx\text{sech}(\eta)\,\tanh^{n}(\eta)+\mathcal{O}(\mu_{2}^{n}).$
+To relate with Eq, (41), we can find one explicit correspondence between $r_{k}$ and $\eta$.
+To include the parameter of $\phi_{k}$, we will follow the guidelines provided in Appendix D of Parker:2018yvk to rigourisly derive the so-called open two-mode squeezed state, which will accurately reproduce the wave function represented in (41).
 For deriving the two-mode squeezed state in open system, we first introduce the concept of orthogonal polynomial sequence (OPS) $\{P_{n}(x)\}_{n\geq 0}$, it satisfied with the following recurrence relation,
@@
 This determinant encodes information about $b_{n}$ as represented by the corresponding matrix.
-Based on the definition of $L$, we can express it in terms of $x^{n}$ or $e_{n}$, which will be used to derive the wave function using the open system method.
-To obtain the similar recurrence relation as (43), we introduce the the natural orthonormal basis $\{e_{n}\}$, and meanwhile $P_{n}$ can be represented in terms of $e_{n}$
+Based on the definition of $L$, we express it in terms of $x^{n}$ or $e_{n}$, which will be used to derive the wave function using the open-system methodology.
+To obtain the similar recurrence relation as (53), we introduce the the natural orthonormal basis $\{e_{n}\}$, and meanwhile $P_{n}$ is represented in terms of $e_{n}$
 $P_{n}(x)=\Bigl(\prod_{k=1}^{n}b_{k}\Bigr)e_{n},\ \ \ \ \mbox{and}\ \ \ x^{n}=\mathcal{L}^{n}e_{0},$ (66)
 where $x$ is the parameter of $P_{n}$, $\mathcal{L}$ represents our Hamiltonian expressed in terms of creation and annihilation operators, and $|e_{n}\rangle\equiv|\mathcal{O}_{n}\rangle=\Bigl|n_{\vec{k}};n_{-\vec{k}}\Bigr\rangle$.
-By combining Equation (55) with Equation (56), we can derive the following recurrence relation.
+By combining Equation (65) with Equation (66), we derive the following recurrence relation.
 $b_{n+1}e_{n+1}+b_{n}e_{n-1}=(x-\tilde{c}_{n})e_{n}.$ (67)
@@
 where $\mu_{n,k}$ refers to a triangular linear transform with matrix elements that relate $x$ to $P_{n}$.
-This formula demonstrates that the polynomials $P_{n}$ can be expressed in terms of $x$.
+This formula demonstrates that the polynomials $P_{n}$ is expressed in terms of $x$.
 All three orthogonal polynomial sequences (OPS) mentioned are examples of Sheffer orthogonal polynomials, which are defined as a set of polynomials $\{P_{n}(x)\}_{n\geq 0}$ characterized by a generating function,
@@
 $\sum_{n=0}^{\infty}Q_{n}(x)\frac{(i\eta)^{n}}{n!}=\frac{\rm sech(\eta)^{\eta_{1}}}{(1+\mu_{2}\tanh(\eta))^{\eta_{1}}}\exp(x\frac{i\tanh(\eta)}{1+\mu_{2}\tanh(\eta)}).$ (78)
-Based on Eqs. (56) and (59), we could obtain
+Based on Eqs. (66) and (69), we obtain
 $(e_{d}|\mathcal{L}^{n}|e_{0})=\mu_{n,d}\prod_{k=1}^{n}b_{k}.$ (79)
-Furthermore, we could have
+Furthermore, we have
 $(e_{d}|e^{i\mathcal{L}\eta}|e_{0})=\prod_{k=1}^{d}b_{k}\sum_{n=0}^{\infty}\frac{(i\eta)^{n}}{n!}\mu_{n,d}.$ (80)
-In light of (59), we could obtain
+In light of (69), we obtain
 $\sum_{n=0}^{\infty}\frac{(i\eta)^{n}}{n!}\mu_{n,d}=\frac{\rm sech\eta}{1+\mu_{2}\tanh\eta}\frac{1}{d!}\frac{(i\tanh\eta)^{d}}{(1+\mu_{2}\tanh\eta)^{d}},$ (81)
@@
 In our calculations, we multiply a phase factor $ie^{2i\phi_{k}(\eta)}$ which does not alter the wave function.
-Finally, we can set $r_{k}=\eta$, allowing us to denote the two-mode squeezed state within the open system as,
+Finally, we set $r_{k}=\eta$, allowing us to denote the two-mode squeezed state within the open system as,
 $\phi_{n}(\eta)=(-1)^{n}e^{2in\phi_{k}}\frac{\rm sechr_{k}}{1+\mu_{2}\tanh r_{k}}|1-\mu_{1}^{2}|^{\frac{n}{2}}\frac{(\tanh r_{k})^{n}}{(1+\mu_{2}\tanh r_{k})^{n}}.$ (83)
-This wave function is our central result in this paper whose leading order of $\phi_{n}$ in terms of $\mu_{2}$ is consistent with the wave function described in (39).
-The idea of changing the variable is borrowed from the construction of coherent states, where a key component is the generation of the displacement operator, defined as $D(\eta)=e^{iHt}$, with $H$ being the corresponding Hamiltonian, where one can change $t$ in terms of $r_{k}$ and $\phi_{k}$.
-In Ref.
-Caputa:2021sib , it is stated that the generalized coherent state is derived by constructing a generalized displacement operator according to their group representation theory.
-6.4 The power of wave function (73)
+This wave function constitutes the central result of this paper.
+Its leading-order expansion of $\phi_{n}$ in terms of $\mu_{2}$ is consistent with the wave function presented in (41).
+The idea of performing a variable change is inspired by the construction of coherent states, where a key element is the displacement operator, defined as
+$D(\eta)=e^{iHt},$
+with $H$ being the corresponding Hamiltonian, and where the parameter $t$ can be re-expressed in terms of $r_{k}$ and $\phi_{k}$.
+As noted in Ref.
+Caputa:2021sib , a generalized coherent state can be derived by constructing a generalized displacement operator based on the appropriate group representation theory.
+VI.4 The power of wave function (83)
 We discuss the two-mode squeezed state in an open system, as described in equation (83).
-Unlike the two-mode squeezed state derived from closed system methods, our wave function (73) is more realistic, as it considers a vast open environment.
-Additionally, although our derivation pertains to the early universe, this wave function can also be applied to fields such as quantum optics, quantum information, and condensed matter physics, where information is encoded in the parameters $\mu_{1}$ and $\mu_{2}$ determined by the specific Hamiltonian.
-Finally, we observe that our calculations hold true by requiring that the recurrence relation (55) which is equivalent to the Lanczos algorithm.
-This connection highlights the power of the Lanczos algorithm, which provides a natural method for constructing the wave function according to their Hamiltonian, even for open systems.
-6.5 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$ in open system’s method
-According to the standard procedure, we need to derive the evolution equations for $r_{k}$ and $\phi_{k}$ in relation to conformal time $\eta$.
-The detailed calculations for this derivation can be found in Appendix A.
-The evolution can be summarized as follows:
+In contrast to the two-mode squeezed state derived from closed-system methedology, our wave function (83) is more physically realistic, since it includes the contribution of open system via Hamiltonian.
+Furthermore, although our derivation is obtained within the context of the early universe, this wave function is also widely applicable in fields such as quantum optics, quantum information, and condensed matter physics, where physical information is encoded in the parameters $\mu_{1}$ and $\mu_{2}$, which are determined by the specific system Hamiltonian.
+Finally, we verify the validity of our results by requiring that the recurrence relation (65), which is equivalent to the Lanczos algorithm, be satisfied.
+This connection underscores the utility of the Lanczos algorithm, which offers a systematic framework for constructing the wave function directly from the Hamiltonian, even for open quantum systems.
+VI.5 The evolution of $r_{k}(\eta)$ and $\phi_{k}(\eta)$ in open-system methodology
+Following the standard procedure, we derive the evolution equations for $r_{k}$ and $\phi_{k}$ with respect to conformal time $\eta$.
+The details of this derivation are provided in Appendix A.
+The resulting evolution is summarized by the following equations:
 $\begin{split}r_{k}^{\prime}=&\ \frac{|1-\mu_{1}^{2}|^{\frac{1}{2}}\sinh 2r_{k}\bigl(\frac{a^{2}V_{,\phi\phi}}{2k}\sin 2\phi_{k}-\frac{a^{\prime}}{a}\cos 2\phi_{k}\bigr)-\sinh 2r_{k}\mu_{2}^{\prime}}{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})}\\ \phi_{k}^{\prime}=&\ -\bigl(k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr)+\frac{1}{2}\bigl(\frac{a^{2}V_{,\phi\phi}}{2k}\cos 2\phi_{k}+\frac{a^{\prime}}{a}\sin 2\phi_{k}\bigr)\\ &\ \Bigl[|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}+|1-\mu_{1}^{2}|^{-\frac{1}{2}}(\tanh^{-1}r_{k}+\mu_{2})\Bigr],\end{split}$ (84)
-where prime denotes the varying with respect to the conformal time $\eta$.
-The same as the closed system, we change variable as $y=\log_{10}a$, the evolution equation of $r_{k}$ and $\phi_{k}$ will become as,
+where prime denotes the varying with respect to the conformal time $\eta$. We change into the variable $y=\log_{10}a$, the evolution equation of $r_{k}$ and $\phi_{k}$ will become as,
 $\begin{split}\frac{H_{0}}{\ln 10}10^{-\frac{y}{2}(1+3\omega)}\frac{dr_{k}}{dy}=&\ \frac{|1-\mu_{1}^{2}|^{\frac{1}{2}}\sinh 2r_{k}\bigl(\frac{10^{2y}V_{,\phi\phi}}{2k}\sin 2\phi_{k}-H_{0}10^{-\frac{y}{2}(1+3\omega)}\cos 2\phi_{k}\bigr)-\sinh 2r_{k}\mu_{2}^{\prime}}{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})}\\ \frac{H_{0}}{\ln 10}10^{-\frac{y}{2}(1+3\omega)}\frac{d\phi_{k}}{dy}=&\ -\bigl(k+\frac{10^{2y}V_{,\phi\phi}}{2k}\bigr)+\frac{1}{2}\bigl(\frac{10^{2y}V_{,\phi\phi}}{2k}\cos 2\phi_{k}+H_{0}10^{-\frac{y}{2}(1+3\omega)}\sin 2\phi_{k}\bigr)\\ &\ \Bigl[|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}+|1-\mu_{1}^{2}|^{-\frac{1}{2}}(\tanh^{-1}r_{k}+\mu_{2})\Bigr]\end{split}$ (85)
@@
 $\begin{split}&\ \mu_{2}=k+\frac{10^{2y}V_{,\phi\phi}}{2k},\ \ \ \ \mu_{2}^{\prime}=H_{0}10^{-\frac{y}{2}(1+3\omega)}\frac{10^{2y}V_{,\phi\phi}}{k}+\frac{10^{2y}V_{,\phi\phi}^{\prime}}{2k}\\ &\ |1-\mu_{1}^{2}|=(\frac{10^{2y}V_{,\phi\phi}}{2k})^{2}+(H_{0}10^{-\frac{y}{2}(1+3\omega)})^{2}.\end{split}$
-During the calculations, we accounted for the background of $\phi$ using conformal time $\eta$, as demonstrated by the various potentials in Eq.
+During the calculations, we accounted for the background of $\phi$ using conformal time $\eta$ via preheating, as demonstrated by the various potentials in Eq.
 (24).
@@
 Fig.
-8 clearly illustrates the evolution of $r_{k}$ across various epochs, including inflation, radiation domination (RD), and matter domination (MD).
+8 clearly illustrates the evolution of $r_{k}$ across various epochs, including inflation, RD, and MD.
 It is evident that the different potentials produce similar trends for $r_{k}$.
@@
 These differences will result in distinctive evolutions of Krylov complexity and Krylov entropy.
-6.6 Krylov complexity and Krylov entropy with the approach of open system
-In this section, we will continue use the wave function (73) to explore the Krylov complexity and Krylov entropy.
-All of these calculation are exactly the same with our previous study Li:2024kfm , the only differences is that we need to use $r_{k}$ to take place of $\eta$ in the formulas for Krylov complexity and Krylov entropy whose formulas can be expressed as following:
+VI.6 Krylov complexity and Krylov entropy with open-system methodology
+In this section, we continue our investigation using the wave function (83) to calculate the Krylov complexity and Krylov entropy.
+The procedure for these calculations follows precisely that of our previous study Li:2024kfm .
+The only modification is the substitution of the parameter $r_{k}$ for $\eta$ in the formulas for Krylov complexity and Krylov entropy, which are given as follows:
 $K=\frac{\rm sech^{2}r_{k}|1-\mu_{1}^{2}|\tanh^{2}r_{k}}{[1+2\mu_{2}\tanh r_{k}+(\mu_{2}^{2}-|1-\mu_{1}^{2}|)\tanh^{2}r_{k}]^{2}}$ (86)
@@
 $\begin{split}S_{k}=&\ \frac{\rm sech^{2}r_{k}}{A^{2}}\bigl[(1+\mu_{2}\tanh r_{k})^{2}\ln(1+\mu_{2}\tanh r_{k})^{2}-\tanh^{2}r_{k}|1-\mu_{1}^{2}|\ln\tanh^{2}r_{k}\\ &\ -[(1+\mu_{2}\tanh r_{k})^{2}-|1-\mu_{1}^{2}|\tanh^{2}r_{k}]\ln\rm sech^{2}r_{k}-\tanh^{2}r_{k}|1-\mu_{1}^{2}|\ln|1-\mu_{1}^{2}|\bigr]\end{split}$ (87)
-where we have defined $A=1+2\mu_{2}\tanh r_{k}+(\mu_{2}^{2}-|1-\mu_{1}^{2}|)\tanh^{2}r_{k}$ for simplicity.
-It straightforwardly demonstrates the leading order of Krylov complexity (76) and Krylov entropy (77) will nicely recover into the case of Eq.
+where we have defined $A=1+2\mu_{2}\tanh r_{k}+(\mu_{2}^{2}-|1-\mu_{1}^{2}|)\tanh^{2}r_{k}$.
+It straightforwardly demonstrates the leading order of Krylov complexity (86) and Krylov entropy (87) will reduce to Eq.
 (42) and Eq.
@@
 Fig.
-9 illustrates the evolution of Krylov complexity during inflation, radiation domination (RD), and matter domination (MD).
-During inflation, we observe that Krylov complexity increases exponentially, which is consistent with the closed system’s method.
-However, the trends in RD and MD differ from those shown in Fig.
-4, where the complexity tends to stabilize around certain constant values, oscillating around these values in both RD and MD.
-In the case of an open system, it is evident that Krylov complexity decreases during RD.
-For MD, Krylov complexity initially peaks as the scale factor $a$ increases and then declines to stabilize at constant values.
-Additionally, it is noteworthy that the order of Krylov complexity in an open system is significantly lower than in a closed system, even in RD and MD.
-Our numerical analysis shows that the term $\rm sech(r_{k})$ in Eq.
-(76) leads to this outcome, regardless of whether the system is weakly dissipative or not.
-We have demonstrated that both RD and MD can be approximated as weakly dissipative systems under the condition $\mu<1$.
-Recalling the meaning of Krylov complexity, it describes the operator growth in a many-body system and our early universe is a perfect chaotic system due to $b_{n}=\alpha n$.
-Our numeric reveals that the operator growth is much smaller compared with the case of a closed system, which is consistent with reality since the dissipation will lead to faster decreasing of operator growth compared with that without the dissipation effects.
-Once one understands the role of dissipation effects, it becomes evident that the order of Krylov entropy in an open system is lower compared to that in a closed system, as illustrated in Figs 10 and 5.
-The evolution trend is particularly distinctive for RD and MD, as shown in Figure 10, where the Krylov entropy initially peaks before stabilizing at constant values.
-In the case of MD, the Krylov entropy initially holds steady at certain values before decreasing to some constant level.
-Overall, the background conditions and dissipative effects have a significant impact on both Krylov complexity and Krylov entropy.
-In this section, we utilized the wave function presented in Eq.
-(73) to investigate Krylov complexity and Krylov entropy.
-By comparing our findings to those from a closed system, we observed that the dissipative effects not only influence the overall evolution of the system but also reduce the order of Krylov complexity and Krylov entropy.
-Since Krylov complexity reflects the growth of operators—a fundamental quantum property of a dynamical system—we can conclude that, when the growth of the operator is diminished in the open system compared to the closed system, the dissipative effects lead to a more rapid decoherence-like behavior.
-Meanwhile, we also could see that the various potentials will almost produce the same trend of Krylov complexity and Krylov entropy in different epochs.
+9 illustrates the evolution of Krylov complexity during inflation, RD, and MD.
+During inflation, Krylov complexity exhibits exponential growth, a behavior consistent with the closed-system methodology.
+In contrast, its evolution in the RD and MD epochs differs significantly with the closed-system methodology as shown in Fig.
+4.
+For the open-system methodology, Krylov complexity in RD shows a decreasing trend, while in MD it first peaks as the scale factor $a$ increases and subsequently declines to a constant value.
+Notably, the magnitude of Krylov complexity in the open-system methodology is substantially lower than in the closed-system methodology across both RD and MD.
+Our numerical analysis indicates that this suppression originates from the $\rm sech(r_{k})$ term in Eq.
+(86), an effect presents the dissipation strength.
+We have shown that both the RD and MD epochs can be approximated as weakly dissipative systems, characterized by $\mu_{2}<1$.
+Recalling that Krylov complexity quantifies operator growth in a many-body system, and our earlier finding $b_{n}=\alpha n$ indicates that the unverse is fully chaotic, the significantly reduced operator growth observed in the open-system methodology reveals that: dissipation generically leads to a faster suppression of operator growth compared to that in a closed-system methodology.
+An understanding of the role of dissipation makes it clear that the magnitude of the Krylov entropy in open-system methodology is lower than that in the closed-system methodology, as demonstrated by a comparison of Figs.
+10 and 5.
+The evolutionary trend is particularly distinct during the RD and MD epochs, as shown in Fig.
+10.
+In RD, the Krylov entropy initially rises to a peak before saturating to a constant value.
+In MD, it first remains at a nearly constant level before declining and settling at a lower asymptotic value.
+Overall, both the cosmological background conditions and dissipative effects significantly influence the behavior of Krylov complexity and Krylov entropy.
+In this section, we employed the wave function (83) to study the Krylov complexity and Krylov entropy.
+By comparing our results with those from a closed-system methodology, we observed that dissipative effects not only alter the overall evolution of the system but also suppress the magnitude of both the Krylov complexity and entropy.
+Since Krylov complexity quantifies operator growth, a fundamental quantum feature of dynamical systems, the suppressed growth in the open-system methodology indicates that dissipation induces more rapid decoherence-like behavior.
+Furthermore, we note that different inflationary potentials yield nearly identical trends for both Krylov complexity and entropy across the various cosmological epochs.
 Figure 9: The numerical of Krylov complexity in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.
 Figure 10: The numerical of Krylov entropy in terms of $\log_{10}a$ for three different periods (inflation, RD, and MD), where we set $H_{0}=1$, $k=1$ in inflation, $k=0.01$ at RD and $k=0.005$ at MD for simplicity.
-7 Summary and outlook
-The definition of Krylov complexity is clear and unambiguous, especially when compared to circuit complexity.
-Its uniqueness eliminates the need for introducing a parametric manifold to illustrate the concept of complexity.
-Additionally, Krylov complexity characterizes the growth of operations within a dynamical system, and the order of Krylov complexity can indicate the quantum level of that system.
-Krylov complexity is rooted in the Lanczos algorithm, a powerful method for assessing the nature of dynamical systems, which can be chaotic, integrable, or free.
-The key criteria for this assessment is the Lanczos coefficient, whose formula indicates the characteristics of a given dynamical system.
-For example, the relation $b=\alpha n$ suggests that the system is chaotic.
-In our paper, we highlight that the Lanczos coefficient effectively reveals that our early universe was a many-body, infinite, and chaotic system.
-From a cosmological perspective, the universe is considered an entirely open system.
-Notably, the Lanczos algorithm offers a natural way to construct the wave function of an open system using the second kind of Meixner polynomials.
-Given these insights, it is particularly important to explore Krylov complexity throughout the early universe, including inflation, RD and MD.
-In this paper, we thoroughly examine Krylov complexity in both open and closed system methods.
-Our primary result is the derivation of the two-mode squeezed state using the open system approach, along with the calculation of its evolution equations for $\phi_{k}(\eta)$ and $r_{k}(\eta)$.
+VII Summary and outlook
+The definition of Krylov complexity is precise and unambiguous, particularly in contrast to the various definitions of circuit complexity.
+Its uniqueness remove the need to introduce a parametric manifold to conceptualize complexity.
+Furthermore, Krylov complexity characterizes the growth of operator supported within a dynamical system, and its magnitude indicates the effective quantum size of that system.
+Krylov complexity is fundamentally based on the Lanczos algorithm, a powerful method for diagnosing the nature of dynamical systems, whether chaotic, integrable, or free.
+The key diagnostic is the behavior of the Lanczos coefficient $b_{n}$, whose asymptotic growth with index $n$ serves as an indicator of the system’s dynamical phase.
+For instance, a linear growth $b_{n}\sim\alpha n$ is a hallmark of quantum chaos.
+In our work, we emphasize that the Lanczos coefficient explicitly reveals the early universe as a many-body, infinite, and chaotic system.
+From a cosmological standpoint, the universe is fundamentally an open system.
+Notably, the Lanczos algorithm provides a natural framework for constructing the wave function of an open system using the second kind of Meixner polynomials.
+Given these insights, it is particularly important to investigate Krylov complexity throughout the evolution of the early universe, encompassing the inflationary, RD, and MD eras.
+In this paper, we thoroughly examine Krylov complexity in both open- and closed-system methodologies.
+Our primary result is the derivation of the two-mode squeezed state using the open-system methodology, along with the calculation of its evolution equations for $\phi_{k}(\eta)$ and $r_{k}(\eta)$.
 Below, we present our key findings.
@@
 In Sec.
-2, we introduce the Lanczos algorithm and the associated Lanczos coefficients, which apply equally to both open and closed systems.
-As noted by Parker:2018yvk , the Liouvillian superoperator acts as the Hamiltonian when it can be expressed in terms of creation and annihilation operators.
-Our exploration of Krylov complexity encompasses the entire universe, including inflation, RD, MD, and the violations of slow-roll conditions, which necessitates considering contributions from potential energy.
+II, we introduce the Lanczos algorithm and the associated Lanczos coefficients, a framework applicable to both open and closed systems.
+As highlighted in Parker:2018yvk , the Liouvillian superoperator effectively acts as a Hamiltonian when expressed in terms of creation and annihilation operators.
+Our investigation of Krylov complexity spans the entire evolution of the universe, including the inflationary epoch, RD and MD eras, where the slow-roll conditions are violated, the latter necessitating the inclusion of potential energy contributions.
 Sec.
-3 discusses the most common potentials, such as the Higgs potential, Starobinsky’s inflationary potential, and chaotic inflationary potential.
-These models encompass a wide range of inflationary scenarios due to the unifying characteristics of $\alpha$ attractors.
-To accurately capture the information pertaining to these potentials, we examine the dynamics of the field $\phi$ in relation to the scale factor during the preheating process, as represented in Eq.
+III examines several representative inflationary potentials, such as the Higgs potential, the Starobinsky inflationary potential, and chaotic inflationary models.
+These models collectively cover a broad ranges of inflationary scenarios, unified through the characteristics of $\alpha$-attractors.
+To accurately extract the dynamical information encoded in these potentials, we analyze the evolution of the field $\phi$ relative to the scale factor during preheating, governed by Eq.
 (22).
-From these calculations, we determine the effective mass for various inflationary potentials, as illustrated in Fig.
-2.
-Our findings indicate that the potential approaches certain constant values during RD and MD.
+From this analysis, we compute the effective mass for the various inflationary potentials, as shown in Fig.
+2, where our results indicate that how the effective mass varies with respect to scale factor.
 $(b)$.
-In light of the scale factor and various potentials, we examine Krylov complexity using the closed system method.
-First, we numerically simulate the $r_{k}$ of the two-mode squeezed state, as depicted in Fig.
+Considering the scale factor and various potentials, we investigate Krylov complexity using the closed-system methodology.
+First, we numerically simulate the parameter $r_{k}$ for the two-mode squeezed state, as shown in Fig.
 3.
-The simulation shows oscillations during inflation, which are enhanced in RD and MD, maintaining some constant values.
-Notably, the different potentials produce almost identical trends for $r_{k}$.
-Based on the numerical results for $r_{k}$, we also simulate the corresponding Krylov complexity and Krylov entropy, as illustrated in Figs.
+The results reveal oscillations before horizon exits and then it exponetially grows during inflation, which become enhanced in the RD and MD eras, eventually settling into constant values.
+Notably, different potentials yield nearly identical trends for $r_{k}$.
+Based on the numerical results for $r_{k}$, we also simulate the corresponding Krylov complexity and Krylov entropy, as depicted in Figs.
 4 and 5.
-The trends of both Krylov complexity and Krylov entropy are similar across the corresponding epochs.
+Both Krylov complexity and entropy exhibit similar trends across the respective epochs.
 During inflation, we observe exponential growth, consistent with our previous study Li:2024kfm .
-In the RD and MD phases, the trends show initial enhancement before saturating to constant values.
-Additionally, various inflationary models exhibit nearly the same trends for Krylov complexity and Krylov entropy, confirming the approximation that the effective mass of the inflaton is approximately constant, as indicated in Li:2024ljz .
+In the RD and MD phases, the trends display an initial enhancement followed by saturation to constant values.
+Moreover, various inflationary models demonstrate nearly identical trends for both Krylov complexity and entropy, confirming the approximation that the effective mass of the inflaton is approximately constant, as indicated in Li:2024ljz .
 $(c)$.
-Within the method of an open system, we rigorously derive the wave function of an open system using the second kind of Meixner polynomials, which we name it as the open two-mode squeezed state.
-This derivation is represented in Eq.
+Within the open-system methodology, we rigorously derive the wave function of an open system using the second kind of Meixner polynomials, dubbed as the open two-mode squeezed state.
+This wave function is represented in Eq.
 (83).
-Throughout this process, we discover that the Lanczos coefficient $b_{n}$ and the dissipative coefficient are precisely determined by the specific Hamiltonian, leading to the conclusion that $b_{n}$ is identical across both methods.
+Throughout this process, we discover that the Lanczos coefficient $b_{n}$ and the dissipative coefficient $\mu_{2}$ are precisely determined by the specific Hamiltonian, leading to the conclusion that $b_{n}$ is identical across both methodologies.
 Our findings indicate that $b_{n}=\alpha n$, suggesting that our universe behaves as an entirely chaotic system.
@@
 From Fig.
-7, we can confirm that the RD and MD systems are weakly dissipative, characterized by $\mu_{2}\ll 1$.
-Additionally, we derive the evolution of $r_{k}$ and $\phi_{k}$, noting that the solution for $r_{k}$ is dependent on the solution for $\phi_{k}$ due to their coupling, as described by Eq.
-(74).
-Using the wave function in Eq.
-(73), we calculate the Krylov complexity and Krylov entropy, which are represented in Eqs.
+7, we confirm that the RD and MD systems are weakly dissipative, characterized by $\mu_{2}\ll 1$.
+Additionally, we derive the evolution of the squeezing parameters $r_{k}$ and $\phi_{k}$.
+Their solutions are coupled, as described by Eq.
+(84), meaning the solution for $r_{k}$ depends on that for $\phi_{k}$.
+Using the wave function given in Eq.
+(83), we compute the Krylov complexity and Krylov entropy, presented in Eqs.
 (86) and (87).
-The leading order of these quantities aligns well with the formulas derived from the closed system method.
-In accordance with standard procedures, we need the numerical values of $r_{k}$ within the open system framework, as illustrated in Fig.
+Their leading-order forms are in good agreement with the corresponding formulas derived within the closed-system methodology.
+Following the standard procedure, we require the numerical values of $r_{k}$ in the open-system framework, as shown in Fig.
 8.
-Unlike the closed system, there is no oscillation observed for $r_{k}$ during inflation, and the trends of Krylov complexity and Krylov entropy are depicted in Figs.
-9 and 10.
-Compared to the closed system method, we observe that the values of Krylov complexity and Krylov entropy are significantly smaller.
-Recalling the definition of Krylov complexity, which pertains to the operator growth in a dynamical system, we find that a smaller operator growth suggests a decoherence-like behavior within the Krylov complexity.
-When dissipative effects are considered, we observe decoherence-like behavior, whether the system is weakly dissipative or not.
-While we have achieved a comprehensive understanding of Krylov complexity in the early universe, many ideas still remain to be explored. Here, we outline several future directions.
-Firstly, we investigate the Krylov complexity in the context of single-field inflation, and it is quite natural to extend our work into the multi-field framework Liu:2020zzv ; Liu:2019xhn ; Liu:2021rgq ; Zhang:2022bde as well as $f(R)$ gravity Liu:2018hno ; Liu:2020zlr ; Liu:2018htf .
-This extension allows us to explore the geometrical effects of fields and higher orders of Ricci curvature on Krylov complexity.
-Our Hamiltonian, given by (33), is Hermitian and includes the component of an open system.
-The traditional approach to handling open systems in quantum mechanics involves constructing what is known as a Lindbladian Lindblad:1975ef ; Gorini:1975nb , where the corresponding Hamiltonian is non-Hermitian.
-In our case, a natural method is to expand the action to higher orders with respect to the Mukhanov-Sasaki variable.
-Secondly, our results indicate that the Krylov complexity in a weak dissipative system (RD or MD) will not follow the general discussions mentioned in Sec.
-6.2.
-One possible explanation for this discrepancy is the neglect of decoherence in curvature perturbation during horizon exits.
-Ref.
-Bhattacharjee:2022lzy suggests that decoherence in circuit complexity may lead to this phenomenon of saturation.
-A similar concept might apply to Krylov complexity, particularly concerning the role of decoherence in the construction of the wave function.
-Thirdly, we have already obtained the open two-mode squeezed state as described in (73).
-In light this wave function, we can explore several additional avenues.
-For instance, we can calculate the two-point correlation functions to derive the power spectrum.
-Since the derivation of this wave function is model-independent, we have the flexibility to utilize different models of this wave function to test its validity.
-Conversely, we can also constrain the parameters $\mu_{1}$ and $\mu_{2}$ based on data from the Cosmic Microwave Background (CMB).
-Furthermore, it is important to highlight that Ref.
-Adhikari:2022oxr indicates the potential application of the CV conjecture to Krylov complexity during inflation; however, their approach is tailored for a closed system.
-In contrast, our universe is fundamentally an open system, which provides us with the opportunity to further investigate the validity of the CV conjecture for Krylov complexity within the context of open systems.
-Acknowledgements.
+In contrast to the closed-system case, no oscillations in $r_{k}$ are observed during inflation.
+The resulting trends for Krylov complexity and Krylov entropy are depicted in Figs.
+9 and 10, respectively.
+Compared to the closed-system results, the magnitudes of both Krylov complexity and entropy are significantly reduced.
+Recalling that Krylov complexity quantifies operator growth in a dynamical system, a smaller value indicates suppressed operator growth.
+This suppression reflects a decoherence-like behavior in the context of Krylov complexity once considering the open-system methodology.
+While we have achieved a comprehensive understanding of Krylov complexity in the early universe, many ideas still remain to be explored.
+Here, we outline several future directions.
+Our investigation begins with Krylov complexity in single-field inflation.
+A natural extension of this work involves generalizing it to multi-field frameworks Liu:2020zzv ; Liu:2019xhn ; Liu:2021rgq ; Zhang:2022bde and to theories of $f(R)$ gravity Liu:2018hno ; Liu:2020zlr ; Liu:2018htf .
+Such extensions would allow us to explore the influence of field-space geometry and higher-order Ricci curvature terms on Krylov complexity.
+Our analysis is based on the Hamiltonian given in (35), which is Hermitian and incorporates open-system methodology.
+In conventional quantum mechanics, open systems are often treated using a Lindbladian master equation Lindblad:1975ef ; Gorini:1975nb , typically involving a non-Hermitian effective Hamiltonian.
+In our present framework, a systematic approach involves expanding the action to higher orders in the Mukhanov-Sasaki variable.
+Secondly, our findings demonstrate that in a weakly dissipative system (such as during RD or MD), the behavior of Krylov complexity deviates from the general behavior outlined in Sec.
+VI.2.
+One plausible explanation for this discrepancy is the neglect of decoherence effects in the curvature perturbations upon horizon exit.
+As suggested in Ref.
+Bhattacharjee:2022lzy , decoherence can lead to a saturation phenomenon in circuit complexity.
+A conceptually similar mechanism may govern Krylov complexity, particularly regarding the influence of decoherence on the structure of the wave function.
+Thirdly, having derived the open two-mode squeezed state wave function presented in (83), several promising avenues for further investigation emerge.
+For instance, this wave function can be used to calculate two-point correlation functions and derive the corresponding power spectrum.
+Given that its derivation is model-independent, it offers the flexibility to test its validity by applying it within different theoretical models.
+Conversely, parameters such as $\mu_{1}$ and $\mu_{2}$ can be constrained using observational data from the Cosmic Microwave Background (CMB).
+Furthermore, it is noteworthy that Ref.
+Adhikari:2022oxr suggests the potential applicability of the Complexity=Volume (CV) conjecture to Krylov complexity during inflation, although their analysis is confined to a closed system.
+In contrast, since our universe is fundamentally an open system, our framework provides a distinct opportunity to investigate the validity of the CV conjecture for Krylov complexity within an open-system methodology.
+Acknowledgements
 LH and KH are funded by NSFC grant NO. 12165009, Hunan Natural Science Foundation NO. 2023JJ30487 and NO. 2022JJ40340.
-Appendix A The calculation of $r_{k}$ and $\phi_{k}$ within the method of open system
-In this appendix, we will give the detailed calculations for $r_{k}$ and $\phi_{k}$ based on the open two-mode squeezed state,
+Appendix A The calculation of $r_{k}$ and $\phi_{k}$ within open-system methodology
+In this appendix, we present the detailed derivations and calculations for the parameters $r_{k}$ and $\phi_{k}$ based on the open two-mode squeezed state,
 $|\psi\rangle=\frac{\rm{sech}\ r_{k}}{1+\mu_{2}\tanh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl(\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr)^{n}|n,n\rangle_{\vec{k},-\vec{k}}$ (88)
@@
 $\hat{H}=\bigl(k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr)(\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}+\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}})+\bigl(\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a}\bigr)\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{\vec{k}}^{\dagger}+\bigl(\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr)\hat{c}_{-\vec{k}}\hat{c}_{\vec{k}}$ (89)
-[…]
+Starting from the Schrödinger equation $\hat{H}\psi=i\partial_{\eta}\psi$ (with $\hbar=1$), we first evaluate its left-hand side as follows.
 $\begin{split}\hat{H}|\psi\rangle=&\ \frac{\rm sech~r_{k}}{1+\mu_{2}\tanh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl(\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr)^{n}\Bigl[\bigl(k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr)(\hat{c}_{\vec{k}}\hat{c}_{\vec{k}}^{\dagger}\\ &\ +\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{-\vec{k}})|n,n\rangle_{\vec{k},-\vec{k}}+\bigl(\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a}\bigr)\hat{c}_{-\vec{k}}^{\dagger}\hat{c}_{\vec{k}}^{\dagger}|n,n\rangle_{\vec{k},-\vec{k}}+\bigl(\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr)\hat{c}_{-\vec{k}}\hat{c}_{\vec{k}}|n,n\rangle_{\vec{k},-\vec{k}}\Bigr]\\ =&\ \frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl(\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr)^{n}\Bigl[\bigl(k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr)(2n+1)\\ &\ |n,n\rangle_{\vec{k},-\vec{k}}+\bigl(\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a}\bigr)(n+1)|n+1,n+1\rangle_{\vec{k},-\vec{k}}+\bigl(\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr)n|n-1,n-1\rangle_{\vec{k},-\vec{k}}\Bigr]\\ =&\ \frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl[\bigl(k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr)-|1-\mu_{1}^{2}|^{\frac{1}{2}}e^{2i\phi_{k}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\bigl(\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr)\Bigr]|0,0\rangle_{\vec{k},-\vec{k}}\\ &\ +\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl[\bigl(k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr)-|1-\mu_{1}^{2}|^{\frac{1}{2}}e^{2i\phi_{k}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\bigl(\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr)\Bigr]\\ &\ \sum_{n=1}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl(\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr)^{n}|n,n\rangle_{\vec{k},-\vec{k}}+\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl[\bigl(2k+\frac{a^{2}V_{,\phi\phi}}{k}\bigr)\\ &\ -|1-\mu_{1}^{2}|^{\frac{1}{2}}e^{2i\phi_{k}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\bigl(\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr)-|1-\mu_{1}^{2}|^{-\frac{1}{2}}e^{-2i\phi_{k}}\Bigl(\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr)^{-1}\\ &\ \bigl(\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a}\bigr)\Bigr]\sum_{n=1}^{\infty}(-1)^{n}n|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl(\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr)^{n}|n,n\rangle_{\vec{k},-\vec{k}}\end{split}$ (90)
@@
 $\begin{split}i\frac{d}{d\eta}|\psi\rangle=&\ i\frac{d}{d\eta}\Bigl[\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl(\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr)^{n}|n,n\rangle_{\vec{k},-\vec{k}}\Bigr]\\ =&\ i\Bigl[-\frac{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})r_{k}^{\prime}+\sinh 2r_{k}\mu_{2}^{\prime}}{2\cosh^{3}r_{k}(1+\mu_{2}\tanh r_{k})^{2}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl(\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr)^{n}\\ &\ |n,n\rangle_{\vec{k},-\vec{k}}\Bigr]+i\Bigl[\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\sum_{n=0}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl(\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr)^{n}\\ &\ \bigl[2i\phi_{k}^{\prime}+\frac{2r_{k}^{\prime}}{\sin 2r_{k}(1+\mu_{2}\tanh r_{k})}+\frac{|1-\mu_{1}^{2}|^{\prime}}{2|1-\mu_{1}^{2}|}-\frac{\tanh r_{k}\mu_{2}^{\prime}}{1-\mu_{2}\tanh r_{k}}\bigr]|n,n\rangle_{\vec{k},-\vec{k}}\Bigr]\\ =&\ -i\frac{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})r_{k}^{\prime}+\sinh 2r_{k}\mu_{2}^{\prime}}{2\cosh^{3}r_{k}(1+\mu_{2}\tanh r_{k})^{2}}|0,0\rangle_{\vec{k},-\vec{k}}\\ &\ -i\frac{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})r_{k}^{\prime}+\sinh 2r_{k}\mu_{2}^{\prime}}{2\cosh^{3}r_{k}(1+\mu_{2}\tanh r_{k})^{2}}\sum_{n=1}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl(\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr)^{n}\\ &\ |n,n\rangle_{\vec{k},-\vec{k}}+\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl(-2\phi_{k}^{\prime}+i\frac{2r_{k}^{\prime}}{\sin 2r_{k}(1+\mu_{2}\tanh r_{k})}+i\frac{|1-\mu_{1}^{2}|^{\prime}}{2|1-\mu_{1}^{2}|}\\ &\ -i\frac{\tanh r_{k}\mu_{2}^{\prime}}{1-\mu_{2}\tanh r_{k}}\Bigr)\sum_{n=1}^{\infty}(-1)^{n}n|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl(\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr)^{n}|n,n\rangle_{\vec{k},-\vec{k}}\end{split}$ (91)
-[…]
+Substituting Eqs. (90) and (91) into the Schrödinger equation yields the following simplified form for the open system’s squeezed state:
 $\begin{split}&\ A_{1}|0,0\rangle_{\vec{k},-\vec{k}}+A_{1}\sum_{n=1}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl(\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr)^{n}|n,n\rangle_{\vec{k},-\vec{k}}\\ &\ +A_{2}\sum_{n=1}^{\infty}(-1)^{n}n|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl(\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr)^{n}|n,n\rangle_{\vec{k},-\vec{k}}\\ =&\ B_{1}|0,0\rangle_{\vec{k},-\vec{k}}+B_{1}\sum_{n=1}^{\infty}(-1)^{n}|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl(\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr)^{n}|n,n\rangle_{\vec{k},-\vec{k}}\\ &\ +B_{2}\sum_{n=1}^{\infty}(-1)^{n}n|1-\mu_{1}^{2}|^{\frac{n}{2}}e^{2in\phi_{k}}\Bigl(\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr)^{n}|n,n\rangle_{\vec{k},-\vec{k}}.\end{split}$ (92)
@@
 $\begin{split}&\ -i\frac{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})r_{k}^{\prime}+\sinh 2r_{k}\mu_{2}^{\prime}}{2\cosh^{3}r_{k}(1+\mu_{2}\tanh r_{k})^{2}}\\ =-\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}&\ \Bigl[|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\bigl(\frac{a^{2}V_{,\phi\phi}}{2k}\sin 2\phi_{k}-\frac{a^{\prime}}{a}\cos 2\phi_{k}\bigr)\Bigr].\end{split}$ (94)
-and therefore, we could obtain the evolution of $r_{k}$ as,
+and therefore, we obtain the evolution of $r_{k}$ as,
 $r_{k}^{\prime}=\frac{|1-\mu_{1}^{2}|^{\frac{1}{2}}\sinh 2r_{k}\bigl(\frac{a^{2}V_{,\phi\phi}}{2k}\sin 2\phi_{k}-\frac{a^{\prime}}{a}\cos 2\phi_{k}\bigr)-\sinh 2r_{k}\mu_{2}^{\prime}}{(\sinh 2r_{k}+2\mu_{2}\cosh^{2}r_{k})}.$ (95)
@@
 $\begin{split}&\ \frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl[-2\phi_{k}^{\prime}+i\frac{2r_{k}^{\prime}}{\sin 2r_{k}(1+\mu_{2}\tanh r_{k})}+i\frac{|1-\mu_{1}^{2}|^{\prime}}{2|1-\mu_{1}^{2}|}-i\frac{\tanh r_{k}\mu_{2}^{\prime}}{1-\mu_{2}\tanh r_{k}}\Bigr]\\ =&\ \frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl[\bigl(2k+\frac{a^{2}V_{,\phi\phi}}{k}\bigr)-|1-\mu_{1}^{2}|^{\frac{1}{2}}e^{2i\phi_{k}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\bigl(\frac{a^{2}V_{,\phi\phi}}{2k}-i\frac{a^{\prime}}{a}\bigr)\\ &\ -|1-\mu_{1}^{2}|^{-\frac{1}{2}}e^{-2i\phi_{k}}\Bigl(\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigr)^{-1}\bigl(\frac{a^{2}V_{,\phi\phi}}{2k}+i\frac{a^{\prime}}{a}\bigr)\Bigr],\end{split}$ (96)
-the real component can read as,
+the real component reads as,
 $\begin{split}-\frac{2\phi_{k}^{\prime}\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}&\ =\frac{\rm sech\ r_{k}}{1+\mu_{2}\tanh r_{k}}\Bigl[\bigl(2k+\frac{a^{2}V_{,\phi\phi}}{k}\bigr)\\ -\bigl(\frac{a^{2}V_{,\phi\phi}}{2k}\cos 2\phi_{k}+\frac{a^{\prime}}{a}\sin 2\phi_{k}\bigr)&\ \Bigl(|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}+|1-\mu_{1}^{2}|^{-\frac{1}{2}}(\tanh^{-1}r_{k}+\mu_{2})\Bigr)\Bigr].\end{split}$ (97)
-Finally, we could obtain the evolution equation of $\phi_{k}$ as follows,
+Finally, we obtain the evolution equation of $\phi_{k}$ as follows,
 $\begin{split}\phi_{k}^{\prime}=&\ -\bigl(k+\frac{a^{2}V_{,\phi\phi}}{2k}\bigr)+\frac{1}{2}\bigl(\frac{a^{2}V_{,\phi\phi}}{2k}\cos 2\phi_{k}+\frac{a^{\prime}}{a}\sin 2\phi_{k}\bigr)\\ &\ \Bigl[|1-\mu_{1}^{2}|^{\frac{1}{2}}\frac{\tanh r_{k}}{1+\mu_{2}\tanh r_{k}}+|1-\mu_{1}^{2}|^{-\frac{1}{2}}(\tanh^{-1}r_{k}+\mu_{2})\Bigr].\end{split}$ (98)
-One can easily check that Eqs. (85) and (88) could nicely recover into Eq. (35) as $\mu_{1}=\mu_{2}=0$.
+It can be verified that Eqs. (95) and (98) correctly reduce to Eq. (37) in the limit $\mu_{1}=\mu_{2}=0$.